A Complete Bibliography of The Journal of the American Statistical Association: 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

26 August 2024 Version 1.10

Title word cross-reference

$105.00 [McD20]. $109.00 [WFL22]. $110.00 [Ber24]. $119.50 [Ism20]. $119.95 [Beb20]. $119.99 [Qiu24]. $120.00 [Ni22, Seo24]. $129.95 [Ché20b]. $130.00 [Muk24]. $140.00 [All20]. $149.99 [WFL22]. $150.00 [Cho22]. $174.95 [Che20a, Zha20]. $200.00 [Ni23]. $250.00 [Won24]. $270.00 [Cha23]. 2^{K} [BP23]. $32.00 [AS20]. $39.99 [Che20d]. $49.95 [Par22]. $52.99 [Sen24]. $53.00 [Rob24]. $59.95 [Cho22]. $64.95 [HM23]. $69.95 [Sch20]. $69.99 [Bor24]. $74.95 [Ye24]. $79.95 [Sto22]. $80.00 [Gio23]. $96.00 [Ni22]. $99.95 [Beh20, Ell23, Hor22, LN20, Len20, Rah24, Wan20, Zha22]. ^{K} [LRY24]. _{2.5} [LZCH21]. c [TDB24]. D [LD20]. i [SLgX20]. K [AS21, PRV22]. L^{2} [ZGZC20]. M [SLZ24]. n [LM21, NS23]. p [Hof22b, LM21, LX20, NS23]. R [HLL22, Jew21]. X [Ano22c, RT20a, RT20b, SH20].

-Estimation [HLL22, SLZ24]. -Norm [ZGZC20, AS21]. -Probabilities [LD20]. -Sample [PRV22]. -Value [LX20, TDB24]. -Values [Hof22b]. -Zero- [LM21].

.T [Rug21].

/978 [WFL22]. /978-1-4939-8853-2 [WFL22].

0 [Ell23, Rah24, Rob24].

1 [Bor24, Ni23, Par22, Seo24, Wan20]. 19 [Jew21, QDL21a, TFCF21, WWLT24]. 1990 [NMSD21].

2 [Che20d, LRY24, McD20, Sen24]. 2023 [Ano24]. 2D [ZNBR22]. 2nd [Len20, Par23, Qiu24].

3 [Gio23, LN20, Len20, Ye24]. 3rd [Sto22].

4 [HM23].

5 [All20].

6 [Ber24, Cha23, Har23, Hor22].

7 [Beb20, Ché20b, Ism20, Qiu24, Zha20].

8 [Beh20, Che20a, Muk24, Won24, Zha22].

9 [AS20, Sch20, Sto22]. 978 [All20, AS20, Beb20, Beh20, Ber24, Bor24, Cha23, Ché20b, Che20d, Che20a, Ell23, Gio23, Har23, HM23, Hor22, Ism20, LN20, Len20, McD20, Muk24, Ni23, Par22, Qiu24, Rah24, Rob24, Sch20, Sen24, Seo24, Sto22, Wan20, Won24, Ye24, Zha20, Zha22]. 978-0-19-284726-3 [Gio23]. 978-0-192-85902-0 [Rob24]. 978-0-36-714639-9 [Sch20]. 978-0-36-756995-2 [Cho22]. 978-0-367-17387-6 [Har23]. 978-0-367-27950-9 [Sto22]. 978-0-367-48875-8 [Muk24]. 978-0-367-54376-1 [Ni23]. 978-0-367-55419-4 [HM23]. 978-0-367-60952-8 [Won24]. 978-0-367-70505-3 [Ye24]. 978-0-367-74845-6 [Hor22]. 978-0-46-509760-9 [AS20]. 978-0-674-97523-1 [Par22]. 978-0-81-537944-7 [Beb20]. 978-0-81-538710-7 [Ché20b]. 978-1-00-310027-0 [Cho22]. 978-1-032-32562-0 [Rah24]. 978-1-10-870111-2 [Che20d]. 978-1-108-48029-1 [Bor24]. 978-1-108-99413-2 [Sen24]. 978-1-11-807858-7 [Ism20]. 978-1-13-873235-3 [LN20]. 978-1-138-05859-0 [Ell23]. 978-1-138-10640-6 [Cha23]. 978-1-4398-0008-9 [Ni22].

978-1-4398-0010-2 [Ni22]. 978-1-46-651520-8 [Beh20]. 978-1-46-655332-3 [Len20]. 978-1-46-655503-7 [Zha20]. 978-1-466-58281-1 [Seo24]. 978-1-482-21945-6 [Ber24]. 978-1-49-870447-2 [McD20]. 978-1-49-876301-1 [Wan20]. 978-1-49-876977-8 [Zha22]. 978-1-49-877321-8 [Che20a]. 978-1-49-878862-5 [All20]. 978-1-4939-8851-8 [WFL22]. 978-1-4939-8853-2 [WFL22]. 978-3-031-17482-7 [Qiu24].

A/B [SWL^{+}23]. Aberrant [WJIS20]. Abnormal [LZS23]. Abrupt [RRRF22]. Absence [CJM^{+}20b, HS22]. Absolute [LTO^{+}21]. Abundance [LTO^{+}21, WJIS20]. Accelerated [NS23]. Accelerating [MML23]. Acceptance [Cho22]. Access [ARN20]. Accommodating [LSN23]. According [ZSL^{+}23]. Accounting [MN22, ZHC21]. Accuracy [Dua23]. Accurate [FGW23, HJO^{+}20, WJIS20, YB24]. ACE2 [SJW23]. Act [NMSD21]. Action [CWF^{+}24, Ber24]. Active [DYFQ23, XWMD24]. Activity [BMDG20, MLH^{+}22]. Acyclic [LSP20]. AdaBoost [LLWL22]. Adaptive [BMS20, BST24, CSX22, FGQ24, GSW23, GCW20, LLM23a, LZZ23b, LTY21, LH22, MQLH20, MLZH24, QIL24, QW21, SZF20, XLYT20, ZDM21, ZWS22, ZC22, ZWLS20, DL21]. Adaptive-to-Model [LZZ23b]. Adaptively [DJM23]. AdaSmooth [MOA24]. Additive [BMA^{+}22, CJK^{+}24, DLL23, FWZ^{+}21, HMP20, HBX23, LLLZ23, LKL20, MR20, MOA24, Mur21, TLX24, YYZ24]. Addressing [PBOD23, TLT^{+}24]. Adequate [FLY24]. Adhesion [SHR^{+}20a]. Adjust [ZD24]. Adjusted [FGZ22, FGJS22, LL23, YRTJ23]. Adjustment [EB21, LRY24, WSM^{+}23, YQL23, YSYZ23]. Adjustments [FNDK23, Won24]. Administrative [BWSB24]. Adolescent [WCZZ24]. Adoption [ST21]. Adults [PLCS21, ZCZZ23]. Advanced [Kan21]. Advantage [SLL^{+}24]. Advections [SLG23]. Adversarial [INGP23]. Advertising [JPR20]. Affect [RBMB21]. Affine [SW21a]. Affine-Equivariant [SW21a]. African [XTK^{+}22]. Against [GPPV20]. Age [CPH22, JCY^{+}21]. Age-Related [JCY^{+}21]. Agencies [Dre23]. Agencies-Are [Dre23]. Aggregating [LXZZ23]. Aggregation [DGSZ23, LXLF20, ZJLD23]. Agnostic [WGSC23]. Agreement [Ism20]. Agresti [Hor22]. Aided [WC24]. Air [GJK^{+}20, LSD21, MKW^{+}24, NMSD21]. al [LZ21b, ZP21]. Alan [Hor22]. Albert [Zha20]. Alert [STC^{+}20]. Algebraic [BF24]. Algorithm [HFR20, MZPD22, MC22, MOA24, QIL24, WGSC23, YRTJ23]. Algorithm-Agnostic [WGSC23]. Algorithms [Ize21, SWS22, Vat20, WZC23]. Alignment [BK20]. Allele [KTDS20]. Allocation [DLS24, KPV22, NMJ20]. Allowing [WJIS20]. ALMOND [QW21]. Almudevar [Muk24]. Alzheimer [SZZ24, YWKZ22]. Amendments [NMSD21]. American [Ano24, SHWB23, XTK^{+}22]. Americans [LWLL23]. Analgesia [ZHC21]. Analogues [SW21a]. Analyses [BWSB24, LTO^{+}21, MV20]. Analysis [ARN20, AL23, BN21, BK22, BIK20, BTHK20, CLX22, Ché20b, CMZ20, CLZ20, CLLY22, CHYL23, CHDP23, Coo20, CMC^{+}20, DL23a, DDX22, DCGM23, DJM23, DG23, Efr22, FMM22, FD21, Fog20, FDF20, Gho22, GX23, GJK^{+}20, GS23, GLLZ22, HJO^{+}20, Har23, HS21, Hof22a, HM23, HY24, IW22a, IW22b, Imb22, JM20, JT22, KML23, KLLL23, Kow22, KHK21, LL22, LLZZ22, LZZ23a, LA20, LZX^{+}24, LZS23, MLL21, MC22, MM21, MV20, MYB^{+}20, MNW^{+}20, PXZ21, PK23, Pen22b, SWZ20, SMMH21, TBQ20, THXZ23, WRG23, WCZZ24, Wei20, XLLJ20, XC23a, XS22, XTK^{+}22, XYY24, YBRM20, YSWB24, YYZZ20, ZSL20, ZK23, ZWX^{+}23a, ZWX^{+}23b, ZBM^{+}24, ZX24, ZDNH23, ZLGY23, ZHC21, Sto22, Che20c]. Analytic [LX20]. Analytics [Mar21, Qiu24, Rah24]. Analyzing [KG23, XCZW23, YSYZ23]. Anastasios [Zha22]. Ancestry [ZYZ24]. Andreas [Bor24]. Angle [FLLZ22, QLFL20, XZZ^{+}22]. Angle-Based [FLLZ22, QLFL20, XZZ^{+}22]. Animal [GBL^{+}21]. Ann [Che20a]. Annual [WHHZ22]. Anomaly [SCACvdH24]. ANOVA [SWPH20]. Antecedents [AM22]. Anthony [Muk24]. Anytime [GHL24]. Anytime-Valid [GHL24]. Application [BGOT20, BMDG20, CXH23, CPH22, CMC^{+}20, DGD21, DP21, DTQ23, FDNZ24, GKB^{+}21, GRLB20, GR21, GYY^{+}22, HL21b, JPR20, JJ21, JZW23, LZ21a, LA20, LLR21a, LLR21b, LLM23b, LDL^{+}22, LZX^{+}24, MZW23, MKV^{+}24, NCMK20, OPI22, OWP20, SLC21, SHWB23, SZZ24, SL20, SHR^{+}20a, TBQ20, TET^{+}23, WW22, WCL21, WST^{+}20, YLXL23, ZSL20, ZDNH23, ZQL^{+}23, ZHC21, ZLGY23]. Applications [BST24, CLS20, DJM23, FL22, HW21, JCY^{+}21, LLL22a, LKM21, LLR21c, LSB^{+}22, LLL22b, MHZ23, McD20, MAZB22, PRV22, PXZ21, PLZ24, SL22b, Ste23, Vat20, YWKZ22, ZC22, ZLQ^{+}21, dMMV21, Ism20, LN20, Wan20]. Applied [CCH20, Har23]. Approach [BMDG20, CSX22, CXH23, CX23, CJZ23, Coo20, CT21b, DLXL23b, DLXL23c, DGD21, DG20, FMW20, GT23, GX23, GCW20, Gug23, GZW24, Jan23, JT22, KRW24, LB23, LS20, LWY^{+}21, LHS22, LCL22, LSN23, LYZ23, LY22, LDL^{+}22, LZH24, Loh20, MLL21, MYB^{+}20, MR23, MNBCCC24, RBMB21, Rob24, SBK20, SCACvdH24, THP^{+}23, WPB^{+}20a, WPB^{+}20b, WST^{+}20, XZWT21, XC23b, YDV24, ZM23, ZB22, ZDNH23, ZJLH23]. Approaches [HL21a]. Approximate [Fra20, HFR20, PT23]. Approximating [LM21]. Approximation [DL22a, DL21, Dua23]. Approximations [JVRS23]. Arbitrary [LX20]. Archetypal [JM20]. Areas [RLS23]. Armed [QLFL20, QIL24]. Arms [NLS^{+}21]. Arrays [CKS23, ST23b]. Arterial [ARN20]. Arthur [Sha21]. Article [gXZ20, JLS20a]. Artificial [MM21]. Assessing [GWL^{+}23, LLYM21, LWLL23, SBY24]. Assessment [ZDY23]. Asset [DLS24]. Assimilated [HLS^{+}21]. Assisted [DGH23, HMW23, IS23a, ST23a, WWF23]. Associate [Ano24]. Associated [FBZ21, GWL^{+}23]. Association [CCLS21, CPH22, JCY^{+}21, LTO^{+}21, LLL22a, LHW23, LZZ23a, LLYM21, LLL22b, SL20, XSP23, ZZ22, ZDNH23, ZQL^{+}23, Ano24]. Associations [LO21]. Assumption [LSRR21, PK23, VDVM24]. Assumption-Lean [PK23, VDVM24]. Astrophysics [DREM^{+}21]. Asymmetric [YRTJ23]. Asymmetry [AM22, CPH22]. Asymptotic [FFHL22, WPTM24]. Asymptotically [BGOT20]. Asymptotics [FNDK23, LW22b]. Asynchronous [CMZ20, LLZZ22]. Athanasios [Beh20]. Atoms [CSdG^{+}23, DCGM23]. Attribute [GX23]. Attributes [WCNL24]. Attribution [CS20, Efr20a, FHT20, WJW^{+}21, gXZ20]. Augmented [BMFR21, DYFQ23, He24]. Aurore [Cha23]. Authors [WHC^{+}24]. Autism [LLR21c, LWSL22]. Auto [TFS21]. Auto-G-Computation [TFS21]. Autocorrelated [RRRF22]. Automated [SLC21]. Automatic [ZMZ24]. Autoregressive [BSM23, LRZ24, WZLL22, WBBM23]. Auxiliary [XS21]. Average [Fog20, Li20, LKM21, SST21, YPL^{+}21, ZD24]. Averages [WBBM23]. Averaging [HMZZ23, LLWL22, ZZLC20].

B [CS20, Zha22, SWL^{+}23]. Babette [Ye24]. Back [Mar21]. Backfitting [HMMMN21, YYZ24]. Backtesting [SQPQ21]. Bagged [IAPK23]. BAGS [LTY21]. Balance [FAS^{+}22]. Balancing [DG23, JRSW22, LH22]. Ball [PWZ^{+}20]. Balser [Beb20]. Banach [PWZ^{+}20]. Band [BTHK20]. Bandit [CLS21a, CWF^{+}24, QIL24]. Bands [Ano23a, LKL20, PS23]. Based [BAXV22, CCLS21, CX23, DS23, DV24, DWHS23, DKPW21, FLLZ22, GVW20, GSS22b, GKB^{+}21, GMVFX22, HMZZ23, JCY^{+}21, LJL23b, LLR21c, MLH^{+}22, MZPD22, MR20, MOA24, MZW23, MKV^{+}24, MR23, MKZ^{+}21, MAZB22, QLFL20, QZ22, SPMK22, SLS20, SWZ20, SCACvdH24, SL20, Wei20, XZZ^{+}22, YLPC20, ZGZC20, ZBM^{+}24, ZLZ24, ZLZ23a, ZLQ^{+}21, KY20, LWLL23, LRY24, SJW23, WPTM24, YPL^{+}21, ZKR23, LN20]. Bases [AHDP24]. Basic [AS20]. Batch [WWW22]. Batches [LZS23]. Bayes [Efr22, Gho22, Imb22, Pen22b, XS22, Cre22, DL22a, DH22, FDAZ22, Hof22a, IW22a, IW22b, ISSM23, RS22, YDV24, YB24]. Bayesian [Har23, Ni23, Rob24, ARN20, AFDZ23, BDGJ23, BMD^{+}24, BCI24, BST24, BHW20, BMSL22, CFMB24, CKK24, CSdG^{+}23, CJM^{+}20b, CMW24, DGD21, DCGM23, FLSZ20, FD21, Fra20, FNDK23, GSS22b, GGR^{+}20, GRLB20, Gug23, GR21, HSB21, HARF^{+}20, HBD24, HL21a, HQYZ20, INGP23, JHB20, KS24, LWY^{+}21, LLM23a, LPR24, LTY21, LKPG21, MLH^{+}22, MJ23, MC22, MCM20, MBCK22, MANW20, MYB^{+}20, MNHG24, MML23, Mur21, NDH^{+}24, NMJ20, NR23, NS23, OWP20, PLL23, PLCS21, PBF22, RMM20, RBMB21, RPLG22, SBK20, SPMC21, SL22c, WBK^{+}22, WST^{+}20, WGK24, XX20, XX21, ZNBR22, ZBM^{+}24, ZHC21, ZROD23, Ni22, Par23]. Being [FBZ21]. Benefits [DL21, FMM22]. Benford [BCCP22]. Benign [LSS23]. Berk [SL20]. Best [CWF^{+}24, RLS23]. Better [YSYZ23]. Between [CPH22, KLLL23, LBN23, LHW23, LLYM21, RLYI20, YZL22, ZQL^{+}23]. Between-Study [RLYI20]. Beyond [AM22, MB23]. Bias [Ano22c, EB21, JRR22, LL23, LM21, PBOD23, RT20a, RT20b, SWR24, SH20, XZLG24]. Bias-Adjusted [LL23]. Bias-Correction [XZLG24]. Bias-Variance [Ano22c, LM21, RT20a, RT20b, SH20]. Biased [FLKK21, RB20, TMZ22].

Biases [KMPT21]. Big [Ban22, KG23, LX24, MNW^{+}20, TJK22, ZXS^{+}22a, Ché20b]. Big-Data [LX24]. Binary [BMSL22, CGM23, GSS23a, GK22, HJO^{+}20, HMW23, PLZ24, SHR^{+}20a, SHR^{+}20b]. Bing [McD20]. Biobank [GKB^{+}21]. Biological [OWP20]. Biomarker [WRG23]. Biomedicine [Zha20]. Biomonitoring [LWLL23]. Biostatistics [Ni22]. Bird [PBOD23]. Bird-Plant [PBOD23]. Bivariate [TET^{+}23, WNK24]. Blind [CFWY23, WGK24]. Blinder [GB23]. Bliss [Beb20]. Block [HQYZ20, HZQ^{+}21, LL23, LRY24, TBC20, XQ21, YLSL20]. Block-missing [YLSL20]. Block-Wise [XQ21]. Blocked [SPMK22]. Blockmodel [OPI22, SCMW24]. Blood [ZLGY23]. Boca [All20, Beb20, Beh20, Ber24, Cha23, Ché20b, Che20a, Cho22, Ell23, HM23, Hor22, LN20, Len20, McD20, Muk24, Ni22, Ni23, Rah24, Sch20, Seo24, Sto22, Wan20, Won24, Ye24, Zha20, Zha22]. Book [Agr23, All20, AS20, Beb20, Beh20, Ber24, Bor24, Cha23, Ché20b, Che20c, Che20d, Che20a, Chi21, Cho22, Ell23, Gio23, Har23, HM23, Hor22, Ism20, LN20, Len20, McD20, Mor21, Muk24, Ni22, Ni23, Par22, Par23, Qiu24, Rah24, Rob24, Sch20, Sen24, Seo24, Sto22, Wan20, WFL22, Won24, Ye24, Zha20, Zha22]. Books [AS20]. Boolean [SL22a]. Boosting [BKR^{+}23]. Bootstrap [KRW24, NR23, PS23, RLY^{+}23, ZKR23, Ano23a]. Bootstrapping [LLM23b, dHZ24]. Border [RBMB21]. Both [CFWY23]. Bottom [LHS22]. Bottom-Up [LHS22]. Boundary [FSV20, YZ20]. Bounded [WQ23]. Bounds [GSS22a, GSS23b, GSS23a, WTW20]. Brad [Cox20]. Bradley [FHT20, gXZ20]. Brain [AL23, CHDP23, FBZ21, GR21, HZLG22, LWY^{+}21, LKPG21, MLH^{+}22, MNW^{+}20, WCZZ24, ZSL20, ZK23, ZWX^{+}23a, ZWX^{+}23b, ZCZZ23]. Brain-Computer [MLH^{+}22]. Branching [LHS22, LTB22]. Break [SS22]. Breaking [GLM23]. Breakpoints [CWW22]. Breast [SL20]. Bridge [CMZ20, WJM22]. Broken [ZWLS20]. Bronchiolitis [HBP^{+}20]. Brumback [Ye24]. Budget [PLSZ23]. Building [CCLS21]. Bulk [KML23, SJW23, SZZ24]. Buyu [Jan23, ZM23].

CACE [ZHC21]. Caiado [Che20a]. Calculation [LX20]. Calibrated [FWZ^{+}21]. Calibration [SHR^{+}20a, THP^{+}23, XX21]. Cambridge [Bor24, Che20d, Par22, Sen24]. Can [LO21]. Cancer [PXZ21, PLZ24, SL20, WBK^{+}22, ZDNH23]. Canonical [LC20, SWZ20]. Capture [CSV21, DKJ24, MA23]. Capture-Recapture [DKJ24, MA23]. Cardinality [DGW23]. Cardiovascular [TTPD20]. Carlo [DHJW22, DL22a, Dua23, GSS22b, NF21]. Carlos [Bor24]. Case [AM21, HS22, LNB22, Ste23]. Cases [TFCF21]. Categorical [Har23, MR23, XLYT20]. Categories [LLWL22, WWLT24]. Category [XLYT20, ZDNH23]. Category-Adaptive [XLYT20]. Catholic [KSR21]. Cauchy [LX20]. Causal [ABD^{+}21, BLLY22, Ché20b, CBK^{+}22, CPS^{+}24, DWHS23, GSS22a, GSS23b, HGC24, IS23b, IJM21, LSP24, LSB^{+}22, NLS^{+}21, OSDvdL24, SBY24, SWL^{+}23, SYD^{+}22, TFS21, VB23, Wes22, Won24, XSP23, YD20, Ye24, NMSD21]. Causality [AM22]. Cause [AS20]. Cautionary [Gio23]. CCA [SWZ20]. Cell [HL21b, JJ21, LLR21a, LLR21b, LWSL22, MKV^{+}24, MAZB22, PLZ24, SZZ24, SHR^{+}20a, WJIS20]. Cell-Line [PLZ24]. Cell-Type-Specific [SZZ24]. Censored [CSX21, CS24, FZHL23, FGRS23, JCYS20, LSX23, MR20, XZZ^{+}22, ZWLS20]. Censoring [Ano22b, DV24, GLLT21, LHW23, MST21, MST22, RB20]. Center [HLL22, HHH23, NLS^{+}21, SDH22]. Center-Outward [HLL22, HHH23, SDH22]. Central [LRY24, SPMK22]. Cesarean [ZHC21]. Chacón [Wan20]. Chain [GSS22b, NF21]. Challenges [CSV21]. Cham [Qiu24]. Chang [Beb20]. Change [BSM23, DG20, DPY22, Fry24, Mie23, ZWS22, CRYZ23]. Change-Points [Fry24, CRYZ23]. Changepoint [CWS24]. Changes [GVW20, RRRF22, WJW^{+}21]. Chapman [All20, Beb20, Beh20, Ber24, Cha23, Ché20b, Che20a, Cho22, Ell23, LN20, Len20, McD20, Muk24, Ni22, Ni23, Rah24, Sch20, Seo24, Sto22, Wan20, Won24, Ye24, Zha20, Zha22]. Characteristic [KY20]. Characteristics [YZL22]. Characterization [Yee22]. Characterizations [BCCP22]. Characterize [GZW24, WBK^{+}22]. Chen [LT22, Rah24]. Chengguang [ZM23]. Chenguang [Jan23]. Chi [DH22]. Chi-Squared [DH22]. Childhood [GLLZ22]. China [LZCH21]. Choudhary [Ism20]. Christou [Beh20]. Chronic [GYY^{+}22]. Churn [BMDG20, GLLT21]. Cities [ILMY22]. Claims [SZS24b]. Class [CJM^{+}20b, DG20, DW21, HHZ21, JGED21a, JGED21b, LV21, MB23, Rug21, SST21, Sha21, Wil21]. Classes [GPPV20]. Classical [HL21a]. Classification [CXH23, FLLZ22, KR23, LX24, WWLT24, XZWT21, ZDY23, ZSJW23, ZLD23, Che20a]. Classified [MJ23]. Clean [NMSD21]. Client [GLLT21]. Climate [DP21, SSB21, WJW^{+}21]. Clinical [Agr23, Beb20, CSdG^{+}23, CMC^{+}20, Ell23, GH21, RLYI20, SMMH21, WSM^{+}23, YSYZ23, YWKZ22]. Clinically [FDNZ24]. Closed [BK20]. Cluster [ACV20, DREM^{+}21, PK23]. Clustered [LZS23, SPMK22, Sug21, XLWG23]. Clustering [BDGJ23, Che20a, GBW24, GYY^{+}22, HTF23, LL23, LPR24, MZPD22, NDH^{+}24, PSBB21, PXZ21, SHO20, WC24, ZPYZ24]. Co [SZZ24, AB20]. Co-Expression [SZZ24]. Co-Movement [AB20]. Coefficient [Cha21, FBFV21, LZX^{+}24, LRZ24, PLL23, WNK24, ZLZ24, ZCM22]. Coefficients [GK22]. Cohesion [NDH^{+}24]. Cohort [XTK^{+}22]. Coletti [SF23, Yon23]. Collaborative [DSWQ21, ZSJW23]. Collaborators [Ano20b]. Collected [DJM23, SMMH21]. Colombia [CBK^{+}22]. Colorectal [ZDNH23]. Combination [LX20]. Combined [YLX24]. Combining [KMPT21, YD20]. Comment [Cre20, FMW20, Han21, Imb22, LV21, LT22, Loh20, Peñ22a, Rug21, Sha21, Wag20, ZM23]. Comments [HHZ21, LYZ23, OY22, TJK22]. Common [CSdG^{+}23, DCGM23]. Communication [FGW23, Par22]. Communication-Efficient [FGW23].

Communities [JKLW23]. Community [LLB^{+}22, SHWB23, WZL^{+}23, YS21, ZHW22, ZW23]. Comparative [NLS^{+}21]. Comparing [AS21, LD20, PK20, TDB24]. Comparison [GCRZ24, HXC23]. Comparisons [LTY21]. Competing [BKR^{+}23, SYD^{+}22]. Completely [ACV20]. Completion [ABD^{+}21, BN21, IS23a]. Complex [CCH20, CHKW24, KRW24, LC20, YZ22]. Complier [RSvdL21]. Component [ASSS21, CMZ20, CLLY22, DSGS22, GS23, HY24, KLLL23, KHK21, MNW^{+}20]. Components [HMMMN21]. Compositional [LTB22, MCM20, SW23, TJH^{+}23, ZMNVF22]. Compositions [ZPZW23]. Compound [SL22b]. Compromise [HVL23]. Computation [Fra20, SSL22, TFS21]. Computational [WZC23]. Computer [MLH^{+}22, NCMK20, SHR^{+}20a, SHR^{+}20b, SWPH20, THP^{+}23, WSX22, XX21]. Computing [HFR20, Len20]. Concentrations [WHHZ22]. Concepts [Gio23]. Concise [Sch20]. Conditional [BHW20, CKK24, DHL20, GHL24, HL24, JLS20a, JLS20b, KY20, KL21b, KKS23, LJL^{+}23a, LLM23a, LX24, SRS^{+}23, SQPQ21, TCYL23, XS21, YPL^{+}21, ZJLH23]. Confidence [Efr22, Gho22, IW22a, Imb22, KL21b, Lei20, LLgX22, LKL20, Pen22b, SZS^{+}24a, XS22, YYLL20, ZLZ23b, Hof22a, IW22b, Agr23]. Confidently [TDB24]. Conflict [CBK^{+}22]. Conflicts [OPI22]. Confluent [MB23]. Conformal [CWZ21, HL24, YSWB24]. Confounded [SZS^{+}24a]. Confounders [LSP24]. Confounding [BK22, HBX23, MV20, MHOZ23, QMZ24, XC23a, ZWSZ21]. Conjugacy [AFDZ23]. Conjugate [BHW20, NS23]. Connection [LBN23]. Connectivity [MZW23, WCZZ24, ZSL20]. Connectome [GR21, HZLG22, ZCZZ23]. Conquer [GT23]. Consistent [SDH22, SSL22]. Constrained [BMDG20, CTC20, DHS23, JPR20, KCPW21, KPS23, XLWG23, ZSP20]. Constraining [SSB21]. Constraint [PLSZ23, QL24]. Constraints [DGW23, KL21b]. Consumed [LCD22]. Contextual [CLS21a]. Continuous [DLG23, HGC24, LDL^{+}22, WMK^{+}24]. Contraction [LSZ24]. Control [AE21, AL21, BMFR21, CFT21, CJZ23, DLXL23a, DLXL23b, DLXL23c, DGSZ23, Fer21, GSW23, GCW20, GRZL23, Jan23, KSR21, KSB23, KMPT21, LB23, Li20, LHS22, LYZ23, LCL23, LKLL22, TCYL23, XZWT21, XC23b, YRTJ23, ZC22, ZM23, CRYZ23]. Controlling [XZL23]. Controls [CSdG^{+}23, CWZ21, Fer21, MM21]. Convection [LYL22]. Convection-Diffusion [LYL22]. Convergence [BST24]. Convex [CFWY23, DW21, HHZ21, JGED21a, JGED21b, LV21, Rug21, Sha21, Wil21]. Convexity [DHS23, KPS23]. Convoluted [ZWZ24]. Coordinatewise [MHZ23]. Copula [DV24, MR20, SPMC21, SL22b, SL22d, YFZ20]. Copulas [ACADF23, NV24, RB20, Vat20]. Coronary [ARN20]. Coronavirus [TTL^{+}21]. Corrected [HQYZ20]. Correcting [JRR22]. Correction [Ano20a, Ano21a, Ano21b, Ano21c, Ano22a, Ano22b, Ano22c, Ano23a, Ano23b, FMM22, KCH24, XZLG24]. Correlated [HS21, LBN23, LLZL23, MN22, RR22]. Correlation [AM22, Cha21, DTQ23, GdM22, LC20, PND23, SPV20, SWZ20, XLWG23, ZZ22, ZYZ24]. Correlations [QZ22]. Corruption [SLC21]. Cortical [MYB^{+}20]. Cortisol [GLLZ22]. Cost [CS23]. Count [DFH^{+}21, JKL^{+}23, JZW23, KTDS20, LLM23b, MNBCCC24, Mur21]. Counterfactual [CLSG24, CWZ21, CT21a, FMM22, MM21]. Counterfactuals [BN21]. Coupled [DSW22b]. Couplings [JLS20a, JLS20b]. Covariance [Ano22c, CS23, DHHK21, DPY22, FGQ24, KML23, LLR21c, LW22a, MB23, MP23, PRV22, PWZ^{+}20, RT20a, RT20b, SLG23, SH20, WWZ22, XLLJ20, YLXL23, YLX24, ZSK23, ZCW23]. Covariance-Based [LLR21c]. Covariances [LXZZ23]. Covariate [GZM21, IS23a, LH22, LRY24, MQLH20, MLZH24, NCL22, PBOD23, WSM^{+}23, YS21, YSYZ23, ZC22]. Covariate-Adaptive [LH22, MQLH20]. Covariate-Assisted [IS23a]. Covariate-Informed [PBOD23]. Covariate-Specific [GZM21]. Covariates [CS24, DSGS22, DHL20, GO22, GZM21, Joc22, LXZZ23, LH22, MN22, PSBB21, QIL24, SL22b, ZL23, ZD24]. Covariation [MBH22a, MBH22b]. Coverage [Hof22a]. COVID [DM21, DY21, QDL21b, Jew21, QDL21a, TFCF21, WWLT24]. COVID-19 [DM21, DY21, QDL21b, Jew21, QDL21a, TFCF21, WWLT24]. Cox [DV24, HyLXZ21, KG23, LSN23, VDVM24]. CRC [All20, Beb20, Beh20, Cha23, Ché20b, Che20a, Cho22, Ell23, Har23, HM23, Hor22, LN20, Len20, McD20, Ni22, Ni23, Rah24, Sch20, Wan20, Zha20, Ber24, Muk24, Seo24, Sto22, Won24, Ye24, Zha22]. Crime [BDGJ23, PSBB21]. Criteria [ZSL^{+}23]. Criterion [HQYZ20]. Criticism [SCACvdH24]. Crop [PLL23]. Cross [BHT24, BPvdL20, Lei20, LDL^{+}22, RR22, SLG23, Wag20, ZZ23b]. Cross-Covariance [SLG23]. Cross-Fitted [ZZ23b]. Cross-Validated [BPvdL20, LDL^{+}22]. Cross-Validation [BHT24, Lei20, RR22, Wag20]. Crowdsourcing [XYWQ24]. Cui [ZP21]. Culture [YB20]. Cumulative [Jew21, LXZZ23, TLT^{+}24, ZZXL20]. Cun [LT22]. Cun-Hui [LT22]. Curvature [HBD24]. Curve [GZM21, JAO23, LRS20, dMMV21]. Curves [ASTZ24, BK20, CvS21, WLGC24]. Cutoff [CKTVB21].

D [Qiu24, SWZ20]. D-CCA [SWZ20]. Dai [CJZ23, Jan23, XC23b, ZM23]. Dan [LT22]. Dana [AS20]. Data [AL21, Agr23, Ano22b, ACV20, ABD^{+}21, BGOT20, BN21, Ban22, BWSB24, BTH21, BHW20, BIK20, CLX22, CCH23a, CSV21, CHKW24, Ché20b, CMZ20, CLZ20, CRYZ23, CCH23b, CX23, CLSG24, CBK^{+}22, CSX21, Coo20, CMC^{+}20, Cre20, DLXL23a, DL23a, DLP23, DP21, DM21, DY21, DHL20, DHHK21, DTQ23, DCGM23, DJM23, DL21, DKPW21, DH22, DGSZ23, DL23b, FGRS23, FBFV21, GT23, GO22, GLPD23, GKB^{+}21, GGR^{+}20, GR21, GYY^{+}22, GWL^{+}23, HMZZ23, HS21, HXGW22, Hor22, HL21b, HQCC21, HMW23, ISSM23, JJ21, JZW23, JVRS23, JHB20, KTDS20, KG23, KCPW21, KRW24, KLLL23, LNB22, LXLF20, LL22, LLZZ22, LZZ23a, LX24, LLR21c, LLR21a, LLR21b, LLM23b, LLgX22, LWSL22, LYL22, LWLL23, LZX^{+}24, LZS23, MST21, MST22, MCM20, MR20, MBCK22, MN22, MYB^{+}20, MNW^{+}20, MZW23, MP21, MNBCCC24, MKZ^{+}21, MJH20, OSDvdL24, OPI22, PLSZ23, PSBB21, PXZ21]. Data [Par22, PLZ24, QZ22, QDL21a, RR22, Rah24, RKJ^{+}20, SBD23, SKL22, SPMC21, SHWB23, SL22d, SR20, SJW23, STC^{+}20, SLSH20, SZZ24, Sug21, SMMH21, SZ21, TBC20, TLX24, TET^{+}23, TJK22, WWZ22, Wan23, WWLT24, WNK24, Wei20, XZWT21, XLYT20, XQ21, XZZ^{+}22, XSP23, XYY24, YPL^{+}21, YD20, YBRM20, Yao23, YNCK21, YYZZ20, YB20, YLSL20, YWAZ22, ZWS22, ZSK23, ZWLS20, ZYZ24, ZZH22, ZXS^{+}22a, ZLD23, ZKZ22, ZLGY23, dMMV21, QDL21b, Gio23, Qiu24]. Data-adaptive [DL21]. Data-driven [CRYZ23]. Data-Optimal [KG23]. Dataset [BLLY22, FDAZ22]. Datasets [KCH23a, KCH23b, Nie23, SF23, SWZ20, Yon23, ZSW22]. Datastreams [RZCL22]. David [Gio23, WFL22]. Davidian [Zha22]. Death [GKB^{+}21]. Debiased [DL23a, WHX20]. Debiasing [DJM23]. Decision [CLS21a, CLS21b, CT21a, GZW24, KT24, Kow22, LLM23a, QPL23, QCS^{+}21a, QCS^{+}21b, SZS^{+}24a, ZWSZ23]. Decomposition [DTQ23, GSS23a, HSZ24, SWZ20, WZLL22]. Decomposition-Based [SWZ20]. Decompositions [Ano22c, RT20a, RT20b, SH20]. Deconvolution [CFWY23, SPMC21]. Decorrelation [HCZY24]. Deductible [SL22b]. Deep [LTB22, RSC20, SZS24b, SSL22, TFCF21, ZMNVF22, ZJLH23]. Definite [CvS21, YDV24]. Definition [HS22]. Degeneration [JCY^{+}21]. Degree [WCNL24]. Degrees [CLS20]. Delaigle [Cha23]. Demand [BTH21, SA20, ZLQ^{+}21]. Dementia [WST^{+}20]. Demographic [KTDS20, TFCF21]. Dempster [Rug21, Sha21]. Dense [He24]. Densities [CKY22, HMP20, QDZ24]. Density [BF24, CJM20a, DL23b, MP24, SPMC21, TJC24, WC24, YSLW22, ZX24]. Density-Aided [WC24]. Dependence [CCH20, CHKW24, DGSZ23, LO21, LHS22, PWZ^{+}20, RB20, SLSH20, XZLG24]. Dependency [LX20]. Dependent [BHW20, CS24, CPH22, DV24, GSS22a, GT22, LRS20, RPLG22, SBD23, ZX24]. Depression [CPH22]. Depth [DLP23, LLgX22, MJH20, dMMV21]. Derandomizing [RWC23]. Derived [RLC^{+}21]. Descent [CLS21b, ZCW23]. Design [Cho21, DAGO22, HS22, LD24, LTY21, OWP20, SPMK22, Sto22, WW22, WSX22]. Design-Based [SPMK22]. Designed [GGR^{+}20]. Designs [CKTVB21, CFWY23, DLG23, EM23, HFR20, HXZ20, KSR21, PK20, RBMB21, ST23b, TZL20, YYLL20]. Detecting [GCW20, LJL23b, RRRF22]. Detection [AGS23, BSM23, CWS24, DG20, DREM^{+}21, LLB^{+}22, LLL22a, LLR21c, Mie23, SS22, SCACvdH24, TLT^{+}24, WJW^{+}21, WZL^{+}23, YS21, Yee22, YZL22, ZHW22, ZW23]. Development [HZLG22]. Developmental [HL21b, JJ21, LLR21a, LLR21b]. Deviation [HZQ^{+}21]. Deviations [MSC21, SHO20]. Diabetes [GWL^{+}23]. Diagonalizability [XDM24]. Dichotomous [JCY^{+}21]. Difference [FDNZ24]. Differences [GVW20, RBMB21, YZL22]. differentiable [GSS22b]. Differential [AS21, DL22b, Dre23, OWP20, THXZ23]. Differentially [BWSB24]. Diffusion [LYL22, PLCS21, QW21]. Diggle [LN20]. Digital [XCZW23]. Dimension [CXH23, CWF^{+}24, McD20, WC24, WWF23, XS21, ZMZ24]. Dimension-Free [WC24]. Dimension-Independent [CWF^{+}24]. Dimensional [AM24, BSM23, BCG20, CLX22, CXH23, CGM23, CZG24, CTC20, CYZ22a, CF23, CWS24, CKS23, DSL22, DJM23, DPY22, DGW23, EB21, FMW20, FZHL23, FDNZ24, GT22, GZM21, GLLZ22, GC22, HMZZ23, HS21, HXZ20, HHLT21, HBX23, JPR20, JWW^{+}24, KSW20, KS24, KAZZ20, LR23a, LS20, LWWftADNI21, LCL22, Li22, LXZZ23, LZCL24, LXJ22, LLM23b, LT22, LLZL23, LZX^{+}24, LZH24, Loh20, LC20, MCL21a, MKftADNI24, MN22, MD20, NCL22, OY22, Peñ22a, QIL24, RLC^{+}21, RS22, SS22, Sen24, SSLL21, SWZ20, SR20, SZ21, TF23b, WHX20, WPB^{+}20a, WZLL22, WCNL24, XLYT20, XCZW23, XYY24, YLXL23, YLX24, ZWS22, ZWKZ22, ZL23, ZPYZ24, ZZL24, ZZ23b, ZLZ23b, ZSP20, ZKZ22, ZLGY23, CWW22, CX23, LPR24, WPB^{+}20b]. Dimensions [FGQ24, HR24, HY24, LYZL24, MM21, MR23, TSW23, YB21, ZGZC20, ZWZ24]. Ding [Rah24]. Ding-Geng [Rah24]. Dinov [Qiu24]. Direct [LSRR21, QLFL20, RSvdL21]. Directed [LSP20, ZHW22]. Directional [LWY^{+}21]. Disaggregated [AL21]. Disclosure [JRSW22]. Discontinuity [CKTVB21, DLG23, RBMB21]. Discoveries [ZROD23]. Discovering [LSD21]. Discovery [CJZ23, DLXL23a, DLXL23b, DLXL23c, DGSZ23, GSW23, Jan23, KSB23, KLSZ21, LB23, LYZ23, LCL23, LSP24, XC23b, XZL23, ZC22, ZM23, Ano21d]. Discrepancy [ACADF23, KLLL23]. Discrete [Agr23, BKR^{+}23, CFVMW20, Ize21, LPR24, TLX24, YFZ20, Zan20]. Discrete-Time [BKR^{+}23]. Discriminant [XYY24]. Discussion [AL23, Ban22, CS20, CHDP23, CJZ23, Cox20, DM21, Dav20, DY21, DW21, Efr22, FHT20, Gho22, Hof22a, HL21b, Jan23, JJ21, KCH23a, LB23, LS20, LLL21, LZ21b, Nie23, Pen22b, SF23, SH20, VD21, Wil21, XC23b, gXZ20, XS22, Yon23, ZP21, ZK23]. Disease [SWR24, SZZ24, WWLT24, YWKZ22]. Diseases [PK23, TTPD20]. Disorders [GZW24]. Disparities [LWLL23]. Dispersion [PSVV24]. Distance [GBL^{+}21, NDH^{+}24, SPV20, WCL21, WSX22, Wei20]. Distance-Based [Wei20]. Distances [LOW21, WPTM24]. Distancing [TFCF21]. Distortion [XZWT21]. Distributed [CLLY22, CLZ22, GT23, HS21, HR24, MKW^{+}24, YWAZ22, ZLZ24]. Distribution [BK20, DS23, DWR23, HHH23, HL24, LCD22, LLM23a, LLgX22, MAZB22, QW24, SDH22, WPTM24]. Distribution-Free [DS23, DWR23, HHH23, MAZB22, SDH22]. Distributional [DFG^{+}24, HKZ23, YBRM20]. Distributionally [MQL21a, MQL21b]. Distributions [AM24, BHW20, DL22a, GPPV20, SW23, SL22b, SHWB23, TET^{+}23]. Districts [RBMB21]. Divergence [ZZXL20]. Diverging [FFHL22, GT22, ZZLC20]. Diverse [LLgX22]. Diversified [FL22]. Divide [GT23]. Divide-and-Conquer [GT23]. Divisible [BSH21]. DNA [GLLZ22]. Do [BHT24, FMM22, RBMB21]. Does [BHT24]. Domain [CZW21, HVL23, LSZ24]. Domains [PBF22]. Donner [Yee22]. Dosing [SM21]. Dot [XX23]. Double [NMJ20]. Doubly [CLW20, DKJ24, MZPD22, PZ21a, SM21, YYZZ20]. Downstream [Wan23]. Drift [PLCS21]. Drift-Diffusion [PLCS21]. Driven [GLPD23, CRYZ23]. Drton [All20]. Dual [Rob24]. Duong [Wan20]. Duplicate [AGS23]. D’Urso [Che20a]. Dylan [Won24]. Dynamic [Ano23b, BAXV22, CSX21, DGD21, DM22, FGY24, HY24, KSW20, LYC23, LRZ24, LLK^{+}20, OPI22, SWL^{+}23, SMN^{+}20, SW21b, XZZ^{+}22, ZLQ^{+}21, ZSZL23, ZZQ24, Zha22]. Dynamically [GT22]. Dynamics [DM21, DY21, QDL21a, QDL21b, SZS24b, YZ22].

e-Book [WFL22, Cho22, Ni22]. Earthquake [ILMY22]. Eastern [JSK^{+}20]. EBLUPs [LW22b]. Ecological [Chi21]. ed [Len20, Sto22]. ed. [Par23, Qiu24]. Edge [WBK^{+}22]. Editorial [Ano20b, Ste23]. Editors [Ano24]. Edlefsen [Rug21, Sha21]. eds [All20, Cha23, Ni23, Won24]. EEG [CPH22, MLH^{+}22]. EEG-Based [MLH^{+}22]. Effect [CFVMW20, Fog20, GZM21, HBX23, LY22, LSRR21, WHX20, Yee22, ZSL20, ZD24, AS20]. Effectiveness [KSR21]. Effects [ARN20, BCG20, CKTVB21, DWHS23, EB21, FVX20, FLKK21, FAM21, GSS23b, GSS23a, GGR^{+}20, GYY^{+}22, HL21a, LR23a, LSD21, Li20, LCL22, LXZZ23, LSB^{+}22, MN23, MHOZ23, MKW^{+}24, NLS^{+}21, PK23, RKJ^{+}20, RSvdL21, SST21, SLS20, SL22a, SWL^{+}23, SYD^{+}22, SRS^{+}23, TFS21, TTPD20, TTL^{+}21, VB23, YPL^{+}21, YD20, YSWB24]. Efficacy [HS22]. Efficiency [KPS23, WZC23, XC23a]. Efficient [FMW20, FGW23, HHH23, Kal21a, LSX23, LS20, LSRR21, LRY24, Loh20, QW24, SMO23, SLL^{+}24, WW22, WPB^{+}20a, WPB^{+}20b, XX23]. Efficiently [KTDS20, SQPQ21]. Efron [CS20, Cox20, FHT20, gXZ20]. EHR [KG23]. Eigen [HTF23]. Eigenspace [CLLY22]. Eigenvalue [KML23]. Eigenvalues [FGZ22, GT22, KML23]. Eigenvector [MSC21]. Eigenvectors [FFHL22]. Elastic [AL23, CHDP23, ZK23, ZWX^{+}23a, ZWX^{+}23b]. Electricity [BTH21]. Electronic [GWL^{+}23, NMJ20, PLSZ23, SMMH21, WZW20]. Elizabeth [Che20a, Won24]. Elucidating [CPH22]. Emanuele [LN20]. Embedded [ILMY22]. Embedding [DSW22a, FLLZ22, HL21b, JJ21, LLR21a, LLR21b, YQ23, ZHW22, ZW23]. Emergency [FLKK21]. Emigration [MA23]. Emil [HM23]. Emissivity [DREM^{+}21]. Empirical [BTHK20, Cre22, DH22, Efr22, Gho22, Hof22a, IW22a, IW22b, ISSM23, Imb22, Pen22b, RLS23, XS22, YDV24, YB24, Zha20]. Emulation [NCMK20]. Encouragements [FLKK21]. Encyclopedia [PLZ24]. Endogeneity [CT21b, Han21]. Endogenous [FVX20, SL22b]. Energy [NCMK20]. Engagement [BMDG20]. Engineering [Cre20]. Enhance [WHC^{+}24]. Enhanced [MZPD22, YLXL23]. Enlarged [CLLY22]. Enough [ZDY23]. Ensemble [KSW20, LDL^{+}22]. Ensembles [HLS^{+}21, HH23, LLM23a, TF23a]. Entry [HZQ^{+}21]. Entry-Wise [HZQ^{+}21]. Environment [LDL^{+}22]. Environmental [Chi21]. Epidemic [DM21, DY21, QDL21a, QDL21b]. Epidemics [BAXV22]. Epidemiology [SPMC21, Mor21]. Epidural [ZHC21]. Epigenetic [LSB^{+}22]. Epigenomic [XTK^{+}22]. Episodically [LCD22]. Equation [OWP20]. Equations [CDJ23, DL22b, NV24, THXZ23]. Equilibrium [ZLQ^{+}21]. Equivariant [SW21a]. Eric [Zha22]. Error [Ano22c, Cha23, FSV20, GRZL23, LCD22, LHS22, LW22b, RT20a, RT20b, SH20, WTW20, WQ23, XZWT21, YRTJ23, CRYZ23]. Errors [ISSM23]. Estimability [CLZ20]. Estimate [BHT24, MKW^{+}24, YD20]. Estimated [FL22, Li20, LWSL22]. Estimates [JRR22, SWR24, TDB24, WJIS20]. Estimating [DKJ24, DHHK21, DPY22, FGZ22, HS22, HSC21, IS23b, LNB22, LLK^{+}20, MSC21, MP24, MN22, NV24, NLS^{+}21, QLFL20, SMN^{+}20, SZZ24, SW21b, TTPD20, XZZ^{+}22, ZM22, ZD24, ZYZ24, ZZQ24]. Estimation [ACADF23, Ano22c, AZ23, BMS20, BPvdL20, CLX23, CFMB24, CS20, CCH20, CSV21, CKY22, CLLY22, CLZ22, DGH23, DHL20, DREM^{+}21, DL23b, Efr20a, FGW23, FGQ24, FVX20, FSV20, FAM21, FHT20, GZM21, HLL22, HVL23, HBX23, ISSM23, JKLW23, KML23, LR23a, LSX23, LCD22, LLM23a, LSN23, LZCL24, LKM21, LW22a, LSRR21, MP23, MBCK22, MNW^{+}20, MNHG24, MCLT24, Mie23, PRV22, Pai23, PZ21a, QDZ24, QL24, RT20a, RT20b, RSvdL21, SS22, SMO23, SM21, SH20, SSLL21, SZS^{+}24a, SLSH20, TJC24, Wag20, WCL21, WWZ22, WNK24, WBBM23, XC23a, gXZ20, XS21, XX23, YPL^{+}21, YFZ20, YZZ23, YDV24, Yao23, YWW^{+}20, ZMZ24, ZSK23, ZX24, ZWLS20, ZZ22, ZB22, ZM22, ZZL24, ZW20, ZCW23, ZKZ22, SLZ24]. Estimator [AL21, BF24, CLZ22, Fer21, GB23]. Estimators [CJM20a, CLS20, SPMK22, YWAZ22, dHZ24]. Ethics [Gio23]. EUR45.00 [Par22]. Evaluating [ARN20, HLS^{+}21, LHW23]. Evaluation [GLLT21, IL23, MBH22b, NMSD21, RLY^{+}23, SWL^{+}23]. Evaluations [BWSB24]. Evarist [Sen24]. Event [LHW23, MR20, SLSH20, WRG23, WNK24, ZZH22, ZWSZ23]. Events [KG23, SYD^{+}22, SMMH21, TMNM22]. Evidence [GWL^{+}23, ILMY22, KSR21]. Evolution [HTA^{+}22]. Exact [CWZ21, FLLZ22]. Exceedances [BGOT20]. Exchange [HFR20]. Exchangeability [HLS^{+}21]. Exchangeable [CKS23, DOH22, GLM23]. Expectation [HJO^{+}20, XS21]. Expected [CMW24, KY20, SQPQ21]. Experimental [IL23, WW22]. Experiments [BRS22, BP23, DAGO22, GSS23b, GGR^{+}20, GCW20, HFR20, IJM21, LD24, LH22, LRY24, SHR^{+}20a, SHR^{+}20b, SWPH20, WSX22, WWL23, WD21, XWMD24, YQL23, ZD24, Cho21]. Exploit [FMM22]. Exponential [BHW20, Che20d, HL21b, JJ21, LLR21a, LLR21b]. Exponential-Family [HL21b, JJ21, LLR21a, LLR21b]. Exposure [LSD21, LWLL23]. Exposures [MKW^{+}24, Wes22, WMK^{+}24]. Expression [LLR21c, SZZ24, WJIS20, WIS20]. Extended [JCY^{+}21, LXLF20].

Extensions [AFDZ23]. External [CCLS21]. Extracting [ZPZW23]. Extrapolating [CKTVB21]. Extrapolation [KMPT21]. Extreme [BSH21, HW21, MN23, dHZ21, dHZ24]. Extremely [HXC23, TZ24]. Extremes [CCH20, HR24, TET^{+}23, WJW^{+}21, ZSW22]. Extremile [DGS22]. Exuberance [JRR22].

Face [JSK^{+}20]. Factor [AZ23, BN21, CTC20, CLZ20, CYZ22a, CF23, DDX22, FMM22, FLY24, FD21, GT22, GT23, He24, LL22, LT22, LLZL23, MC22, OY22, Peñ22a, XYWQ24, ZPYZ24]. Factorial [BP23, HS22, LRY24]. Factorization [JM20, LZQ^{+}21]. Factors [DL22a, FGZ22, FL22, GT22, GDP22, KSR21, XWMD24]. Failure [Coo20]. Fair [VB24]. Fairness [CLSG24, FWW23]. Fairness-Oriented [FWW23]. False [CJZ23, DLXL23a, DLXL23b, DLXL23c, DGSZ23, GSW23, Jan23, KSB23, LB23, LYZ23, LCL23, XC23b, XZL23, ZC22, ZM23]. Families [Che20d]. Family [BHW20, HL21b, JJ21, LLR21a, LLR21b, MLZH24]. Fast [CLX23, FWZ^{+}21, HJO^{+}20, Kow22, MOA24, SWS22, WZL^{+}23, YB24]. FDR [GCW20, LKLL22, TCYL23]. Feasibility [BWSB24]. Feature [GRZL23, LX24, LKLL22, LC20, NCL22, NMJ20, TCYL23, YB21, ZQL^{+}23]. Features [LSS23, LKLL22, MD20]. Female [LNB22]. Field [LHGY23]. Fields [CS23, Ize21, KS24, YZL22]. Filtering [KSB23, Par23]. Filters [IAPK23, JLS20a, JLS20b]. Financial [AB20]. Finding [LBN23]. Fine [GJK^{+}20, TTPD20, SMO23]. Fine-Scale [GJK^{+}20]. Fingerprinting [WJW^{+}21]. Finite [CHYL23, DGH23, LPR24]. Finite-dimensional [LPR24]. Finite-Sample [CHYL23]. First [CLZ22, CDJ23]. First-Order [CLZ22]. Fisher [MKV^{+}24, YLX24]. Fit [HZQ^{+}21, SBY24, XLWG23, ZNBR22, ZDY23]. Fitted [ZZ23b]. Five [CMZ20, ILMY22]. Fixed [Ano22c, BCG20, LSZ24, RT20a, RT20b, SH20]. Fixed- [SH20, Ano22c, RT20a, RT20b]. Fixed-Domain [LSZ24]. FL [All20, Beb20, Beh20, Ber24, Cha23, Ché20b, Che20a, Cho22, Ell23, HM23, Hor22, LN20, Len20, McD20, Muk24, Ni22, Ni23, Rah24, Sch20, Seo24, Sto22, Wan20, Won24, Ye24, Zha20, Zha22]. Flexible [FDF20, JZW23, ZSW22]. Florida [Har23]. Flow [QW24]. Fluctuations [RRRF22]. fMRI [MYB^{+}20]. Focuses [MN23]. Foods [LCD22]. Forecasting [BTH21, FLSU20, MANW20, YNCK21]. Forecasts [HHLT21]. Forest [CBK^{+}22, GLLT21, KCPW21]. Forests [DGH23, SBD23]. Formula [LTB22]. Forward [ZZXL20]. Foundations [Sen24, Hor22]. Fractional [SKL22]. Fragmented [DHHK21, MBCK22]. Frailty [GKB^{+}21]. Frailty-Based [GKB^{+}21]. Frames [RKJ^{+}20]. Framework [CCLS21, HSC21, MBCK22, SLG23, SWL^{+}23, SLZ24, TF23a, WGSC23, XCS20]. Francis [Har23]. Fréchet [TWM23]. Fredholm [CDJ23]. Free [CJZ23, DLXL23c, DS23, DWR23, FMW20, HHH23, Jan23, LB23, LS20, LYZ23, LKLL22, Loh20, MMR24, MAZB22, SDH22, TCYL23, WPB^{+}20a, WC24, XC23b, ZM23, ZZXL20, DLXL23b, WPB^{+}20b, ZLZ23a]. Freedom [CLS20]. Freemium [BMDG20]. Frequencies [MP24]. Frequency [BTHK20, CMZ20, KLLL23]. Frequentist [Har23, Rob24]. Friendly [Par22]. Frontal [CPH22]. Full [BHW20]. Full-Conditional [BHW20]. Fully [HHH23]. Function [DL23b, FGQ24, KY20, LYC23, SJW23, WWZ22, XLWG23, YBRM20, ZWKZ22]. Function-based [KY20]. Function-on-Scalar [LYC23, ZWKZ22]. Functional [Ano23a, CCH23a, CCH23b, CPH22, CMC^{+}20, DHHK21, FGQ24, GYY^{+}22, HMP20, HSZ24, HyLXZ21, JCYS20, JCY^{+}21, JAO23, KCPW21, KPV22, LJL^{+}23a, LLLZ23, LLZZ22, LJL23b, LZCH21, LHGY23, LW22a, LLM23b, LRZ24, LKPG21, MBCK22, MZW23, Mie23, PS23, PLL23, SBJ20, SL22d, SWPH20, THP^{+}23, WWZ22, WCZZ24, XYY24, Yao23, ZCM22]. Functional-Coefficient [LRZ24]. Functionals [GC22, HW21, ZZL24]. Functions [CS23, CFVMW20, FBFV21, KHK21, LA20, MB23, SLG23]. Fundamentals [Ye24]. Fused [ZLZ24]. Fused-Lasso [ZLZ24]. Fusion [BMSL22, LLgX22, SLgX20]. Future [JAO23, Mar21, SSB21, TMNM22].

G [Ni23, TFS21]. Gaining [SWS22]. Galaxy [DREM^{+}21]. Games [BMDG20]. Gamma [JZW23]. Gang [PSBB21]. Gang-Related [PSBB21]. GAP [XCS20]. Gary [Ni22]. Gaussian [CS23, DKPW21, JKL^{+}23, JT22, KS24, LCL23, LSZ24, MLL21, MML23, PBF22, SKL22, SL22d, SHR^{+}20b, WGK24, XX20, XZL23, ZL23]. Gaussianization [MHZ23]. GEE [HMW23]. GEE-Assisted [HMW23]. Gene [GCW20, JCY^{+}21, LLR21c, RLYI20, SZZ24, WJIS20, YLXL23]. Gene-Based [JCY^{+}21]. Gene-Set [YLXL23]. General [CJK^{+}24, DG20, DGSZ23, FLKK21, HXC23, MYB^{+}20, SLZ24, WGSC23, XCS20, ZW23]. Generalisability [KCH23a, KCH23b, Nie23, SF23, Yon23]. Generalization [DDX22]. Generalized [BMA^{+}22, CLX23, CGM23, CJZ23, DLXL23b, DLXL23c, FDAZ22, GB23, Jan23, JCY^{+}21, LB23, LYZ23, LZZ23a, LZCL24, LLZL23, Pai23, PSVV24, SL20, SHR^{+}20b, TF23b, WMK^{+}24, XC23b, YYZ24, YWW^{+}20, ZM23, ZDNH23, Cho21, AM22]. Generated [MKZ^{+}21]. Generating [CSdG^{+}23]. Generation [DSW22b, WJM22]. Generative [ZJLH23]. Generic [CFVMW20, PWZ^{+}20]. Genes [LWSL22, YSL24]. Genetic [CFMB24, SL20, YWKZ22, ZZ22, ZCZZ23, ZYZ24]. Genetic-Imaging-Clinical [YWKZ22]. Geng [Rah24]. Genome [CCLS21, LLL22a, LSB^{+}22, LLL22b, SL20, ZZ22]. Genome-Wide [CCLS21, SL20, ZZ22, LSB^{+}22]. Genomic [HSB21]. Genomics [MKV^{+}24, MAZB22]. Geoadditive [YWW^{+}20]. Geographic [PBOD23, RBMB21]. Geometric [SBK20]. Geostatistical [PBF22]. Geostatistics [LN20]. Gibbs [DW21, HHZ21, JGED21b, LV21, Rug21, Sha21, Wil21, JGED21a, NS23]. Giné [Sen24]. Gini [HW21]. Giorgi [LN20]. Glioblastoma [CMC^{+}20]. Global [KLLL23, LN20, LLL22b, MCL21a, TWM23]. Goal [HL21a]. Gold [CJM^{+}20b, HS22]. Gong [Rug21, Sha21]. Good [ZDY23, Pai23]. Goodness [HZQ^{+}21, SBY24, XLWG23, ZDY23]. Goodness-of-Fit [XLWG23, ZDY23]. Google [GJK^{+}20]. Government [Dre23]. Grace [Cha23]. Gradient [CXH23, CLS21b, GSS22b, NF21, NS23, ZCW23]. Gradient-Based [GSS22b]. Graph [LSP20, LOW21, SPV20, ZW23, ZLQ^{+}21]. Graph-Based [ZLQ^{+}21]. Graphic [Par22]. Graphical [All20, FDLL20, Ize21, LJL^{+}23a, LLLZ23, LCL23, LWSL22, MLL21, MCM20, MML23, MP21, SL22d, TLX24, TJH^{+}23, WBK^{+}22, ZL23]. Graphs [KL21a, XX23, YS21]. Gray [SMO23]. Great [Ste23]. Greedy [CCH23b, HMZZ23]. Group [BMA^{+}22, CCgX23, DHL20, Har23, LTY21, LC20]. Grouped [ACV20, BMA^{+}22, Sug21]. Groups [DGD21, KSR21]. Guarantee [GCW20]. Guaranteed [HSZ24]. Guided [HTF23]. Gustafson [Cha23]. Gut [TJH^{+}23]. GWAS [XSP23].

H [All20, AS20, Beb20, Beh20, Ber24, Bor24, Cha23, Ché20b, Che20a, Cho22, Ell23, Gio23, Ism20, LN20, Len20, McD20, Muk24, Ni22, Ni23, Qiu24, Rah24, Seo24, Sto22, Wan20, Won24, Ye24, Zha20, Zha22]. Haikady [Ism20]. Hall [All20, Beb20, Beh20, Cha23, Ché20b, Che20a, Cho22, Ell23, LN20, Len20, McD20, Ni22, Ni23, Rah24, Sch20, Wan20, Zha20, Ber24, Muk24, Seo24, Sto22, Won24, Ye24, Zha22]. Hall/CRC [All20, Beb20, Beh20, Cha23, Ché20b, Che20a, Cho22, Ell23, LN20, Len20, McD20, Ni22, Ni23, Rah24, Sch20, Wan20, Zha20, Ber24, Muk24, Seo24, Sto22, Won24, Ye24, Zha22]. Hamiltonian [ST23a]. Hamiltonian-Assisted [ST23a]. Handbook [Fra20, All20, Cha23, Che20c, Chi21, Mor21, Ni23, Won24]. Handling [SKL22]. Harezlak [WFL22]. Harvard [Par22]. Hastings [KR23, WWW22]. Hauck [Yee22]. Hauck-Donner [Yee22]. Hazards [DV24, HMMMN21, HBX23]. Health [DGD21, GLLT21, GWL^{+}23, HQCC21, LN20, LKM21, LLK^{+}20, NMJ20, PLSZ23, Rah24, SMMH21, WZW20, NMSD21, Zha20]. Heavy [BCT24, TJC24]. Heavy-Tailed [BCT24, TJC24]. Hellinger [GdM22]. Hens [Yon23, SF23]. Hermitian [CvS21]. Heterogeneity [AB20, LSN23, RLYI20, WBK^{+}22, WCZZ24, WCNL24, ZLZ24]. Heterogeneous [ARN20, CLX22, FVX20, FLKK21, LSD21, LJL23b, LLR21c, MA23, MKW^{+}24, QDZ24, Sug21, XLYT20, XTK^{+}22]. Heteroscedasticity [FGJS22, Joc22, ZZ23b]. Heteroscedasticity-Adjusted [FGJS22]. Heteroscedasticity-Robust [Joc22]. Hidden [AM24, WST^{+}20]. Hierarchical [ACV20, BGOT20, BTH21, BSH21, BHW20, CMW24, DGD21, DOH22, DWR23, FLLZ22, GT23, GBW24, KSW20, LW20, LLB^{+}22, WWLT24, ZSW22, ZHC21]. High [AM24, BSM23, BCG20, CLX22, CXH23, CGM23, CZG24, CTC20, CMZ20, CYZ22a, CWW22, CX23, CF23, CWS24, CKS23, DSL22, DJM23, DPY22, DGW23, Dua23, EB21, FMW20, FGQ24, FZHL23, FDNZ24, GT22, GZM21, GLLZ22, HMZZ23, HS21, HR24, HHLT21, HBX23, HY24, JPR20, JWW^{+}24, KSW20, KS24, KAZZ20, KLLL23, LR23a, LS20, LCL22, Li22, LXZZ23, LZCL24, LXJ22, LLM23b, LT22, LLZL23, LYZL24, LZX^{+}24, LZH24, Loh20, LC20, MCL21a, MKftADNI24, MM21, MN22, MR23, MD20, OY22, Peñ22a, QIL24, RLC^{+}21, RS22, SS22, SSLL21, SWZ20, SR20, SZ21, TSW23, TF23b, WHX20, WPB^{+}20a, WPB^{+}20b, WZLL22, WCNL24, WWF23, XLYT20, XYY24, YB21, YLXL23, YLX24, YQ23, ZGZC20, ZWS22, ZWKZ22, ZL23, ZPYZ24, ZZL24, ZZ23b, ZLZ23b, ZWZ24, ZSP20, ZKZ22, ZLGY23]. High-Accuracy [Dua23]. High-Dimensional [AM24, BCG20, CLX22, CGM23, CZG24, CTC20, CYZ22a, CF23, CWS24, CKS23, DSL22, DJM23, DPY22, DGW23, EB21, FMW20, FZHL23, FDNZ24, GT22, GZM21, GLLZ22, HMZZ23, HS21, HHLT21, HBX23, JPR20, KSW20, KS24, KAZZ20, LR23a, LS20, LCL22, Li22, LXZZ23, LZCL24, LXJ22, LLM23b, LT22, LZX^{+}24, LZH24, Loh20, LC20, MCL21a, MKftADNI24, MN22, MD20, OY22, Peñ22a, QIL24, RLC^{+}21, RS22, SS22, SSLL21, SWZ20, SR20, SZ21, TF23b, WHX20, WPB^{+}20a, WZLL22, WCNL24, XYY24, YLXL23, YLX24, ZWS22, ZWKZ22, ZL23, ZPYZ24, ZZL24, ZZ23b, ZLZ23b, ZSP20, ZKZ22, CWW22, CX23, WPB^{+}20b]. High-Frequency [KLLL23]. High-Order [YQ23]. Higher [SBY24, SCACvdH24]. Higher-Order [SBY24]. Highly [PBF22]. Hilbert [THP^{+}23]. Hilbertian [CJK^{+}24]. History [KTDS20, Par22]. Hoboken [Ism20]. Holloway [Zha22]. Homophilous [IS23b]. Honest [WWF23, ZLZ23b]. Horizon [ZZQ24]. Horizons [SLL^{+}24]. Horseshoe [SBJ20]. House [RBMB21]. Housing [Ban22, TJK22, YLPC20, ZXS^{+}22a]. Howard [Par22]. Hub [HZB^{+}24]. Huber [SZF20, TSW23]. Hui [LT22]. Hvitfeldt [HM23]. Hybrid [LZZ23b]. Hypergeometric [MB23]. Hypergraph [ZW23]. Hyperspherical [GPPV20]. Hypotheses [DSL22, KRW24, LHS22, WD21]. Hypothesis [Li22, LLYM21, LDL^{+}22, LSB^{+}22, LLL22b, LZH24, MCL21a, PND23].

ICeD [WJIS20]. ICeD-T [WJIS20]. Ideas [GV21]. Identifiability [CLZ20, CHYL23]. Identification [FAM21, LTO^{+}21, LWWftADNI21, RSvdL21, WBBM23, XC23a]. Identified [FBZ21]. Identifying [MHOZ23, YSL24, RBMB21]. Idiosyncratic [FMM22]. IFAA [LTO^{+}21]. Ignatiadis [Efr22, Gho22, Hof22a, Imb22]. Ignorability [RMM20]. II [Ano21d, GWL^{+}23]. Illness [GKB^{+}21]. Image [FLSZ20, LWWftADNI21]. Image-on-Scalar [LWWftADNI21]. Images [MKftADNI24, ZBM^{+}24]. Imaging [CHKW24, Ché20b, DTQ23, FBZ21, LA20, LZX^{+}24, TBQ20, YWKZ22]. Imbalanced [ZLD23]. Immune [WJIS20]. Impact [GLLT21, Kaf20, TJH^{+}23]. impacts [NMSD21]. Imperfect [GSS23b, LD20, YRTJ23]. Implications [CLZ20, FDF20, KC22]. Implicit [DAGO22, FYY23]. Importance [WGSC23]. Important [FDNZ24, GV21]. Improve [WSM^{+}23]. Improved [BPvdL20, HVL23, PZ21a, RLYI20]. Improvement [CMW24]. Improves [Pai23]. Improving [MN23]. Impure [WIS20]. Imputation [CPH22, JRSW22, MJH20, SKL22, YLSL20]. Inchausti [Rob24]. Incidence [HBP^{+}20, Jew21]. Income [SHWB23]. Incomplete [DM21, DY21, HTA^{+}22, QDL21a, QDL21b, RKJ^{+}20, ZHC21]. Incorporating [CCLS21, GBL^{+}21]. Independence [FGRS23, GHL24, HGC24, KY20, LZ21a, MZW23, SDH22, TMZ22]. Independent [BGOT20, CWF^{+}24, GS23, MNW^{+}20]. Index [FYY23, HKZ23, KPS23, LZH24, YLPC20, ZWZ22]. Indexed [KHK21]. India [JSK^{+}20]. Indices [dHZ21]. Indirect [Hof22b]. Individual [BLLY22, CLX22, FVX20, YSWB24, ZZH22]. Individual-Level [ZZH22]. Individualized [Ano21d, CCgX23, CZW21, CT21a, DLS24, FWW23, FDNZ24, GZM21, Han21, HSC21, IL23, KLSZ21, LM21, MQL21a, MQL21b, PZ21a, QLFL20, QPL23, QMZ24, QCS^{+}21a, QCS^{+}21b, SLgX20, TBQ20, TXQ21, WZW20, ZWSZ21, ZWSZ23]. Induced [XC23a, YSL24]. Infection [KL21a, TJH^{+}23]. Infectious [PK23]. Inference [AS22, Ano22b, AM21, BRS22, BCT24, BST24, BAXV22, CGM23, CLW20, CLS21a, CLS21b, CLZ22, CWW22, CF23, Che23, CWS24, CFVMW20, CWZ21, CKS23, CBK^{+}22, CPS^{+}24, DP21, DFG^{+}24, DHS23, FLSU20, FDLL20, FZHL23, FD21, FNDK23, Fry24, GBW24, GSS22b, GMVFX22, GH21, HyLXZ21, HZB^{+}24, HR24, HW21, HGC24, IJM21, Joc22, KC22, KKS23, KCH23a, KCH23b, LR23a, LSD21, LO21, Li20, LTO^{+}21, LM21, LL22, LCL22, LZCL24, LXJ22, LLPY24, MW20, MQLH20, MST21, MST22, MM21, MKZ^{+}21, Nie23, OSDvdL24, PT23, PSVV24, QZ22, RB20, RLY^{+}23, RLS23, RMM20, SBK20, SF23, SST21, SSLL21, SYD^{+}22, STC^{+}20, SZ21, TBC20, WHX20, WSM^{+}23, WPTM24, WLGC24, WGSC23, WWF23, WZC23, XCS20, XLLJ20, XSP23, YZ22, Ye24, Yon23, YWW^{+}20, YB24, ZDM21, ZWS22, ZKR23, ZSP20, ZJLD23, NMSD21, Beh20, Ché20b, Muk24, Won24]. Inferences [CHKW24, GC22, LLgX22, MLH^{+}22]. Inferential [CSV21, JRSW22]. Inferring [HTA^{+}22, IS23b, KTDS20]. Infinite [LM21, Sen24, SLL^{+}24, ZZQ24]. Infinite- [LM21]. Infinite-Dimensional [Sen24]. Infinitely [BSH21]. Inflated [JZW23, SPMC21, ZPZW23]. Influence [HLS^{+}21, IS23b]. Information [BMS20, CCLS21, FMM22, Hof22b, HQYZ20, NCL22, XCS20, XS21]. Informative [LHW23, RB20, SMMH21]. Informed [PBOD23, Zan20]. Innovative [Beb20]. Input [SWPH20]. Insights [Mar21]. Instrumental [ARN20, BP23, CT21b, FLKK21, MCLT24, QCS^{+}21a, QCS^{+}21b, VB23]. Instruments [KSR21, XSP23]. Insurance [GLLT21, SL22b, SZS24b]. Integrated [Ché20b, HS21]. Integrating [PXZ21, XQ21]. Integration [DDX22, JWW^{+}24, LZZ23a, ZDNH23]. Integrative [CLX22, LL22, XLLJ20]. Integrity [JRSW22]. Intelligence [CT21a]. Intent [PK23]. Intent-to-Treat [PK23]. Intentional [XZWT21]. Inter [MLL21]. Inter-Subject [MLL21]. Interacting [IAPK23]. Interaction [DOH22, LYC23, PBOD23, TJH^{+}23]. Interactions [FD21, LDL^{+}22, PBOD23, RKJ^{+}20]. Interest [MN23]. Interface [MLH^{+}22]. Interference [FAM21, HXZ20, IJM21]. Internal [FBZ21]. International [OPI22]. Internationally [PK20]. Internet [YNCK21]. Interpatient [WBK^{+}22]. Interplay [TFCF21]. Interpolating [SHWB23]. Interpolation [KMPT21]. Interpretable [FLSU20, MB23]. Interquantile [CS24]. Interruption [PSBB21]. Interstellar [HMB^{+}20]. Interval [SZS^{+}24a, ZWLS20, ZQL^{+}23]. Interval-Censored [ZWLS20]. Interval-Valued [ZQL^{+}23]. Intervals [Agr23, CFT21, Efr22, Gho22, GRLB20, Hof22a, IW22a, IW22b, Imb22, KL21b, Pen22b, XS22, YYLL20]. Interventions [CP23, TTL^{+}21]. Intraday [ASTZ24]. Intrinsic [CvS21]. Introduction [AE21, Ano21d, Cho22, KLSZ21, Sch20]. Invalid [LO21, XSP23]. Invariance [NCMK20]. Invariance-Preserving [NCMK20]. Invariant [ZWSZ23]. Inventory [KCPW21]. Inverse [DG23, GZW24, MW20]. Investigating [HZLG22, PSBB21]. Invitation [DHJW22]. Involving [SZ21]. IPAD [FLSU20]. iProMix} [SJW23]. Irrational [JRR22]. Irregular [LRZ24]. Irregularly [PLSZ23]. ISBN [All20, AS20, Beb20, Beh20, Ber24, Bor24, Cha23, Ché20b, Che20d, Che20a, Cho22, Ell23, Gio23, Har23, HM23, Hor22, Ism20, LN20, Len20, McD20, Muk24, Ni22, Ni23, Par22, Qiu24, Rah24, Rob24, Sch20, Sen24, Seo24, Sto22, Wan20, WFL22, Won24, Ye24, Zha20, Zha22]. Issue [Ano21d, KLSZ21]. Ivo [Qiu24]. IVX [YLPC20]. IVX-AR [YLPC20]. ix [Ell23].

J [LN20]. Jacob [Rug21, Sha21]. Japanese [ILMY22]. Jaroslaw [WFL22]. JASA [WHC^{+}24]. Jeffrey [Rah24]. Jihnhee [Zha20]. Jim [Beb20]. John [Beb20, Cho22, Ism20]. Johnson [Ni22]. Joint [BMDG20, GX23, LRY24, LKPG21, SS22, YQ23]. Jones [SL20]. Jorge [Che20a]. José [Wan20, Won24]. Journal [Ano24]. JSM [FWB22]. Juan [Bor24]. Judea [AS20]. Julia [HM23]. Jump [CFS23, CDF20]. Jumping [MC22]. Jumps [CDF20]. Jun [Har23, Jan23, ZM23].

Kallus [LLL21, LZ21b, VD21]. Kalman [KSW20]. Karl [Rah24]. Kateri [Hor22]. Kernel [DL22b, DLL23, DL23a, FGRS23, LZ21a, LJL23b, LKL20, THP^{+}23, Wan20, WNK24]. Kernel-Based [LJL23b]. Kernels [DL21, LTB22]. Kidney [CJM^{+}20b]. Kind [CDJ23]. Knockoff [BCJW21, LKLL22]. Knockoffs [DLL23, FLSU20, FDLL20, RWC23, RSC20]. Knots [YZZ23]. Knowledge [LZCL24, SLC21]. Kriging [WTW20]. Krivitsky [SF23, Yon23]. Kyprianou [Bor24].

L [Len20, Ni22, Sto22]. L2RM [KAZZ20]. Label [BMSL22, FLLZ22, YRTJ23]. Label-Noise-Adjusted [YRTJ23]. Laber [Zha22]. Lag [MKW^{+}24]. Lagrangian [SLG23]. Lamperti [Bor24]. Lamperti-Type [Bor24]. Landmark [ZBM^{+}24]. Landmark-Based [ZBM^{+}24]. Landmarks [DKPW21]. Langevin [QW21]. Large [AZ23, BMDG20, BF24, CLZ20, FDLL20, FGJS22, GT23, GCRZ24, GX23, HTA^{+}22, KT24, KLLL23, LSP20, LCL23, LSB^{+}22, MBH22b, NS23, RZCL22, SCACvdH24, SWPH20, WBBM23]. Large-Panel [KLLL23]. Large-Scale [BMDG20, FDLL20, FGJS22, GT23, GCRZ24, GX23, LCL23, LSB^{+}22, MBH22b, RZCL22, SWPH20, WBBM23, CLZ20]. Lasso [DREM^{+}21, TJH^{+}23, ZLZ24, NR23, ZDM21]. Lassos [BMA^{+}22]. Latent [AB23, CZG24, CJM^{+}20b, CLZ20, CZW21, FL22, FLY24, GX23, HMW23, IS23b, JKL^{+}23, PBOD23, QW21, RMM20, XYWQ24]. Laud [Ni22]. Lauritzen [All20]. Law [BCCP22]. LAWS [CSX22]. Lawson [Cho22]. Layer [DWR23, LL23]. Lead [LO21]. Lean [PK23, VDVM24]. Learning [Ban22, BLLY22, CCgX23, CZG24, CP23, CZW21, CHYL23, CLSG24, DSW22a, DL23a, DYFQ23, DLS24, DL21, EMOS21, FL22, FWW23, FDAZ22, GZW24, HSB21, HM23, HQCC21, INGP23, Kal21a, Kal21b, LLL21, LWWftADNI21, LSN23, LCL23, LLgX22, LLK^{+}20, MKftADNI24, MBH22a, MQL21a, MQL21b, MML23, NBW21, PZ21a, PLCS21, QLFL20, QMZ24, RLY^{+}23, SLgX20, SWL^{+}23, SZS24b, SLL^{+}24, SSL22, TBQ20, TF23b, TJK22, VD21, Wan23, WLGC24, WZW20, XWMD24, XZZ^{+}22, ZDY23, ZLZ24, ZXS^{+}22a, ZWSZ23, ZZQ24, NMSD21]. Least [GC22, SBY24]. Length [GLM23, KL21b, RB20]. Length-Biased [RB20]. LESA [CHDP23, ZK23, AL23, ZWX^{+}23a, ZWX^{+}23b]. Lesion [MNHG24]. Level [EM23, NLS^{+}21, YQ23, ZZH22, MNW^{+}20, QZ22]. Leverage [ZLZ23a]. Leveraging [SZS24b, ZZH22]. Lévy [Bor24]. Li [McD20]. Lieshout [Sch20]. Life [PZ21b, Rob24]. Likelihood [BAXV22, CCH20, DG20, FNDK23, GK22, LSP20, LCL22, MR23, PT23, WZL^{+}23, YWAZ22, YB24, Zha20, ZSP20]. Likelihood-Based [BAXV22, MR23]. Limit [LRY24, SPMK22]. Limited [Har23]. Limits [TLT^{+}24]. Lin [CJZ23, Jan23, XC23b, ZM23]. Line [PLZ24]. Linear [ABS^{+}22, BCG20, CGM23, CS23, CWF^{+}24, CJK^{+}24, CJZ23, DLXL23b, DLXL23c, Fry24, Jan23, JCY^{+}21, Joc22, KAZZ20, KKS23, LB23, LR23a, LCL22, Li22, LYZ23, LZCL24, MYB^{+}20, Mur21, PSVV24, RS22, SBY24, SL22c, TZ24, TF23b, XC23b, XYY24, YZZ23, YLSL20, ZM23, ZZL24, ZKZ22, Cho21]. Linear-Cost [CS23]. Link [MR20, YQ23]. Link-Based [MR20]. Linkage [AGS23]. Liquid [LZZ23a]. Lists [CSV21]. Liu [CJZ23, Jan23, LLL21, VD21, XC23b, ZM23]. Local [CCH20, CJM20a, Che23, DHS23, JT22, KLLL23, RRRF22, YZL22, Zan20]. Locally [CSX22, DP21]. Location [LRZ24]. Locations [IS23b, ZSW22]. Loci [GLLZ22]. Log [FGRS23, Mur21]. Log-Linear [Mur21]. Log-Rank [FGRS23]. Logic [SL22a]. Logistic [HXGW22, MCL21a, Mur21, ZPZW23]. Lohr [Sto22]. Long [LRS20, LKM21]. Long-Range [LRS20]. Long-Term [LKM21]. Longitudinal [AL23, CHDP23, DWHS23, FBFV21, GYY^{+}22, HQCC21, LLZZ22, PLCS21, WRG23, WNK24, ZK23, ZWX^{+}23a, ZWX^{+}23b, ZLGY23]. Look [ZX24]. Loss [CBK^{+}22]. Low [CLX23, KAZZ20, PLZ24, WWZ22, ZKZ22]. Low-Rank [CLX23, KAZZ20, PLZ24, WWZ22].

M [Sch20]. MA [Par22]. Maathuis [All20]. Machine [Ban22, CT21a, DL23a, HM23, INGP23, TJK22, WQ23, WLGC24, ZXS^{+}22a, NMSD21]. Mackenzie [AS20]. Macroeconomic [MANW20]. Macular [JCY^{+}21]. Maharaj [Che20a]. Mahlet [Ni23]. Main [RKJ^{+}20]. Makes [Ste23]. Making [CLS21a, CLS21b, CT21a, GZW24]. Malaria [HS22]. Malawi [LNB22]. Mallows [LXLF20, ZJLD23]. Manifold [YDV24]. Manifolds [LD24, YZ20]. MANOVA [HHH23, LLM23b]. Many [Fer21, HTA^{+}22, Joc22, KTDS20, SWPH20, ZSW22]. Many-Input [SWPH20]. Mapping [LWY^{+}21, WIS20, YWKZ22]. Maps [BK20, KS24]. Marginal [Coo20, KKS23, MCLT24]. Marginalized [GKB^{+}21]. Maria [Hor22, Len20]. Marie [Zha22]. Marina [Ni23]. Mark [Beb20, XZLG24]. Mark-Point [XZLG24]. Marked [LNB22, XZLG24]. Markets [AB20]. Markov [AHDP24, AM24, BCI24, GSS22b, Ize21, LXJ22, NF21, SZS^{+}24a, WST^{+}20]. Markov-Switching [BCI24]. Marloes [All20]. Martens [Gio23]. Martin [All20]. Martingale [Seo24]. Masking [JRSW22]. Mass [Pai23, ZRPB24]. Mass-Shifting [ZRPB24]. Massive [YWAZ22]. Matched [BRS22, PK20, WZW20]. Matching [AS22, AB23, KML23, KMPT21, NLS^{+}21, SW23, WPTM24, Won24, WMK^{+}24, ZSL^{+}23]. Matchings [MAZB22]. Materials [YSL24]. Matérn [MB23]. Maternal [MKW^{+}24]. Mathematical [Sen24]. Mathias [All20]. Matrices [BF24, CvS21, DPY22, FFHL22, PND23, SL22a, YDV24, YLXL23, YLX24]. Matrix [ABD^{+}21, BN21, CTC20, CF23, HCZY24, JM20, KML23, KAZZ20, LZQ^{+}21, MCL21b, ZSK23, ZCW23]. Matrix-valued [ZSK23]. Matrix-Variate [CTC20, CF23]. Matt [WFL22]. Matter [TTPD20]. Max [BSH21]. Max-Infinitely [BSH21]. Maximin [WSX22]. Maximum [ACADF23, CCH23a, GK22, HZQ^{+}21, PT23, YWAZ22, ZSP20]. MCMC [Zan20]. Mean [ACADF23, DL23b, ISSM23, JWW^{+}24, LW22a, LYZL24, MCLT24, PSVV24, YLXL23]. Measure [DS23, PWZ^{+}20]. Measured [LCD22, PLSZ23]. Measurement [Cha23, HW21]. Measurements [HTA^{+}22]. Measures [AM22, ACV20, KY20, RB20]. Measuring [Ban22, TJK22, ZXS^{+}22a, Ism20]. Mechanism [ZM22]. Mechanisms [AS21, GLLT21]. Mechanistic [BIK20, WHHZ22]. Median [SW21a]. Mediating [GLLZ22]. Mediation [DSL22, GLLZ22, LSB^{+}22, SL22a, XC23a, XTK^{+}22]. Mediators [GSS23a]. Medical [LA20]. Medicine [Ano21d, Ber24, KLSZ21, RLC^{+}21, Zha22]. Medium [HMB^{+}20]. Meets [LKL20]. MEG [MZW23]. Mehler [LTB22]. Membership [SCMW24]. Memberships [MSC21]. Memory [QL24]. Mendelian [HS22]. Mental [GZW24]. Meshed [PBF22]. Meta [MV20, ZHC21]. Meta-Analyses [MV20]. Meta-analysis [ZHC21]. Meter [BTH21]. Method [BMFR21, CMW24, CWZ21, CDJ23, DH22, HS21, LBN21, NCMK20, NS23, ZLZ23a]. Methodological [DFH^{+}21]. Methods [AE21, Ano21d, ABD^{+}21, Ber24, CFT21, DKJ24, GK22, HHLT21, Ism20, KSW20, KC22, KLSZ21, LN20, Li20, McD20, QCS^{+}21a, QCS^{+}21b, Seo24, WZL^{+}23, Zha20, Zha22]. Methylation [GLLZ22]. Metric [DYFQ23, Wan23]. Metrics [BPvdL20, ZLQ^{+}21]. Metropolis [KR23, ST23a, WWW22]. Metropolized [BCJW21]. Michael [Ell23, Par22]. Micheas [Beh20]. Microbial [TJH^{+}23, ZPZW23]. Microbiome [GGR^{+}20, JZW23, LTO^{+}21, MCM20, WCL21]. Microclustering [BZS22]. Mild [TTL^{+}21]. Militarized [OPI22]. MIMIX [GGR^{+}20]. Mini [WWW22]. Mini-Batch [WWW22]. Minimal [FDNZ24].

Minimax [CFWY23, LLL22b, WW22]. Minimax-Optimal [CFWY23]. Mirrors [XZL23]. Mismatch [SLC21]. Missing [Ano22b, BN21, DHL20, MST21, MST22, MP21, Pai23, SKL22, STC^{+}20, XQ21, ZD24, YLSL20]. Missing-Not-at-Random [Ano22b, MST21, MST22]. Missingness [ZM22]. Mixed [BCG20, GGR^{+}20, GYY^{+}22, HJO^{+}20, JCY^{+}21, KKS23, LR23a, LCL22, LLZL23, MJ23, MSC21, PLCS21, RKJ^{+}20, SCMW24, XCZW23, ZSL20]. Mixed-Effect [ZSL20]. Mixed-Effects [GGR^{+}20, LCL22]. Mixed-Response [XCZW23]. Mixing [DL22a]. Mixture [Che20c, DSGS22, LZCH21, SKL22, SJW23, Sug21, TET^{+}23, XX20]. Mixtures [MKV^{+}24, ZSW22, ZW20]. MLE [GX23]. Mo [LLL21, LZ21b, VD21]. Mobile [BMDG20, GJK^{+}20, HQCC21, LKM21, LLK^{+}20]. Modal [ZKR23]. modality [YLSL20]. Mode [MC22]. Mode-Jumping [MC22]. Model [AB20, BSH21, CJM^{+}20b, CS24, DGH23, DSGS22, DCGM23, DKPW21, FYY23, GT22, GDP22, GKB^{+}21, GGR^{+}20, GHL24, HARF^{+}20, HJO^{+}20, HXC23, HyLXZ21, HMZZ23, HZB^{+}24, HXZ20, HZLG22, JRSW22, JZW23, KL21b, KKS23, KPS23, LBN23, LW20, LXLF20, LLWL22, LZZ23b, LSN23, LSZ24, LY22, LKLL22, LYC23, LLZL23, LKL20, MLL21, MCL21b, MJ23, MZPD22, MST21, MST22, MMR24, MKZ^{+}21, SMO23, SJW23, SHR^{+}20b, TLX24, TJH^{+}23, TCYL23, Wag20, WW22, WHX20, WSM^{+}23, WNK24, WHHZ22, WWF23, XX20, XCZW23, XQ21, YLPC20, ZPZW23, ZSL20, ZZLC20, ZNBR22, ZLZ23a, ZZXL20, ZHC21, ZCM22, ZJLD23, Ano22b, LN20]. Model-Assisted [DGH23, WWF23]. Model-Based [MZPD22, LN20]. Model-Free [LKLL22, MMR24, TCYL23, ZZXL20, ZLZ23a]. Model-Robust [WW22, WSM^{+}23]. Model-X [GHL24]. Modeling [Ano23b, BGOT20, BMDG20, BKR^{+}23, BFLZ22, Coo20, DGD21, DKPW21, DM22, FBFV21, GT22, GT23, GLPD23, HBD24, HBP^{+}20, HR24, KW24, LNB22, LLLZ23, LZQ^{+}21, LZCH21, LYL22, LKPG21, LOW21, MBH22a, MYB^{+}20, MNBCCC24, PBF22, PSVV24, QW21, RLYI20, Rob24, Sug21, TET^{+}23, WZLL22, WZC23, XWMD24, XYWQ24, ZSW22, ZPYZ24, ZB22, ZROD23]. Modelling [Che20d]. Models [ABD^{+}21, BMA^{+}22, BSM23, BKR^{+}23, BZS22, BFLZ22, BCG20, BHW20, BIK20, CGM23, CKK24, CTC20, CCLS21, CYZ22a, CHYL23, CF23, CJZ23, DLL23, DLXL23c, DM21, DY21, DOH22, DHS23, DV24, DAGO22, DWR23, FMM22, FGY24, Fry24, GK22, GMVFX22, HLL22, He24, HKZ23, HL21a, HBX23, HQYZ20, HZQ^{+}21, HMW23, Ism20, Ize21, Jan23, Jew21, JCY^{+}21, JVRS23, Joc22, Kan21, KSW20, KAZZ20, LB23, LR23a, LJL^{+}23a, LL23, LD20, LCL22, Li22, LYZ23, LCL23, LZCL24, LZCH21, LT22, LWSL22, LZH24, LRZ24, LC20, LW22b, MCL21a, MR20, MA23, MCLT24, MML23, MP21, MKW^{+}24, Mur21, NCMK20, OY22, OWP20, PBOD23, PLL23, PLZ24, PLCS21, Peñ22a, PSVV24, PZ21b, QDL21a, QDL21b, RPLG22, SS22, SKL22, SSB21, SBY24, SSLL21, SA20, SL22d, TBC20, TZ24, TF23b, TLT^{+}24, THP^{+}23, Vat20, WBK^{+}22, WST^{+}20, XC23b, XX21, YFZ20, YZZ23, YYZ24].

Models [YWW^{+}20, YSLW22, ZMNVF22, ZM23, ZL23, ZZL24, ZKZ22, Cho21, DLXL23b, All20, Cha23, Sen24]. Moderate [GC22]. Moderate-Dimensional [GC22]. Modern [Beb20, Har23]. Modification [LY22]. Modified [DWHS23]. Moments [HS21]. Momiao [Ché20b]. Monitors [GJK^{+}20]. Monotone [MCL21b, Ano22b]. Monsoon [JSK^{+}20]. Monte [NF21, DHJW22, DL22a, Dua23, GSS22b]. Mortality [BKR^{+}23, NLS^{+}21]. Most [GV21, GWL^{+}23]. Movement [AB20, GBL^{+}21]. Moving [WBBM23]. MPHIA [BLLY22]. Multi [Ban22, CKTVB21, HL21a, LL23, LM21, MKftADNI24, QLFL20, QIL24, QZ22, TTPD20, TJK22, Wan23, WCZZ24, XCZW23, YLSL20, YQ23, ZDNH23, ZXS^{+}22a, ZKZ22]. Multi-Armed [QLFL20, QIL24]. Multi-category [ZDNH23]. Multi-Cutoff [CKTVB21]. Multi-Dimensional [XCZW23]. Multi-Goal [HL21a]. Multi-Layer [LL23]. Multi-Level [YQ23, QZ22]. Multi-modality [YLSL20]. Multi-Outcome [TTPD20]. Multi-resolution [LM21]. Multi-Response [ZKZ22]. Multi-Source [Ban22, TJK22, ZXS^{+}22a]. Multi-State [WCZZ24]. Multi-subject [QZ22]. Multi-Task [MKftADNI24]. Multi-View [Wan23]. Multicategory [XZZ^{+}22]. Multicollinearity [Yon23]. Multidimensional [PXZ21, WWZ22]. Multidirectional [TXQ21]. Multifile [AGS23]. Multilabel [ZSJW23]. Multilayer [TBQ20]. Multilayered [HSB21]. Multilevel [HZLG22]. Multimodal [CHKW24, DL23a, HZLG22, LL22, LZZ23a, QW24, XLLJ20]. Multinomial [AFDZ23, HXGW22, Mur21, ZPZW23]. Multiparameters [LLgX22]. Multiple [BSM23, CSX22, CSV21, CZW21, CPH22, DSL22, DGD21, Fry24, FGJS22, Gug23, HHH23, JRSW22, KSR21, KKS23, LLWL22, LKPG21, MHOZ23, PRV22, PLZ24, SLG23, SSB21, TLT^{+}24, YD20, ZC22]. Multiple-Domain [CZW21]. Multiple-Output [Gug23, HHH23]. Multiple-Test [PRV22]. Multiple-Testing [DSL22]. Multiplicative [HMMMN21]. Multiresolution [SWPH20]. Multiresponse [LC20]. Multisample [MAZB22]. Multiscale [SPV20, VBM22]. Multisource [XQ21]. Multistate [WST^{+}20]. Multivariate [BCT24, CZG24, CLS20, Coo20, DS23, HHLT21, HXGW22, HMW23, MANW20, MR23, MKZ^{+}21, PZ21b, SKL22, TL23, ZW20, ZRPB24, ZKZ22, Wan20]. Mutational [LZ21a].

N [Ism20, Sch20, SF23]. Nagaraja [Ism20]. Narrowest [Fry24]. National [KCPW21]. Natural [BHW20]. Near [MP24, WZC23]. Near-Optimal [WZC23]. Nearest [CX23, DFG^{+}24]. Nearly [CWF^{+}24]. Neighbor [CX23]. Neighborhood [LXJ22]. Neighbors [DFG^{+}24]. Nested [BFLZ22, DCGM23, LW22b]. Network [CZG24, CP23, DOH22, GR21, HZB^{+}24, JKLW23, LBN21, LBN23, LO21, LOW21, OSDvdL24, OPI22, PK23, SLSH20, TFS21, WZL^{+}23, ZSL20, ZHW22, ZCM22]. Networks [Ano23b, BCI24, BAXV22, CKY22, DL21, DM22, FAS^{+}22, FAM21, HSB21, HXC23, IS23b, KCH23a, KCH23b, LWY^{+}21, LTB22, LKPG21, MNW^{+}20, Nie23, QW21, SF23, SZZ24, TJH^{+}23, WCNL24, Yon23, ZLZ24]. Neural [DL21, LTB22, MLH^{+}22, QW21]. Neurocognition [WCZZ24].

Neuroimaging [BMSL22, HZLG22]. Neuronized [SL22c]. Newton [CLZ22]. Newton-Type [CLZ22]. Next [WJM22]. Neyman [WWLT24, XZWT21, YRTJ23]. NHANES [LWLL23]. Nickl [Sen24]. Niel [SF23]. Nikolaos [Efr22]. Nishiyama [Seo24]. NJ [Ism20]. No [Ano22b, LSRR21, MST21, MST22]. Nodal [WCNL24]. Noise [RRRF22, YRTJ23]. Noisy [CKY22, CMZ20, CFWY23, LSS23, MKftADNI24, MBCK22]. Non [Ano22b, GSS22b, KS24]. Non-differentiable [GSS22b]. Non-Gaussian [KS24]. Non-monotone [Ano22b]. Nonasymptotic [BMD^{+}24]. Nonbifurcating [ZDM21]. Noncompliance [BP23, IJM21, ZHC21]. Nonconvex [CFWY23]. Nondiscrete [Wes22]. Nonhomogeneous [LW20]. Nonignorable [FLSZ20, ZM22]. Nonlinear [BPvdL20, FDLL20, GS23, HH23, JT22, LSP24, LDL^{+}22]. Nonlinearity [AM22]. Nonmonotone [MST21, MST22]. Nonnegative [JM20, LZQ^{+}21]. Nonoverlapping [CSV21]. Nonparametric [ARN20, BC23, CFMB24, CSdG^{+}23, CLS20, Cre22, DLL23, DS23, DFG^{+}24, DCGM23, DREM^{+}21, DWHS23, Efr22, GSS23b, GSS23a, Gho22, GK22, GRLB20, Hof22a, IW22a, IW22b, ISSM23, Imb22, JWW^{+}24, LLLZ23, LLgX22, MKV^{+}24, MJH20, Pen22b, QDZ24, QL24, RB20, RBMB21, RPLG22, Wes22, XS21, XS22, YFZ20, ZW20]. Nonprobability [CLW20]. Nonregular [EM23]. Nonresponse [FLSZ20]. Nonsparsity [ZZL24]. Nonstationarity [LRZ24, MM21]. Nonstationary [BCT24, BTHK20, DL23b, HARF^{+}20, Mie23, SS22]. Nonzero [DL23b]. Norm [AS21, ZGZC20]. Normal [ZPZW23, ZRPB24]. Normality [CX23]. Normalized [ACV20]. Notes [LZQ^{+}21]. Nugget [LSZ24]. Nuisance [SZ21]. Null [Wes22, WD21]. Number [FGZ22, JKLW23, KML23, NCL22, SCACvdH24, TMNM22, ZZLC20, CRYZ23]. Numerically [MOA24]. Nutrition [LDL^{+}22]. Nutrition-Environment [LDL^{+}22]. Nutritional [SPMC21]. NY [Bor24, Che20d, Sen24, WFL22]. NYC [RBMB21].

O [Ni22]. Oaxaca [GB23]. Oaxaca-Blinder [GB23]. Object [DLP23, WPTM24]. Objectives [Kal21b, LLL21, VD21]. Objects [Ano23b, Beh20, DM22]. Observable [FDF20]. Observation [JAO23]. Observational [EB21, Fog20, FAM21, FDF20, GSS22a, KSR21, ST21, YPL^{+}21, YD20, ZSL^{+}23]. Observations [KCPW21, LHW23, SSB21]. Observed [BAXV22, FSV20, IAPK23, ZSZL23]. Obstruction [CJM^{+}20b]. Oceanographic [TET^{+}23]. Off [LKM21, SZS^{+}24a]. Off-Policy [LKM21, SZS^{+}24a]. Offline [SLL^{+}24]. Older [BKR^{+}23]. Omics [Ché20b, ZC22]. Omics-Wide [ZC22]. One [QL24, XX23, YB20, ZSL^{+}23]. One-Pass [QL24]. One-Step [XX23]. Online [CLS21a, CLS21b, CWS24, DJM23, GSW23, MOA24, RLY^{+}23, SSLL21, YYZ24, Yao23, ZCW23]. Online-Score [SSLL21]. Only [LNB22, SR20]. Open [BK20]. Operator [LLLZ23, YSL24]. Operator-Induced [YSL24]. Optima [Che23]. Optimal [CFWY23, CT21b, DFG^{+}24, DAGO22, FWW23, GPPV20, GO22, GZM21, HFR20, HXZ20, HLW23, JKLW23, Kal21a, KG23, Kow22, LD24, LSRR21, LLL22b, MCL21b, MQL21a, MQL21b, MAZB22, OWP20, PZ21a, PK20, QLFL20, QCS^{+}21a, QCS^{+}21b, QL24, SM21, SMN^{+}20, SW21b, TZL20, WCL21, WJW^{+}21, WWF23, WZC23, XZZ^{+}22, XYY24, YLSL20, YWAZ22, ZWSZ23, ZZQ24]. Optimization [BDGJ23, CLX23, CFWY23, CMW24, FGY24, JPR20, QW21]. Optimizing [FWB22, WZW20]. Order [CLZ22, ISSM23, SBY24, YQ23]. Ordinal [LLYM21, Wei20]. Ordinary [DL22b, GC22, OWP20, THXZ23]. Oriented [FWW23, RZCL22]. Origins [WRG23]. Orthogonal [ST23b]. Orthogonality [WSX22]. Orthogonalized [DL23a]. Other [DHHK21]. Our [Mar21]. Outcome [GSS22a, TTPD20, ZZH22, ZDNH23]. Outcome-Dependent [GSS22a]. Outcomes [BFLZ22, CGM23, CZW21, CFVMW20, CMC^{+}20, FLKK21, GO22, HSC21, LKM21, QLFL20, SMN^{+}20, YFZ20, ZWSZ23]. Outer [CXH23]. Outer-Product-of-Gradient [CXH23]. Outlier [SWS22]. Output [Gug23, HHH23]. Outward [HLL22, HHH23, SDH22]. Over-Estimated [FL22]. Over-Parameterized [FYY23]. Overall [SLS20]. Overcoming [SWR24]. Overfitting [LSS23]. Oxford [Gio23, Rob24].

P [Che20d, HM23, Rob24, Rug21, Sch20, Sen24, Sha21, WFL22]. Pablo [Rob24]. pages [Par22]. Pain [GYY^{+}22]. Pair [XLWG23]. Paired [Fog20]. Pairs [BRS22]. Pairwise [GVW20, HXC23]. Paleoclimate [HLS^{+}21]. Pandemic [TTL^{+}21]. Panel [AB20, ABD^{+}21, BIK20, JVRS23, KLLL23]. Panels [AZ23]. Pankaj [Ism20]. Paper [CS20, Cox20, Ste23, Wag20]. Parameter [SS22, Yee22]. Parameterized [FYY23, Kow22]. Parameters [DHS23, DG20, ZZLC20]. Parametric [AM21, FBFV21]. Parasitic [TJH^{+}23]. Pardo [Bor24]. Parsimonious [ZZLC20]. Part [HSC21, Ano21d]. Partial [JAO23, LJL23b, LLYM21, MLL21, MP23, QZ22, SBY24]. Partially [BAXV22, CJK^{+}24, IAPK23, ZSZL23]. Particle [BDGJ23, CDJ23, JLS20a, JLS20b, MOA24]. Particle-Based [MOA24]. Particulate [TTPD20]. Partition [BZS22, ZJLD23]. Partition-Mallows [ZJLD23]. Partitioned [PBF22]. Partitioning [AGS23, LLB^{+}22]. Partitions [SA20]. Partnering [WHC^{+}24]. Pass [QL24]. Past [GV21]. Pathology [ZBM^{+}24]. Pathway [YWKZ22]. Patient [PXZ21, RLC^{+}21, STC^{+}20, WWLT24]. Patient-Derived [RLC^{+}21]. Patients [BKR^{+}23]. Pattern [HXGW22, RZCL22]. Pattern-Oriented-Sampling [RZCL22]. Patterns [WJIS20]. Paul [Cha23, Sha21, Won24]. Pavel [SF23]. PBDEs [LWLL23]. PCBs [LWLL23]. PDMP [CFS23]. Peace [Rah24]. Pearl [AS20]. Pearson [WWLT24, XZWT21, YRTJ23]. Pedigrees [JCY^{+}21]. Peer [EB21, IS23b]. Pelvic [GYY^{+}22]. Penalization [TZ24]. Penalized [AL21, CCLS21, JPR20, SZ21, YYLL20]. Penalties [Ano22c, DREM^{+}21, RT20a, RT20b, SH20]. Periodic [ASTZ24, HARF^{+}20]. Periodogram [CCH23a]. Periodontal [GRLB20]. Periods [Fer21].

Permutation [BC23, KL21a, LJL23b, MCL21b, MKV^{+}24, SCACvdH24]. Permutation-Based [SCACvdH24]. Permuted [MCL21b]. Personalized [DSWQ21, HQCC21, LWSL22, MKW^{+}24]. Perspective [GDP22, Wan23]. Perspectives [INGP23]. Perturbation [GCW20]. Peter [LN20]. Pharmacogenomics [Ber24]. Phase [TZL20]. Phenomenon [ZRPB24]. Phenotypes [XCZW23]. Phenotypic [HTA^{+}22]. Phenotyping [NMJ20]. Philadelphia [BDGJ23]. Phylogenetic [ZDM21]. Pierpaolo [Che20a]. Pierre [Sha21]. Pietro [SF23]. Pitman [MKV^{+}24]. Plant [PBOD23]. Platforms [ZLQ^{+}21]. Plug [MKZ^{+}21]. Plug-in [MKZ^{+}21]. Plus [CLX23, KSR21]. PM [LZCH21]. Point [BSM23, CZG24, DG20, DPY22, DREM^{+}21, HXGW22, MNBCCC24, PSBB21, TL23, XLWG23, XZLG24]. Points [Fry24, Yee22, ZWS22, CRYZ23]. Poisson [JZW23, LW20, MKV^{+}24, MNBCCC24]. Poisson-Gamma [JZW23]. Policies [DWHS23, NBW21]. Policy [Ano21d, FGY24, GRLB20, HQCC21, Kal21a, Kal21b, KLSZ21, LLL21, LKM21, RLY^{+}23, SZS^{+}24a, VD21, VB24]. Pollution [GJK^{+}20, LSD21, MKW^{+}24]. Pólya [AM24]. Polygenic [CCLS21]. Polyhedral [KL21b]. Polynomial [CJM20a]. Polytopes [DW21, HHZ21, JGED21a, LV21, Rug21, Sha21, Wil21, JGED21b]. Pooled [LWLL23]. Pooling [XCS20]. Population [DGH23, DKJ24, MNW^{+}20, SHWB23]. Populations [KTDS20, LOW21]. Portfolio [DGW23]. Positive [CvS21, YDV24]. Positive-Definite [YDV24]. Possible [LZS23]. Post [AS22, KL21b, LKL20]. Post-Matching [AS22]. Post-Model-Selection [KL21b]. Post-Regularization [LKL20]. Posted [ZQL^{+}23]. Posterior [Dua23, LSZ24]. Postoperative [BKR^{+}23]. Postprocessing [HHLT21]. Potential [DDX22]. Poverty [RLS23]. Power [GCW20, Yon23, YLXL23]. Power-Enhanced [YLXL23]. Powerful [LX20]. pp [All20, AS20, Beb20, Beh20, Ber24, Bor24, Cha23, Ché20b, Che20d, Che20a, Cho22, Ell23, Gio23, Har23, HM23, Hor22, Ism20, LN20, Len20, McD20, Muk24, Ni22, Ni23, Qiu24, Rah24, Rob24, Sch20, Sen24, Seo24, Sto22, Wan20, WFL22, Won24, Ye24, Zha20, Zha22]. Practice [HTF23, YSYZ23]. Precipitation [BGOT20, BSH21, CCH20]. Precision [Ano21d, Ber24, KLSZ21, RLC^{+}21, WSM^{+}23, Zha22]. Preconditioned [NS23]. Predict [CJM^{+}20b]. Predictability [YLPC20]. Predicting [BMDG20, CMC^{+}20, PBOD23, TMNM22]. Prediction [Ano22c, Ano23a, BPvdL20, CS20, CFT21, Cre22, DSWQ21, DWR23, Efr20a, FHT20, HL24, JAO23, KT24, LLWL22, MJ23, MMR24, MD20, NCMK20, PLSZ23, PS23, PLL23, RLYI20, RT20a, RT20b, SH20, STC^{+}20, TFCF21, WW22, WTW20, gXZ20, YLSL20, YQ23]. Predictions [Kow22, LM21, MN23]. Predictive [DP21, LLPY24, MANW20, Qiu24]. Predictors [GR21, KKS23, PLL23, RLS23]. Predominantly [XTK^{+}22]. Pregnancy [BFLZ22]. Preliminary [PRV22]. Preprocessing [CLSG24]. Presence [DSGS22, GDP22, HBP^{+}20, LNB22, MHOZ23, RRRF22, SR20, SYD^{+}22]. Presence-Only [LNB22, SR20]. Preserving [AM21, KSB23, NCMK20].

Press [All20, Beb20, Beh20, Ber24, Bor24, Cha23, Ché20b, Che20d, Che20a, Cho22, Ell23, Gio23, Har23, HM23, Hor22, LN20, Len20, McD20, Muk24, Ni22, Ni23, Par22, Rah24, Rob24, Sch20, Sen24, Seo24, Sto22, Wan20, Won24, Ye24, Zha20, Zha22]. Pressure [ZLGY23]. Prevalence [DHL20, Jew21, SWR24]. Price [YLPC20]. Prices [ILMY22, RBMB21]. Pricing [FGY24, SL22b]. Principal [ASSS21, CMZ20, CLLY22, HY24, KLLL23, KHK21, YZ20]. Prior [HL21a, MNHG24, NS23, ZNBR22]. Prior-Preconditioned [NS23]. Prioritizing [LWSL22, WWLT24]. Priors [CFMB24, GSS22b, MNW^{+}20, RS22, RPLG22, SBJ20, SL22c, ZRPB24]. PRISM [YNCK21]. Privacy [AM21, AS21, Dre23]. Privacy-Preserving [AM21]. Private [BWSB24]. Probabilistic [BTH21]. Probabilities [LD20]. Probability [Beh20, JRR22, MW20, TLT^{+}24, YLX24]. Probit [AFDZ23]. Problem [QIL24]. Problems [NCL22]. Procedure [DSL22, KG23, XX23, ZDY23, ZM22]. Procedures [MQLH20, MLZH24]. Process [CFMB24, LSZ24, LTB22, MNBCCC24, SZS^{+}24a, SHR^{+}20b, TL23, WGK24, YB20]. Processes [Beh20, Bor24, CZG24, DOH22, GLM23, HBD24, HLS^{+}21, LW20, LYL22, PSBB21, PBF22, XLWG23, XZLG24]. Processing [MP21]. Product [CXH23, XX23]. Professor [gXZ20]. Profile [WZL^{+}23]. Profile-Pseudo [WZL^{+}23]. Program [FWB22]. Progression [WST^{+}20]. Progressive [SWS22]. Projected [XX21]. Projection [CLS20, LYZL24, LZH24, MD20, SLS20, ST23b, WSX22, ZZH22, ZLZ23b]. Projection-Based [SLS20]. Projections [FL22, SSB21]. Projects [MBH22b]. Propagation [HJO^{+}20]. Propensity [DG23, WMK^{+}24]. Properties [Fer21, Hof22a, MLZH24, NCMK20, WTW20]. Property [ILMY22]. Proportion [BK22]. Proportional [DV24, HMMMN21]. Proportions [HBP^{+}20]. Proposal [WWW22]. Proposals [Zan20]. Proschan [Ell23]. Protected [CLX22]. Proteogenomic [SJW23]. Provable [DL21]. Provides [WJIS20]. Proximal [CPS^{+}24, QMZ24]. Proxy [HLS^{+}21]. Pseudo [MJ23, WZL^{+}23]. Pseudo-Bayesian [MJ23]. pT [ZZQ24]. pT-Learning [ZZQ24]. PTSD [XTK^{+}22]. Public [LN20, SHWB23]. Public-Use [SHWB23]. PUlasso [SR20]. Pursuit [Fry24, ZZ23b]. Purushottam [Ni22]. Python [Hor22].

Q [EMOS21]. Q-Learning [EMOS21]. Qi [LLL21, VD21]. Qiu [ZP21]. QTLs [WIS20]. Quadratic [GC22, LLL22a]. Quadrature [BST24]. Quadruply [XC23a]. Qualitative [SLS20]. Quantification [FDAZ22, LLYM21, ZX24]. Quantile [AB20, CFVMW20, CSX21, DG23, FWZ^{+}21, FZHL23, FBFV21, Gug23, JCYS20, LSX23, LLPY24, VBM22, YBRM20, ZWKZ22, ZWZ22, ZKR23, ZLGY23]. Quantile-based [ZKR23]. Quantiles [SWS22]. Quantitative [XWMD24, Ber24]. Quantitative-Sequence [XWMD24]. Quasi [LCL22, YWAZ22]. Quasi-Likelihood [LCL22, YWAZ22]. Query [DYFQ23]. Query-Augmented [DYFQ23].

R [Har23, Hor22, Rah24, Rob24, Rug21, Won24, Cho22, HM23, Len20, McD20, Rah24, WFL22, Ye24]. R2 [ZNBR22]. R2-2D [ZNBR22]. Rainfall [JSK^{+}20]. Random [Ano22b, Ano22c, Ano23b, ACV20, BZS22, CFWY23, CS23, DGH23, DW21, Dua23, DM22, FFHL22, GLLT21, GK22, HMB^{+}20, HL21a, HHZ21, Ize21, JGED21a, JGED21b, JWW^{+}24, LR23a, LV21, LPR24, LZH24, MST21, MST22, MP23, MN23, MD20, RT20a, RT20b, Rug21, SBD23, Sha21, SH20, SLS20, SA20, TF23a, WW22, Wil21, XX23, YYLL20, YZL22]. Random- [SH20, Ano22c, RT20a, RT20b]. Randomization [LSD21, LRY24, MQLH20, MLZH24, ST21, WSM^{+}23, WD21, ZZ23a]. Randomization-based [LRY24]. Randomized [GSS23b, HFR20, IJM21, LH22, LRY24, PK23, WWL23, WD21, YSYZ23, ZD24]. Range [LRS20]. Rank [BSM23, CLX23, DS23, FGRS23, HHH23, HMZZ23, HL24, KAZZ20, LX24, PLZ24, PND23, TSW23, WWZ22, ZWZ24, ZJLD23, ZKZ22, FDLL20]. Rank-Based [DS23, HMZZ23]. Ranked [JHB20, LXLF20]. Ranking [DSWQ21, FGJS22, ZWSZ21]. Ranks [SDH22]. Rare [KG23, YB21]. RaSE [TF23a]. Rate [CJZ23, DLXL23a, DLXL23b, DLXL23c, DGSZ23, GSW23, GRZL23, Jan23, KSB23, LB23, LHS22, LYZ23, LCL23, XC23b, XZL23, ZC22, ZM23, CRYZ23]. Rates [BC23, BST24, LSZ24, WQ23]. Ratio [DG20, LSP20, SPMK22, YPL^{+}21]. Ratio-based [YPL^{+}21]. Raton [All20, Beb20, Beh20, Ber24, Cha23, Ché20b, Che20a, Cho22, Ell23, HM23, Hor22, LN20, Len20, McD20, Muk24, Ni22, Ni23, Rah24, Sch20, Seo24, Sto22, Wan20, Won24, Ye24, Zha20, Zha22]. RCTs [SPMK22]. Reactivity [GLLZ22]. Real [LZS23]. Real-Time [LZS23]. Reasoning [FBZ21]. Recall [GRLB20]. Recapture [DKJ24, MA23]. Recoding [GO22]. Recommender [LBF^{+}24]. Reconcile [HL21a]. Reconstructions [HLS^{+}21]. Reconstructions-Testing [HLS^{+}21]. Record [AGS23, GWL^{+}23]. Records [NMJ20, PLSZ23, SMMH21, WZW20]. Recovering [AB23]. Recovery [MCL21b]. Recurrent [SMMH21]. Recursive [LLB^{+}22, SSLL21]. Reduced [BSM23, TSW23]. Reduction [CXH23, McD20, XS21, ZMZ24]. Referenced [MNBCCC24]. Regimes [CT21b, LSRR21, LLK^{+}20, QMZ24, SMN^{+}20, SW21b, WWF23, XZZ^{+}22, Zha22, ZZQ24]. Region [JSK^{+}20]. Regionalization [LZCH21]. Regions [FBZ21, LLL22a]. Registration [MBCK22, WRG23]. Regression [ASSS21, Ano22c, ABS^{+}22, BMA^{+}22, BWSB24, BFLZ22, BCI24, BMSL22, CLX22, CKTVB21, CSdG^{+}23, CvS21, CLS20, CCLS21, CLM23, CJK^{+}24, CSX21, CS24, CPH22, DGS22, DM21, DY21, DLG23, DKPW21, FMW20, FLY24, FWZ^{+}21, FZHL23, FLSZ20, FBZ21, FBFV21, Gug23, GR21, HHH23, HMP20, HH23, He24, HVL23, HXGW22, ISSM23, JCYS20, Joc22, KG23, KCPW21, KC22, KAZZ20, LSX23, LLLZ23, LS20, LWWftADNI21, LL22, LLZZ22, LXJ22, LZX^{+}24, LLPY24, Loh20, LZS23, LW22b, MCL21a, MCM20, MR23, MKZ^{+}21, Mur21, NS23, OPI22, PLZ24, PZ21b, QDL21a, RS22, RBMB21, RPLG22, RT20a, RT20b, SH20, SLC21, SL22b, SL22c, SZF20, SZ21, TSW23, TL23, TZ24, TTPD20, TWM23, VDVM24, WPB^{+}20a, WPB^{+}20b, WBK^{+}22, WFL22, YFZ20, YBRM20, YQL23, YYZZ20, YYLL20, ZNBR22, ZWKZ22, ZWZ22, ZKR23, ZL23, ZLZ24, ZWLS20, ZB22, ZZ23b, ZLZ23b, ZSZL23, ZWZ24, ZLGY23]. Regression [QDL21b]. Regressions [Har23, LZZ23b]. Regularization [FYY23, LKL20, ZLD23]. Regularized [GO22, HH23, YS21]. Reinforced [KSR21, PLSZ23]. Reinforcement [GZW24, RLY^{+}23, SWL^{+}23, SLL^{+}24]. Reinforcing [Kaf20]. Rejection [KSB23]. Rejective [YQL23]. Rejoinder [CYZ22b, CT21a, DLXL23b, Efr20b, IW22b, JGED21b, Kal21b, KCH23a, LLR21b, MQL21b, QCS^{+}21b, QDL21b, RT20b, WPB^{+}20a, ZWX^{+}23b, ZXS^{+}22b]. Related [DGD21, DHHK21, Ize21, JCY^{+}21, PSBB21]. Relationship [LJL23b]. Reliable [MNW^{+}20]. Remote [KCPW21]. Repeated [QDZ24, SWR24]. Replicability [MBH22b, RLYI20]. Replicated [ISSM23, XZLG24]. Replication [DDX22, MBH22b]. Representation [DL21]. Representations [Bor24]. Representativeness [KCH23a, KCH23b, Nie23, SF23, Yon23]. Reproducibility [WHC^{+}24]. Reproducing [THP^{+}23]. Repulsion [NDH^{+}24]. Repulsive [XX20]. Required [ZQL^{+}23]. Rerandomization [WWL23, YQL23]. Research [Agr23]. Residual [ZZ23b]. Resistance [SWS22]. resolution [LM21]. Resolves [TJH^{+}23]. Resources [ZYZ24]. Response [CJK^{+}24, GK22, HJO^{+}20, Har23, MR23, XCZW23, ZQL^{+}23, ZSZL23, ZKZ22]. Responses [HMP20, KAZZ20, Li22, PLZ24, SHR^{+}20a]. Restricted [KC22]. Results [AFDZ23]. Retargeting [Kal21a]. Retirement [DGD21]. Retracted [JLS20a]. Returns [YLPC20]. Retweets [LW20]. Revealing [JHB20]. Reversible [CFS23, WWW22]. Review [Agr23, All20, AFDZ23, AS20, Beb20, Beh20, Ber24, Bor24, Cha23, Ché20b, Che20c, Che20d, Che20a, Chi21, Cho22, DFH^{+}21, Ell23, Gio23, Har23, HM23, Hor22, Ism20, LN20, LBF^{+}24, Len20, McD20, Mor21, Muk24, Ni22, Ni23, Par22, Par23, Qiu24, Rah24, Rob24, Sch20, Sen24, Seo24, Sto22, Wan20, WFL22, Won24, Ye24, Zha20, Zha22]. Richard [Sen24]. Ride [ZLQ^{+}21]. Ride-sourcing [ZLQ^{+}21]. Ridge [He24, LLL22b, ZWLS20]. Ridge-Type [LLL22b]. Riemannian [CLX23, LD24, YZ20]. Right [FGRS23, MR20]. Right-Censored [FGRS23, MR20]. Risk [CCLS21, HW21, ILMY22, JRSW22, LSN23, LLR21c, LWSL22, PLSZ23, SQPQ21, Wag20, ZZH22]. Risks [BKR^{+}23]. Rizzo [Len20]. RNA [HL21b, JJ21, LLR21b, LLR21a, LWSL22]. RNA-Seq [HL21b, LLR21b, JJ21, LLR21a, LWSL22]. Roach [Beb20]. Robin [Beb20]. Robust [AS22, AZ23, AM21, CLW20, CWZ21, DKJ24, EMOS21, FMW20, GMVFX22, Joc22, LS20, LTO^{+}21, LDL^{+}22, LZX^{+}24, Loh20, MW20, MQL21a, MQL21b, PZ21a, PSVV24, RSvdL21, SM21, SHO20, WW22, WPB^{+}20a, WPB^{+}20b, WSM^{+}23, XC23a]. Robustness [ASSS21, BMD^{+}24, QPL23, WTW20]. Rolf [Che20d]. Rong [LT22]. Rongling [Ber24]. Rosenbaum [Won24]. Rosner [Ni22]. Rosset [Wag20]. Rotational [GPPV20]. Rules [CZW21, FWW23, GZM21, Han21, HSC21, IL23, MQL21a, MQL21b, PZ21a, QLFL20, QPL23, QCS^{+}21a, QCS^{+}21b, ZWSZ21, ZWSZ23]. Ruobin [Sha21].

Ruppert [WFL22].

S [Jan23, Won24, ZM23]. Saddlepoint [JVRS23]. Salary [ZQL^{+}23]. Sample [BPvdL20, CHYL23, Fog20, GCRZ24, HL24, JWW^{+}24, LLR21c, PRV22, QDZ24, SST21, XCS20, ZGZC20, ZSL^{+}23]. Sampled [TBC20]. Sampler [DW21, HHZ21, JGED21a, JGED21b, LV21, Rug21, Sha21, Wil21]. Samplers [CFS23, DHJW22]. Samples [CLW20, KCH23a, KCH23b, Nie23, SF23, WJIS20, WIS20, Yon23]. Sampling [BCJW21, Cho22, DGH23, GSS22a, GBL^{+}21, Ize21, LR23b, MKZ^{+}21, NS23, QW24, RB20, RZCL22, ST23a, TMZ22, Vat20, WJM22, YQL23, ZJLH23, Sto22]. SAS [Rah24]. Scalable [DSWQ21, KS24, PBF22]. Scalar [FLSZ20, LWWftADNI21, LLZZ22, LYC23, YBRM20, ZWKZ22]. Scale [BMDG20, CJZ23, DLXL23b, DLXL23c, DFG^{+}24, FDLL20, FGJS22, GT23, GCRZ24, GVW20, GX23, GJK^{+}20, Jan23, KT24, LBN21, LBN23, LB23, LYZ23, LCL23, LSB^{+}22, MBH22b, RZCL22, SWPH20, WBBM23, XC23b, ZSW22, ZM23, CLZ20]. Scale-Free [CJZ23, DLXL23c, Jan23, LB23, LYZ23, XC23b, ZM23, DLXL23b]. Scale-Up [LBN23, LBN21]. Scaled [CFMB24]. Scan [LLL22a]. Schedule [SWR24]. School [RBMB21]. Schools [KSR21]. Science [AS20, Cre20, Gio23, OWP20, Qiu24, Rah24, YB20]. Scientists [Hor22, Rob24]. Score [GLPD23, SW23, SSLL21, ZLZ23a]. Score-Driven [GLPD23]. Scores [TFCF21, WMK^{+}24]. Screening [CSX22, GRZL23, LX24, LKLL22, NCL22, TF23a, TCYL23, XLYT20, ZQL^{+}23, ZLZ23a, ZZXL20]. Search [GRLB20, HARF^{+}20, LC20, SW21b, YNCK21]. Seasonal [HHLT21]. Section [AE21, ZHC21]. Segmentation [CCH23b, VBM22]. Segregation [SHWB23]. Selected [GH21]. Selecting [GLLZ22, ZWSZ21]. Selection [CWF^{+}24, CFS23, DLL23, DGW23, GDP22, GRZL23, HTF23, HL21a, HMW23, HLW23, KL21b, LLR21c, LR23b, LC20, RLYI20, SR20, TXQ21, TWM23, Wag20, WCNL24, XQ21, YB21, YSL24, ZMZ24, ZWLS20, CRYZ23, Ni23]. Selective [GBW24, PT23]. Self [Ano22b, MST21, MST22, Wan23]. Self-Censoring [Ano22b, MST21, MST22]. Self-supervised [Wan23]. Semi [ABS^{+}22, SLZ24]. Semi-Supervised [ABS^{+}22, SLZ24]. Semicontinuous [HSC21]. Semiorthogonal [LZQ^{+}21]. Semiparametric [CT21b, CPS^{+}24, FGY24, GZW24, HLL22, HyLXZ21, HXGW22, KPS23, LZ21a, LCD22, LLWL22, LY22, LYC23, MST21, MST22, PLCS21, SKL22, TLT^{+}24, XC23a, XLWG23, YSLW22, Ano22b, WFL22]. Semiparametrically [LRZ24]. Sensing [KCPW21]. Sensitivity [AS21, BK22, DG23, Fog20, FDF20, MV20, YSWB24]. Separability [LHGY23]. Separable [SYD^{+}22, SRS^{+}23]. Separation [WGK24]. Seq [HL21b, LLR21a, LLR21b, JJ21, LWSL22]. Sequence [CCH23a, CCH23b, DPY22, JZW23, XWMD24]. Sequencing [LLL22b, ZDNH23]. Sequential [DHJW22, DG20, GSW23, KPV22, LTY21, RMM20, ZROD23].

Sequentially [BFLZ22]. Series [Ano23a, BCT24, BTHK20, Che20a, CTC20, CYZ22a, CWW22, DP21, DFH^{+}21, DJM23, GT22, GT23, GMVFX22, HARF^{+}20, HCZY24, JKL^{+}23, JAO23, LRS20, LT22, Mie23, OY22, PS23, Peñ22a, QZ22, SHR^{+}20b, WZLL22, Wei20, ZPYZ24, ZX24]. Set [KSB23, LLL22b, SL20, WQ23, YLXL23]. Set-Based [SL20]. Set-Valued [WQ23]. Sets [DWR23, GCRZ24, ZLZ23b]. Setting [CLS21a]. Severe [WWLT24]. Sex [CPH22, LNB22]. Sex-Dependent [CPH22]. SGLD [WWW22]. Shannon [Zha22]. Shape [AL23, CHDP23, DKPW21, XLWG23, ZK23, ZWX^{+}23a, ZWX^{+}23b, ZBM^{+}24]. Shape-Constrained [XLWG23]. Sharon [Sto22]. Sharp [DG23, GSS23a, MSC21]. Shift [FDAZ22]. Shifted [MP23]. Shifting [ZRPB24]. Shortfall [SQPQ21]. Shrinkage [SBJ20, TTPD20, YDV24, ZNBR22, ZB22, ZLZ23b]. Side [BMS20]. Sieve [LKL20]. Signal [LLL22a]. Signals [GCW20]. Signature [RLYI20]. Signatures [LZ21a]. Significance [FDNZ24, Fry24]. Signs [SDH22]. Silge [HM23]. Simple [CJM20a, ZGZC20]. Simulating [Vat20]. Simulations [Beb20, SSB21]. Simulator [HLW23]. Simultaneous [HCZY24, LLL22a, MCL21a, MBCK22, RLS23, XLLJ20, XDM24, YLXL23, ZWLS20]. Single [FYY23, HKZ23, HL21b, JJ21, KPS23, LLR21a, LLR21b, LWSL22, LZH24, MKV^{+}24, MAZB22, ZWZ22]. Single-Cell [HL21b, JJ21, LLR21a, LLR21b, LWSL22]. Single-Index [KPS23, LZH24, ZWZ22]. Singular [HSZ24]. Site [ARN20]. Size [DKJ24, DKPW21, LNB22]. Size-and-Shape [DKPW21]. Sizes [QDZ24]. Skeleton [WC24]. Skills [ZQL^{+}23]. Slab [BMA^{+}22, MNHG24, NR23]. Small [BPvdL20, CWF^{+}24, HVL23, RLS23, Won24, Yon23]. Small-Sample [BPvdL20]. Smaller [Hof22b]. Smart [BTH21]. Smoking [ZDNH23]. Smooth [HMMMN21, YYZ24]. Smoothing [JLS20a, JLS20b, MOA24, Wan20, YSLW22, Par23]. Snippets [LW22a]. Social [FAS^{+}22, IS23b, OSDvdL24, SLSH20, TFCF21, WCNL24]. Society [Kaf20]. Solutions [Beb20]. Solving [CDJ23, NV24]. Some [AM22]. Sons [Ism20]. Source [Ban22, TJK22, WGK24, ZZH22, ZXS^{+}22a]. Sources [DREM^{+}21, LLgX22, YD20]. sourcing [ZLQ^{+}21]. Space [BGOT20, CXH23, KSW20, LWY^{+}21, PWZ^{+}20, THP^{+}23, XCZW23, Yee22]. Space-Time [BGOT20, KSW20]. Spaces [CWF^{+}24, Zan20]. Spanning [ZLZ24]. Spanning-Tree-Based [ZLZ24]. Sparse [BMA^{+}22, BMS20, CLX23, CSV21, CWF^{+}24, FLY24, FGQ24, HXC23, IS23a, LWWftADNI21, LLM23b, LKL20, MBCK22, MML23, NS23, RS22, SLS20, SL22c, SSL22, TSW23, WBBM23, WZC23, XCS20, YS21, YYLL20, YLSL20, ZWZ24]. Sparsity [ZZL24]. Spatial [BSH21, BMSL22, CSX22, DTQ23, HBD24, HLS^{+}21, HR24, JVRS23, KS24, KC22, LSZ24, LZCH21, MMR24, Mor21, MNBCCC24, PSBB21, SW21a, Sch20, WGK24, YZL22, ZMNVF22, ZWKZ22, ZSW22]. Spatial-Functional [LZCH21]. Spatial-Temporal [PSBB21]. Spatially [KHK21, LHGY23, PLL23, SBD23]. Spatio [CBK^{+}22, GLPD23, HBP^{+}20, LYL22, LRZ24, SLG23]. Spatio-Temporal [CBK^{+}22, GLPD23, HBP^{+}20, LYL22, LRZ24, SLG23]. Spatiotemporal [GJK^{+}20, YZL22]. SPC [Cho22]. Special [AE21, Ano21d, KLSZ21]. Specific [GZM21, LTY21, SHO20, SZZ24, WLGC24, WIS20]. Spectral [BF24, DL23b, HTF23, JT22, LL23, MP24, YZ22, ZX24]. Sphere [SW21a]. Spherical [SLC21]. Spike [BMA^{+}22, MNHG24, NR23]. Spike-and-Slab [BMA^{+}22, MNHG24, NR23]. Spiked [KML23]. Spikes [FFHL22]. Spillover [VB23]. Spillovers [CDF20]. Spline [YZZ23, YSLW22]. Splitting [DLXL23a]. Springer [Qiu24, WFL22]. Spurious [LO21]. Squared [DH22]. Squares [GC22, SBY24]. Stable [FLSU20, MOA24, Bor24]. Stage [IJM21]. Staggered [ST21]. Stan [Har23]. Standard [CJM^{+}20b, HS22]. State [LWY^{+}21, WCZZ24, XCZW23]. State-Space [LWY^{+}21, XCZW23]. States [WHHZ22]. Statin [GWL^{+}23]. Stationary [DP21, LHGY23]. Statistic [ISSM23, SL20]. Statistical [Ano24, Beb20, CGM23, CHKW24, CLS21a, CLS21b, CF23, Che23, Coo20, DLS24, FGW23, GV21, HMB^{+}20, JSK^{+}20, Jew21, Li20, LA20, LYL22, MQLH20, Muk24, NCMK20, Sen24, SSLL21, WZC23, Zha22, Che20d, Ell23, Len20, Rah24, Rob24]. Statistically [SLL^{+}24]. Statistics [AM21, BWSB24, CCLS21, Hor22, Kaf20, LLL22a, Sch20, ZZ22, ZZH22, Seo24, Chi21]. Stefan [Efr22]. Steffen [All20]. Step [GMVFX22, XX23]. Stick [GLM23]. Stick-Breaking [GLM23]. Stochastic [Beh20, BST24, CLS21b, HH23, HQYZ20, HZQ^{+}21, LL23, LYL22, NF21, OPI22, RSvdL21, SW21b, TBC20, Vat20, ZCW23]. Strategies [ARN20, Beb20, SM21, WW22, WZW20]. Stratified [WSM^{+}23, WWL23]. Streaming [LZS23]. Streams [SCACvdH24]. Street [GJK^{+}20]. Stress [GLLZ22, XTK^{+}22]. Stringent [TTL^{+}21]. Strong [GCW20, ST23b]. Structural [BLLY22, BMSL22, MCLT24, NCMK20, SS22, YSL24, ZCZZ23]. Structure [AS21, CZG24, CP23, EM23, GSW23, HSB21, JHB20, LWWftADNI21, LHS22, MML23, XYWQ24]. Structure-Adaptive [GSW23]. Structured [CLZ20, DOH22, GX23, MNHG24, PND23, TTPD20, ZB22]. Structures [AL23, AZ23, CCH20, CHDP23, LPR24, LX20, LC20, ZK23, ZWX^{+}23a, ZWX^{+}23b]. Stuart [Won24]. Studentized [Fog20]. Studies [CCLS21, EB21, Fog20, FAM21, FDF20, GSS22a, LSD21, LLL22a, LDL^{+}22, LSB^{+}22, LLL22b, ST21, SWS22, Ste23, TZL20, WCL21, XSP23, ZSL^{+}23, ZZ22]. Study [BWSB24, DGD21, FLKK21, KSR21, LNB22, NLS^{+}21, RLYI20, RLS23, SL20, SHR^{+}20a, ZDNH23, ZQL^{+}23]. Studying [SJW23]. Subcortical [AL23, CHDP23, ZK23, ZWX^{+}23a, ZWX^{+}23b]. Subgraph [CKY22]. Subgroup [GWL^{+}23, JHB20, LTY21, XYWQ24]. Subgroup-Specific [LTY21]. Subgroups [GH21]. Subject [LHW23, MLL21, MNW^{+}20, QZ22]. Subject-level [MNW^{+}20]. Subpopulation [SHO20]. Subpopulation-Specific [SHO20]. Subpopulations [LBN23]. Subsampling [KG23, YWAZ22]. Subset [CWF^{+}24]. Subspace [SBJ20, TF23a, ZMZ24]. Sufficient [ZMZ24, McD20]. Sulfate [WHHZ22]. Sum [HL24]. Summary [BWSB24, CCLS21, XSP23, ZZ22, ZZH22]. Sundberg [Che20d]. Supervised [ABS^{+}22, NCL22, SLZ24, Wan23, HM23]. Supervision [YRTJ23]. Supply [ZLQ^{+}21]. Supply-Demand [ZLQ^{+}21]. Support [GCRZ24, WQ23]. Surface [MP23, MYB^{+}20, ZLD23]. Surface-to-Volume [ZLD23]. Surfaces [LA20]. Surgeons [PK20]. Surgical [FLKK21]. Surrogate [DL22a]. Surveillance [RZCL22]. Survey [BWSB24, CLW20, KRW24, SHWB23, YQL23]. Survival [GKB^{+}21, HMZZ23, Kan21, LTY21, RPLG22, SMN^{+}20, SCMW24, THXZ23, WLGC24]. SVD [KW24]. Switching [BCI24]. Switzerland [Qiu24]. Symmetric [FSV20, YDV24]. Symmetrized [DGSZ23]. Symmetry [GPPV20]. Symptom [GYY^{+}22]. Syndrome [GYY^{+}22]. Synthesis [MANW20]. Synthesize [LLgX22]. Synthetic [AE21, AL21, BMFR21, CFT21, CSdG^{+}23, CWZ21, Fer21, FNDK23, KMPT21, Li20, MKZ^{+}21]. Systems [CSV21, IAPK23, LBF^{+}24, STC^{+}20, ZLQ^{+}21].

T [Sha21, Zha22, Jew21, WJIS20]. Tables [PZ21b]. Tadesse [Ni23]. Tail [CCH20, HW21, ZX24]. Tailed [BCT24, TJC24]. Tale [KCH23a, KCH23b, Nie23, SF23, Yon23]. Tales [Gio23]. Targeted [JRSW22, Kow22, MNW^{+}20, MD20, SZ21]. Targeting [VB24]. Tarn [Wan20]. Task [MKftADNI24, Wan23]. Tasks [BZS22]. Taxonomic [PBOD23]. Taylor [Har23]. Tchetgen [ZP21]. Techniques [Gio23]. Tempered [QW24]. Template [MNW^{+}20]. Temporal [CBK^{+}22, GLPD23, HBP^{+}20, LYL22, LRZ24, PSBB21, SLG23, TL23, YZ22]. Temporary [MA23]. Tensor [BCI24, CLX23, CYZ22a, DTQ23, HSZ24, IS23a, LT22, MZPD22, OY22, Peñ22a, TBQ20, WZLL22, ZSZL23]. Term [LKM21]. Terminal [WNK24]. Test [CX23, FDNZ24, FGRS23, HZQ^{+}21, HL24, LZ21a, LJL23b, LHGY23, LX20, LLL22b, LYZL24, MZW23, PRV22, SLS20, XLWG23, XZLG24, YLXL23, YLX24, ZGZC20, ZZ23a]. Testing [BCG20, CSX22, CLSG24, CDF20, DSL22, DS23, DHL20, FAS^{+}22, FGJS22, GSW23, HLS^{+}21, JCY^{+}21, KY20, KRW24, LHS22, Li22, LXZZ23, LLYM21, LDL^{+}22, LSB^{+}22, MCL21a, SWR24, SL22a, SWL^{+}23, SLSH20, TMZ22, XDM24, YLPC20, YLXL23, ZC22]. Tests [BCCP22, BC23, GPPV20, GVW20, GHL24, HHH23, JWW^{+}24, KL21a, LSP20, LZZ23b, LZH24, MKV^{+}24, MAZB22, PND23, ST21, SDH22, SL20, Wes22, WD21, Yee22]. Text [HM23, Mar21, XZWT21]. textbook [Hor22]. Their [CLZ20, MLZH24, PLZ24]. Theorem [LRY24]. Theorems [SPMK22]. Theoretical [SWS22]. Theories [MHZ23]. Theory [Ano21d, FFHL22, HTF23, KLSZ21, LM21, SLZ24, SSL22, Beh20, Muk24, Sch20]. Theory-Guided [HTF23]. There [CSV21, Dre23]. Thinking [Ell23, Ni22]. Thirty [LBN21]. Thompson [LR23b]. Three [NLS^{+}21]. Threshold [GRZL23]. Thresholded [WGK24]. Thresholding [FGZ22, FGQ24, FGJS22, LZX^{+}24, TZ24, ZWZ22]. Thresholds [FDNZ24]. Tibshirani [Wag20]. Time [Ano23a, Ano23b, BGOT20, BCT24, BKR^{+}23, BCI24, BTHK20, CTC20, CYZ22a, CWW22, CS24, Coo20, DGD21, DP21, DFH^{+}21, DJM23, DM22, GT22, GT23, GMVFX22, HARF^{+}20, HCZY24, JKL^{+}23, JAO23, KSW20, KCPW21, LRS20, LSN23, LT22, LZS23, MR20, MCLT24, Mie23, OY22, PS23, Peñ22a, QZ22, SLSH20, SHR^{+}20b, WZLL22, WRG23, WNK24, Wei20, ZSL20, ZPYZ24, ZX24, ZZH22, ZWSZ23, Che20a]. Time-Dependent [CS24]. Time-to-Event [WRG23, ZZH22, ZWSZ23]. Time-Varying [Ano23b, BCI24, DGD21, DM22, LSN23, MCLT24, WNK24, ZSL20]. Times [LHW23]. Tipping [Yee22]. Tobit [AFDZ23]. Tone [PLCS21]. Tool [ZDY23]. Topic [CHYL23, KW24, LZQ^{+}21, WZC23]. Topological [CMC^{+}20]. Topology [HMB^{+}20]. Tracing [MP23]. Tradeoff [KMPT21, LM21]. Tradeoffs [PK20]. Trained [PK20]. Training [DL21]. Trait [HTA^{+}22]. Traits [JCY^{+}21]. Trajectories [HL21b, JJ21, LLR21a, LLR21b]. Trans [ZYZ24]. Trans-Ancestry [ZYZ24]. Transcriptome [XSP23]. Transcriptome-Wide [XSP23]. Transfer [LSN23, LCL23, LZCL24, TF23b]. Transformation [CKK24, ZWSZ23]. Transformation-Invariant [ZWSZ23]. Transformed [CSX21, ZSW22]. Transitional [LM21]. Translation [SLC21]. Transport [Dua23, GO22, KS24]. Transportation [DS23]. Trauma [GLLZ22, NLS^{+}21]. Traumatic [XTK^{+}22]. Treat [NBW21, PK23]. Treatment [CKTVB21, CZW21, CT21b, DWHS23, DLG23, FWW23, FVX20, Fog20, FAM21, GZM21, Han21, HBX23, HSC21, IL23, KPV22, Li20, LLK^{+}20, MCLT24, MQL21a, MQL21b, NLS^{+}21, PZ21a, QLFL20, QMZ24, SST21, ST21, SLS20, SMN^{+}20, SW21b, WHX20, WLGC24, WZW20, WWF23, XC23a, XZZ^{+}22, YPL^{+}21, YSWB24, ZWSZ21, Zha22, ZD24, ZZQ24]. Treatment-Induced [XC23a]. Treatment-Specific [WLGC24]. Treatments [FVX20, HGC24, MHOZ23, RMM20]. Tree [HH23, LHS22, LLM23a, SW21b, TTPD20, WCL21, ZDM21, ZLZ24]. Tree-Structured [TTPD20]. Trees [AM24, HTA^{+}22, KT24, Mur21, ZLD23]. Trends [dHZ21]. Triage [LZQ^{+}21, SZS24b]. Trial [LTY21]. Trials [CSdG^{+}23, GH21, PK23, WSM^{+}23, YSYZ23, Beb20, Ell23]. Triplet [NLS^{+}21]. Trolls [CMZ20]. Truncated [TZ24, YYZZ20, ZRPB24]. Tsiatis [Zha22]. Tukey [DLP23]. Tumor [WJIS20, WIS20, ZBM^{+}24, ZDNH23]. Tumor-Specific [WIS20]. Tuning [FMW20, LS20, Loh20, WPB^{+}20a, WPB^{+}20b]. Tuning-Free [FMW20, LS20, Loh20, WPB^{+}20a, WPB^{+}20b]. Turing [Pai23]. Twitter [LW20]. Two [DSGS22, DFG^{+}24, DREM^{+}21, DWR23, EM23, GSS23a, GCRZ24, GMVFX22, HLS^{+}21, HXZ20, HL24, HSC21, IJM21, JWW^{+}24, KCH23a, KCH23b, LHW23, Nie23, SF23, TZL20, TZ24, XCS20, Yon23, YZL22, ZGZC20, ZSL^{+}23]. Two-Component [DSGS22]. Two-Dimensional [HXZ20]. Two-Layer [DWR23]. Two-Level [EM23]. Two-Part [HSC21]. Two-Phase [TZL20]. Two-Sample [GCRZ24, HL24, JWW^{+}24, XCS20, ZGZC20]. Two-Scale [DFG^{+}24]. Two-Stage [IJM21]. Two-Step [GMVFX22]. Two-Way [TZ24]. Type [Bor24, CLZ22, GWL^{+}23, LLL22b, SZZ24, XZWT21]. Types [DHHK21, LLZL23].

U.S. [CCH20, TFCF21, YLPC20]. UK [Gio23, Rob24, GKB^{+}21]. UK-Biobank [GKB^{+}21]. Ultra [LLZL23, XLYT20, ZLGY23]. Ultra-High [LLZL23, XLYT20, ZLGY23]. Ultrahigh [LWWftADNI21, NCL22]. Ultrahigh-Dimensional [LWWftADNI21, NCL22]. Umbrella [YRTJ23]. Unbalanced [ZYZ24]. Uncertain [CP23]. Uncertainty [HBP^{+}20, ZX24]. Unconscious [XZWT21]. Underpinnings [ZCZZ23]. Understanding [DM21, DY21, FYY23, QDL21a, QDL21b]. Undirected [GR21, WBK^{+}22]. Unemployment [YNCK21]. Unified [KRW24, LLPY24, MLZH24]. Uniform [ST23b, WTW20]. Uniformity [WSX22]. Uniformly [WWF23]. United [WHHZ22]. University [Bor24, Che20d, Gio23, Par22, Rob24, Sen24]. Unknown [WRG23]. Unmeasured [BK22, MV20, MHOZ23, QMZ24, ZWSZ21]. Unobserved [AB20, LH22, MN22]. Unseen [CFMB24]. Update [AHDP24]. Urologic [GYY^{+}22]. US-Trained [PK20]. USA [Har23]. Usage [GWL^{+}23]. Use [FBZ21, SHWB23]. User [BMDG20]. Using [DS23, FLLZ22, FGY24, FNDK23, GLLT21, GJK^{+}20, GZM21, HZB^{+}24, HQCC21, HZQ^{+}21, HL24, JCY^{+}21, JHB20, KW24, KCPW21, KRW24, Kow22, LSD21, LD20, LLL22a, LZCH21, LLgX22, LWSL22, LLK^{+}20, MW20, MNW^{+}20, MKZ^{+}21, PLSZ23, PSBB21, PLL23, QCS^{+}21a, QCS^{+}21b, RBMB21, SKL22, SSB21, SL22a, SZS24b, SL20, SZ21, TTPD20, VB23, WLGC24, WZW20, XZL23, YPL^{+}21, YNCK21, ZNBR22, NMSD21, SHWB23]. Utility [LX24]. Utilizing [XYWQ24].

V [LLK^{+}20]. V-Learning [LLK^{+}20]. Vaccine [HS22]. Valid [GHL24, MMR24]. Validated [BPvdL20, LDL^{+}22]. Validation [BHT24, Lei20, RR22, Wag20, YPL^{+}21]. Value [GLLT21, HSZ24, LX20, SQPQ21, dHZ21, dHZ24, TDB24]. Value-at-Risk [SQPQ21]. Valued [WQ23, ZQL^{+}23, ZSK23]. Values [Hof22b]. Vannucci [Ni23]. VAR [SS22]. Variable [ARN20, CFS23, CT21b, FSV20, FLKK21, GDP22, HMW23, LR23b, MCLT24, Ni23, QCS^{+}21a, QCS^{+}21b, SR20, TXQ21, TF23a, TWM23, WCNL24, WGSC23, XLYT20, YSL24, ZMZ24, ZWLS20, ZLZ23a]. Variables [AB23, BP23, GLM23, LLYM21, LLZL23, TFCF21, VB23, Har23]. Variance [Ano22c, LM21, RT20a, RT20b, SH20]. Variant [SL20]. Variants [HZB^{+}24]. Variate [CTC20, CF23]. Variation [CFMB24, FBZ21, MBH22a, MBH22b]. Variational [RS22, SBK20, TBC20, YB24]. Various [GLLT21, QLFL20]. VARMA [HLL22]. Varying [Ano23b, BCI24, DGD21, DM22, LSN23, MCLT24, PLL23, WNK24, ZSL20, ZCM22]. Vector [BSM23, LYZL24, WZLL22, WQ23, WBBM23]. Vectors [JWW^{+}24, YLXL23]. Vehicles [GJK^{+}20]. Verbal [FBZ21]. Versatile [ZM22]. Very [DPY22]. Vexler [Zha20]. Via [ISSM23, JM20, JRSW22, LLM23b, LR23b, ZW23, ZZXL20, ACADF23, BK22, Bor24, BIK20, CLS21b, CRYZ23, CWF^{+}24, DL22a, DLXL23a, DG23, Dua23, FBZ21, HJO^{+}20, HVL23, Hof22b, JWW^{+}24, KR23, Kal21a, LXZZ23, LOW21, PT23, PBF22, QW21, QW24, RZCL22, SSLL21, SDH22, THXZ23, TF23a, WZLL22, WRG23, WGK24, XX23, YNCK21, ZDM21, ZZQ24]. View [Wan23, GJK^{+}20]. Viewpoint [BMD^{+}24]. Violence [PSBB21]. Violent [PSBB21]. Visits [SMMH21]. Visualization [LLYM21, Par22]. Vitality [Ban22, TJK22, ZXS^{+}22a]. Volatility [ASTZ24]. Volume [ZLD23]. Vulnerable [GWL^{+}23].

W [Ni22]. Wager [Hof22a, Efr22, Gho22, Imb22]. Wainer [Par22]. Wainwright [All20]. Wald [Yee22]. Wand [WFL22, JHB20]. Warp [BK20, WJM22]. Wasserstein [CLM23, WCL21]. Wavelet [CvS21, GMVFX22, MZW23]. Wavelet-Based [GMVFX22, MZW23]. Way [HL21a, TZ24]. Weak [FL22, LHGY23, WD21]. Weather [HHLT21, SZS24b]. Website [JPR20]. Weighted [HL24]. Weighting [CSX22, DG23, LD20, MW20, Won24]. Weights [HVL23, HGC24, KRW24]. Well [BHT24]. Wesley [Ni22]. When-to-Treat [NBW21]. While [KSB23]. Whole [LLL22a, LLL22b]. Wide [CCLS21, LRZ24, SL20, XSP23, ZC22, ZZ22, LSB^{+}22]. Wiley [Ism20]. Wilson [Rah24]. Wise [HZQ^{+}21, XQ21]. Without [CDF20, FDF20, LM21, YLSL20, ZM22]. Worker [LNB22]. World [CT21a, LM21]. Wu [Ber24].

x [AS20, GHL24]. Xenografts [RLC^{+}21]. xi [WFL22, Won24]. xii [Gio23, Muk24, Ye24]. xiii [Sch20, Seo24]. Xin [Jan23, ZM23]. Xing [Jan23, XC23b, ZM23]. Xiong [Ché20b]. xiv [Beb20, Cha23, Che20d, Len20, Sen24]. xix [HM23, Ni22]. Xu [Har23]. xv [Ber24, Che20a, Har23, Zha22]. xvi [Rob24]. xvii [Cho22, Ism20]. xviii [All20, Zha20]. xx [Bor24]. xxi [McD20, Rah24, Wan20]. xxiii [Ni23, Sto22]. xxix [Beh20, Ché20b]. xxvi [LN20]. xxxiv [Qiu24].

Yang [LT22]. Year [AHDP24]. Years [GV21, LBN21]. Yi [Cha23]. Yield [PLL23]. Yoichi [Seo24]. York [AS20, Bor24, Che20d, Sen24, WFL22]. Young [ZCZZ23]. Yu [Zha20].

Zebrafish [TJH^{+}23]. Zero [JZW23, LM21, MP24, SPMC21, ZPZW23]. Zero-Inflated [JZW23, SPMC21, ZPZW23]. Zhang [LT22]. Zing [CJZ23]. Zubizarreta [Won24].

References

Ando:2020:QCM

[AB20] Tomohiro Ando and Jushan Bai. Quantile Co-movement in financial markets: a panel quantile model with unobserved heterogeneity. Journal of the American Statistical Association, 115(529):266–279, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Arellano:2023:RLV

[AB23] Manuel Arellano and Stéphane Bonhomme. Recovering latent variables by matching. Journal of the American Statistical Association, 118(541):693–706, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Athey:2021:MCM

[ABD^{+}21] Susan Athey, Mohsen Bayati, Nikolay Doudchenko, Guido Imbens, and Khashayar Khosravi. Matrix completion methods for causal panel data models. Journal of the American Statistical Association, 116(536):1716–1730, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Azriel:2022:SSL

[ABS^{+}22] David Azriel, Lawrence D. Brown, Michael Sklar, Richard Berk, Andreas Buja, and Linda Zhao. Semi-supervised linear regression. Journal of the American Statistical Association, 117(540):2238–2251, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Alquier:2023:ECM

[ACADF23] Pierre Alquier, Badr-Eddine Chérief-Abdellatif, Alexis Derumigny, and Jean-David Fermanian. Estimation of copulas via maximum mean discrepancy. Journal of the American Statistical Association, 118(543):1997–2012, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Argiento:2020:HNC

[ACV20] Raffaele Argiento, Andrea Cremaschi, and Marina Vannucci. Hierarchical normalized completely random measures to cluster grouped data. Journal of the American Statistical Association, 115(529):318–333, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Abadie:2021:ISS

[AE21] Alberto Abadie and Matias D. Cattaneo (Guest Editors). Introduction to the special section on synthetic control methods. Journal of the American Statistical Association, 116 (536):1713–1715, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anceschi:2023:BCP

[AFDZ23] Niccolò Anceschi, Augusto Fasano, Daniele Durante, and Giacomo Zanella. Bayesian conjugacy in probit, tobit, multinomial probit and extensions: a review and new results. Journal of the American Statistical Association, 118(542):1451–1469, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Agresti:2023:BRC

[Agr23] Alan Agresti. Book review: Confidence Intervals for Discrete Data in Clinical Research. Journal of the American Statistical Association, 118(544):2945, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Aleshin-Guendel:2023:MPR

[AGS23] Serge Aleshin-Guendel and Mauricio Sadinle. Multifile partitioning for record linkage and duplicate detection. Journal of the American Statistical Association, 118(543):1786–1795, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Almendra-Hernandez:2024:MBY

[AHDP24] Félix Almendra-Hernández, Jesús A. De Loera, and Sonja Petrovi’c. Markov bases: a 25 year update. Journal of the American Statistical Association, 119(546):1671–1686, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Abadie:2021:PSC

[AL21] Alberto Abadie and Jérémy L’Hour. A penalized synthetic control estimator for disaggregated data. Journal of the American Statistical Association, 116(536):1817–1834, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Aston:2023:DLL

[AL23] John A. D. Aston and Eardi Lila. Discussion of LESA: Longitudinal elastic shape analysis of brain subcortical structures.

Journal of the American Statistical Association, 118(541):18– 19, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [ZWX^{+}23a].

Allen:2020:BRH

[All20] Genevera I. Allen. Book review: Handbook of Graphical Models, by Marloes Maathuis, Mathias Drton, Steffen Lauritzen, and Martin Wainwright, eds. Boca Raton, FL: Chapman & Hall/CRC Press, 2018, xviii + 536 pp., $140.00(H), ISBN: 978-1-49-878862-5. Journal of the American Statistical Association, 115(531):1555–1557, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Avella-Medina:2021:PPP

[AM21] Marco Avella-Medina. Privacy-preserving parametric inference: a case for robust statistics. Journal of the American Statistical Association, 116(534):969–983, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Allen:2022:G

[AM22] David E. Allen and Michael McAleer. “Generalized measures of correlation for asymmetry, nonlinearity, and beyond”: Some antecedents on causality. Journal of the American Statistical Association, 117(537):214–224, 2022. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Awaya:2024:HMP

[AM24] Naoki Awaya and Li Ma. Hidden Markov Pólya trees for highdimensional distributions. Journal of the American Statistical Association, 119(545):189–201, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2020:C

[Ano20a] Anonymous. Correction. Journal of the American Statistical Association, 115(530):1035–1036, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2020:EC

[Ano20b] Anonymous. Editorial collaborators. Journal of the American Statistical Association, 115(532):2105–2113, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2021:Ca

[Ano21a] Anonymous. Correction. Journal of the American Statistical Association, 116(534):1039, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2021:Cb

[Ano21b] Anonymous. Correction. Journal of the American Statistical Association, 116(535):1560, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2021:Cc

[Ano21c] Anonymous. Correction. Journal of the American Statistical Association, 116(536):2100, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2021:ITM

[Ano21d] Anonymous. Introduction to Theory and Methods Special Issue on Precision Medicine and Individualized Policy Discovery Part II. Journal of the American Statistical Association, 116(534):645, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2022:C

[Ano22a] Anonymous. Correction. Journal of the American Statistical Association, 117(538):1043, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2022:CSI

[Ano22b] Anonymous. Correction to: Semiparametric Inference for Non-monotone Missing-Not-at-Random Data: the No Self-Censoring Model. Journal of the American Statistical Association, 117(537):530, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [MST21].

Anonymous:2022:FRR

[Ano22c] Anonymous. From fixed-X to random-X regression: Biasvariance decompositions, covariance penalties, and prediction error estimation: Correction. Journal of the American Statistical Association, 117(537):529, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [RT20b].

Anonymous:2023:CBP

[Ano23a] Anonymous. Correction: Bootstrap prediction bands for functional time series. Journal of the American Statistical Association, 118(543):2211, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [PS23].

Anonymous:2023:CMT

[Ano23b] Anonymous. Correction to “Modeling Time-Varying Random Objects and Dynamic Networks”. Journal of the American Statistical Association, 118(541):778, 2023. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [DM22].

Anonymous:2024:JAS

[Ano24] Anonymous. The Journal of the American Statistical Association 2023 Associate Editors. Journal of the American Statistical Association, 119(545):792–793, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Adhikari:2020:NBI

[ARN20] Samrachana Adhikari, Sherri Rose, and Sharon-Lise Normand. Nonparametric Bayesian instrumental variable analysis: Evaluating heterogeneous effects of coronary arterial access site strategies. Journal of the American Statistical Association, 115(532):1635–1644, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Aronow:2020:BRB

[AS20] Peter M. Aronow and Fredrik Sävje. Book review: The Book of Why: The New Science of Cause and Effect, Judea Pearl and Dana Mackenzie. New York: Basic Books, 2018, x + 418 pp., $32.00 (H), ISBN: 978-0-46-509760-9. Journal of the American Statistical Association, 115(529):482–485, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Awan:2021:SSD

[AS21] Jordan Awan and Aleksandra Slavković. Structure and sensitivity in differential privacy: Comparing K-norm mechanisms. Journal of the American Statistical Association, 116 (534):935–954, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Abadie:2022:RPM

[AS22] Alberto Abadie and Jann Spiess. Robust post-matching inference. Journal of the American Statistical Association, 117 (538):983–995, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Agarwal:2021:RPC

[ASSS21] Anish Agarwal, Devavrat Shah, Dennis Shen, and Dogyoon Song. On robustness of principal component regression. Journal of the American Statistical Association, 116(536):1731– 1745, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Andersen:2024:IPV

[ASTZ24] Torben G. Andersen, Tao Su, Viktor Todorov, and Zhiyuan Zhang. Intraday periodic volatility curves. Journal of the American Statistical Association, 119(546):1181–1191, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Avarucci:2023:REL

[AZ23] Marco Avarucci and Paolo Zaffaroni. Robust estimation of large panels with factor structures. Journal of the American Statistical Association, 118(544):2394–2405, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Banerjee:2022:DMH

[Ban22] Sudipto Banerjee. Discussion of “Measuring Housing Vitality from Multi-Source Big Data and Machine Learning”. Journal of the American Statistical Association, 117(539):1063–1065, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [ZXS^{+}22a] and rejoinder [ZXS^{+}22b]..

Bu:2022:LBI

[BAXV22] Fan Bu, Allison E. Aiello, Jason Xu, and Alexander Volfovsky. Likelihood-based inference for partially observed epidemics on dynamic networks. Journal of the American Statistical Association, 117(537):510–526, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bertanha:2023:PTN

[BC23] Marinho Bertanha and Eunyi Chung. Permutation tests at nonparametric rates. Journal of the American Statistical Association, 118(544):2833–2846, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Barabesi:2022:CTB

[BCCP22] Lucio Barabesi, Andrea Cerasa, Andrea Cerioli, and Domenico Perrotta. On characterizations and tests of Benford’s Law. Journal of the American Statistical Association, 117(540):1887–1903, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bradic:2020:FET

[BCG20] Jelena Bradic, Gerda Claeskens, and Thomas Gueuning. Fixed effects testing in high-dimensional linear mixed models. Journal of the American Statistical Association, 115(532): 1835–1850, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Billio:2024:BMS

[BCI24] Monica Billio, Roberto Casarin, and Matteo Iacopini. Bayesian Markov-switching tensor regression for time-varying networks. Journal of the American Statistical Association, 119(545):109–121, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bates:2021:MKS

[BCJW21] Stephen Bates, Emmanuel Candès, Lucas Janson, and Wenshuo Wang. Metropolized knockoff sampling. Journal of the American Statistical Association, 116(535):1413–1427, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Barigozzi:2024:IHT

[BCT24] Matteo Barigozzi, Giuseppe Cavaliere, and Lorenzo Trapani. Inference in heavy-tailed nonstationary multivariate time series. Journal of the American Statistical Association, 119 (545):565–581, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Balocchi:2023:CPB

[BDGJ23] Cecilia Balocchi, Sameer K. Deshpande, Edward I. George, and Shane T. Jensen. Crime in Philadelphia: Bayesian clustering with particle optimization. Journal of the American Statistical Association, 118(542):818–829, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bebu:2020:BRI

[Beb20] Ionut Bebu. Book review: Innovative Strategies, Statistical Solutions and Simulations for Modern Clinical Trials, by Mark Chang, John Balser, Jim Roach, and Robin Bliss. Boca Raton, FL: Chapman & Hall/CRC Press, 2019, xiv + 361 pp., $119.95(H), ISBN: 978-0-81-537944-7. Journal of the American Statistical Association, 115(530):1029–1030, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Behme:2020:BRT

[Beh20] Anita D. Behme. Book review: Theory of Stochastic Objects: Probability, Stochastic Processes and Inference, by Athanasios Christou Micheas. Boca Raton, FL: Chapman & Hall/ CRC Press, 2018, xxix + 378 pp., $99.95(H), ISBN: 978-1-46-651520-8. Journal of the American Statistical Association, 115(529):486–487, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Berg:2024:BRQ

[Ber24] Arthur Berg. Book review: Quantitative Methods for Precision Medicine: Pharmacogenomics in Action, Rongling Wu, Boca Raton, FL: Chapman & Hall/CRC Press, 2023, xv + 289 pp., $110.00(H), ISBN 978-1-482-21945-6. Journal of the American Statistical Association, 119(546):1690–1691, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Barigozzi:2024:AEL

[BF24] Matteo Barigozzi and Matteo Farnè. An algebraic estimator for large spectral density matrices. Journal of the American Statistical Association, 119(545):498–510, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bi:2022:MPO

[BFLZ22] Xuan Bi, Long Feng, Cai Li, and Heping Zhang. Modeling pregnancy outcomes through sequentially nested regression models. Journal of the American Statistical Association, 117(538):602–616, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bacro:2020:HST

[BGOT20] Jean-Noël Bacro, Carlo Gaetan, Thomas Opitz, and Gwladys Toulemonde. Hierarchical space-time modeling of asymptotically independent exceedances with an application to precipitation data. Journal of the American Statistical Association, 115(530):555–569, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bates:2024:CVW

[BHT24] Stephen Bates, Trevor Hastie, and Robert Tibshirani. Crossvalidation: What does it estimate and how well does it do it? Journal of the American Statistical Association, 119(546): 1434–1445, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bradley:2020:BHM

[BHW20] Jonathan R. Bradley, Scott H. Holan, and Christopher K. Wikle. Bayesian hierarchical models with conjugate fullconditional distributions for dependent data from the natural exponential family. Journal of the American Statistical Association, 115(532):2037–2052, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Breto:2020:PDA

[BIK20] Carles Bretó, Edward L. Ionides, and Aaron A. King. Panel data analysis via mechanistic models. Journal of the American Statistical Association, 115(531):1178–1188, 2020. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bharath:2020:DWM

[BK20] Karthik Bharath and Sebastian Kurtek. Distribution on warp maps for alignment of open and closed curves. Journal of the American Statistical Association, 115(531):1378–1392, 2020.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bonvini:2022:SAP

[BK22] Matteo Bonvini and Edward H. Kennedy. Sensitivity analysis via the proportion of unmeasured confounding. Journal of the American Statistical Association, 117(539):1540–1550, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Berger:2023:MPM

[BKR^{+}23] Moritz Berger, Ana Kowark, Rolf Rossaint, Mark Coburn, Matthias Schmid, and the Group Pose-Study. Modeling postoperative mortality in older patients by boosting discrete-time competing risks models. Journal of the American Statistical Association, 118(544):2239–2249, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bao:2022:CSL

[BLLY22] Le Bao, Changcheng Li, Runze Li, and Songshan Yang. Causal structural learning on MPHIA individual dataset. Journal of the American Statistical Association, 117(540): 1642–1655, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bai:2022:SSG

[BMA^{+}22] Ray Bai, Gemma E. Moran, Joseph L. Antonelli, Yong Chen, and Mary R. Boland. Spike-and-slab group lassos for grouped regression and sparse generalized additive models. Journal of the American Statistical Association, 117(537):184–197, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bhatia:2024:BRN

[BMD^{+}24] Kush Bhatia, Yi-An Ma, Anca D. Dragan, Peter L. Bartlett, and Michael I. Jordan. Bayesian robustness: a nonasymptotic viewpoint. Journal of the American Statistical Association, 119(546):1112–1123, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Banerjee:2020:LSC

[BMDG20] Trambak Banerjee, Gourab Mukherjee, Shantanu Dutta, and Pulak Ghosh. A large-scale constrained joint modeling approach for predicting user activity, engagement, and churn with application to freemium mobile games. Journal of the American Statistical Association, 115(530):538–554, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ben-Michael:2021:ASC

[BMFR21] Eli Ben-Michael, Avi Feller, and Jesse Rothstein. The augmented synthetic control method. Journal of the American Statistical Association, 116(536):1789–1803, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Banerjee:2020:ASE

[BMS20] Trambak Banerjee, Gourab Mukherjee, and Wenguang Sun. Adaptive sparse estimation with side information. Journal of the American Statistical Association, 115(532):2053–2067, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Brown:2022:BSB

[BMSL22] D. Andrew Brown, Christopher S. McMahan, Russell T. Shinohara, and Kristin A. Linn. Bayesian spatial binary regression for label fusion in structural neuroimaging. Journal of the American Statistical Association, 117(538):547–560, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). For the Alzheimer’s Disease Neuroimaging Initiative.

Bai:2021:MCC

[BN21] Jushan Bai and Serena Ng. Matrix completion, counterfactuals, and factor analysis of missing data. Journal of the American Statistical Association, 116(536):1746–1763, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bormetti:2024:BRS

[Bor24] Giacomo Bormetti. Book review: Stable Lévy Processes via Lamperti-Type Representations, Andreas E. Kyprianou and Juan Carlos Pardo, New York, NY: Cambridge University Press, 2022, xx + 463 pp., $69.99(H), ISBN 978-1-108-48029-1. Journal of the American Statistical Association, 119(545): 789–790, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Blackwell:2023:NIV

[BP23] Matthew Blackwell and Nicole E. Pashley. Noncompliance and instrumental variables for 2^{K} factorial experiments. Journal of the American Statistical Association, 118(542):1102– 1114, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Benkeser:2020:ISS

[BPvdL20] David Benkeser, Maya Petersen, and Mark J. van der Laan. Improved small-sample estimation of nonlinear crossvalidated prediction metrics. Journal of the American Statistical Association, 115(532):1917–1932, 2020. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bai:2022:IEM

[BRS22] Yuehao Bai, Joseph P. Romano, and Azeem M. Shaikh. Inference in experiments with matched pairs. Journal of the American Statistical Association, 117(540):1726–1737, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bopp:2021:HMI

[BSH21] Gregory P. Bopp, Benjamin A. Shaby, and Raphaël Huser. A hierarchical max-infinitely divisible spatial model for extreme precipitation. Journal of the American Statistical Association, 116(533):93–106, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bai:2023:MCP

[BSM23] Peiliang Bai, Abolfazl Safikhani, and George Michailidis. Multiple change point detection in reduced rank high dimensional vector autoregressive models. Journal of the American Statistical Association, 118(544):2776–2792, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bilodeau:2024:SCR

[BST24] Blair Bilodeau, Alex Stringer, and Yanbo Tang. Stochastic convergence rates and applications of adaptive quadrature in Bayesian inference. Journal of the American Statistical Association, 119(545):690–700, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

BenTaieb:2021:HPF

[BTH21] Souhaib Ben Taieb, James W. Taylor, and Rob J. Hyndman. Hierarchical probabilistic forecasting of electricity demand with smart meter data. Journal of the American Statistical Association, 116(533):27–43, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bruce:2020:EFB

[BTHK20] Scott A. Bruce, Cheng Yong Tang, Martica H. Hall, and Robert T. Krafty. Empirical frequency band analysis of nonstationary time series. Journal of the American Statistical Association, 115(532):1933–1945, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Barrientos:2024:FSD

[BWSB24] Andrés F. Barrientos, Aaron R. Williams, Joshua Snoke, and Claire McKay Bowen. A feasibility study of differentially private summary statistics and regression analyses with evaluations on administrative and survey data. Journal of the American Statistical Association, 119(545):52–65, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Betancourt:2022:RPM

[BZS22] Brenda Betancourt, Giacomo Zanella, and Rebecca C. Steorts. Random partition models for microclustering tasks. Journal of the American Statistical Association, 117(539): 1215–1227, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Christiansen:2022:TCI

[CBK^{+}22] Rune Christiansen, Matthias Baumann, Tobias Kuemmerle, Miguel D. Mahecha, and Jonas Peters. Toward causal inference for spatio-temporal data: Conflict and forest loss in Colombia. Journal of the American Statistical Association, 117(538):591–601, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cai:2023:IGL

[CCgX23] Chencheng Cai, Rong Chen, and Min ge Xie. Individualized group learning. Journal of the American Statistical Association, 118(541):622–638, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Castro-Camilo:2020:LLE

[CCH20] Daniela Castro-Camilo and Raphaël Huser. Local likelihood estimation of complex tail dependence structures, applied to U.S. precipitation extremes. Journal of the American Statistical Association, 115(531):1037–1054, 2020. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cerovecki:2023:MPS

[CCH23a] Clément Cerovecki, Vaidotas Characiejus, and Siegfried Hörmann. The maximum of the periodogram of a sequence of functional data. Journal of the American Statistical Association, 118(544):2712–2720, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2023:GSF

[CCH23b] Yu-Ting Chen, Jeng-Min Chiou, and Tzee-Ming Huang. Greedy segmentation for a functional data sequence. Journal of the American Statistical Association, 118(542):959– 971, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2021:PRF

[CCLS21] Ting-Huei Chen, Nilanjan Chatterjee, Maria Teresa Landi, and Jianxin Shi. A penalized regression framework for building polygenic risk models based on summary statistics from genome-wide association studies and incorporating external information. Journal of the American Statistical Association, 116(533):133–143, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Corradi:2020:TJS

[CDF20] Valentina Corradi, Walter Distaso, and Marcelo Fernandes. Testing for jump spillovers without testing for jumps. Journal of the American Statistical Association, 115(531):1214–1226, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Crucinio:2023:PMS

[CDJ23] Francesca R. Crucinio, Arnaud Doucet, and Adam M. Johansen. A particle method for solving Fredholm equations of the first kind. Journal of the American Statistical Association, 118(542):937–947, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2023:SIH

[CF23] Elynn Y. Chen and Jianqing Fan. Statistical inference for high-dimensional matrix-variate factor models. Journal of the American Statistical Association, 118(542):1038–1055, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Camerlenghi:2024:SPP

[CFMB24] Federico Camerlenghi, Stefano Favaro, Lorenzo Masoero, and Tamara Broderick. Scaled process priors for Bayesian nonparametric estimation of the unseen genetic variation. Journal of the American Statistical Association, 119(545):320– 331, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chevallier:2023:RJP

[CFS23] Augustin Chevallier, Paul Fearnhead, and Matthew Sutton. Reversible jump PDMP samplers for variable selection. Journal of the American Statistical Association, 118(544):2915– 2927, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cattaneo:2021:PIS

[CFT21] Matias D. Cattaneo, Yingjie Feng, and Rocio Titiunik. Prediction intervals for synthetic control methods. Journal of the American Statistical Association, 116(536):1865–1880, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chernozhukov:2020:GIQ

[CFVMW20] Victor Chernozhukov, Iván Fernández-Val, Blaise Melly, and Kaspar Wüthrich. Generic inference on quantile and quantile effect functions for discrete outcomes. Journal of the American Statistical Association, 115(529):123–137, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2023:CNO

[CFWY23] Yuxin Chen, Jianqing Fan, Bingyan Wang, and Yuling Yan. Convex and nonconvex optimization are both minimaxoptimal for noisy blind deconvolution under random designs.

Journal of the American Statistical Association, 118(542): 858–868, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cai:2023:SIH

[CGM23] T. Tony Cai, Zijian Guo, and Rong Ma. Statistical inference for high-dimensional generalized linear models with binary outcomes. Journal of the American Statistical Association, 118(542):1319–1332, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chatterjee:2021:NCC

[Cha21] Sourav Chatterjee. A new coefficient of correlation. Journal of the American Statistical Association, 116(536):2009–2022, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chan:2023:BRH

[Cha23] Kwun Chuen Gary Chan. Book review: Handbook of Measurement Error Models, Grace Y. Yi, Aurore Delaigle, and Paul Gustafson, eds. Boca Raton, FL: Chapman & Hall/ CRC Press, 2022, xiv + 577 pp., $270.00(H), ISBN: 978-1-138-10640-6. Journal of the American Statistical Association, 118(541):776–777, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chung:2023:DLL

[CHDP23] Moo K. Chung, Jamie L. Hanson, Richard J. Davidson, and Seth D. Pollak. Discussion of “LESA: Longitudinal Elastic Shape Analysis of Brain Subcortical Structures”. Journal of the American Statistical Association, 118(541):20–21, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [ZWX^{+}23a].

Chen:2020:BRT

[Che20a] Ming Chen. Book review: Time Series Clustering and Classification, by Elizabeth Ann Maharaj, Pierpaolo D’Urso, and Jorge Caiado. Boca Raton, FL: Chapman & Hall/CRC Press, 2019, xv + 228 pp., $174.95(H), ISBN: 978-1-49-877321-8. Journal of the American Statistical Association, 115(531): 1558, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2020:BRB

[Ché20b] Oliver Y. Chén. Book review: Big Data in Omics and Imaging: Integrated Analysis and Causal Inference, by Momiao Xiong. Boca Raton, FL: Chapman & Hall/CRC Press, 2018, xxix + 736 pp., $129.95 (H), ISBN: 978-0-81-538710-7. Journal of the American Statistical Association, 115(529):487– 488, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2020:BRH

[Che20c] Yen-Chi Chen. Book review: Handbook of Mixture Analysis. Journal of the American Statistical Association, 115(532): 2101–2102, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2020:BRS

[Che20d] Yen-Chi Chen. Book review: Statistical Modelling by Exponential Families, by Rolf Sundberg. New York, NY: Cambridge University Press, 2019, xiv + 282 pp., $39.99(P), ISBN: 978-1-10-870111-2. Journal of the American Statistical Association, 115(530):1032, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2023:SIL

[Che23] Yen-Chi Chen. Statistical inference with local optima. Journal of the American Statistical Association, 118(543):1940– 1952, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chiu:2021:BRH

[Chi21] Grace S. Chiu. Book review: Handbook of Environmental and Ecological Statistics. Journal of the American Statistical Association, 116(533):453–455, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chang:2024:SIC

[CHKW24] Jinyuan Chang, Jing He, Jian Kang, and Mingcong Wu. Statistical inferences for complex dependence of multimodal imaging data. Journal of the American Statistical Association, 119(546):1486–1499, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Choe:2021:DEG

[Cho21] Youngjun Choe. Design of experiments for generalized linear models. Journal of the American Statistical Association, 116 (534):1038, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Choe:2022:BRI

[Cho22] Youngjun Choe. Book review: An Introduction to Acceptance Sampling and SPC with R, John Lawson, Boca Raton, FL, Chapman & Hall/CRC Press, 2021, xvii + 280 pp., $150.00 (H), $59.95 (e-book), ISBN 978-0-36-756995-2 (H), 978-1-00-310027-0 (e-book). Journal of the American Statistical Association, 117(537):528, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2023:LTM

[CHYL23] Yinyin Chen, Shishuang He, Yun Yang, and Feng Liang. Learning topic models: Identifiability and finite-sample analysis. Journal of the American Statistical Association, 118 (544):2860–2875, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cho:2024:PLA

[CJK^{+}24] Sungho Cho, Jeong Min Jeon, Dongwoo Kim, Kyusang Yu, and Byeong U. Park. Partially linear additive regression with a general Hilbertian response. Journal of the American Statistical Association, 119(546):942–956, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cattaneo:2020:SLP

[CJM20a] Matias D. Cattaneo, Michael Jansson, and Xinwei Ma. Simple local polynomial density estimators. Journal of the American Statistical Association, 115(531):1449–1455, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chang:2020:BLC

[CJM^{+}20b] Changgee Chang, Jeong Hoon Jang, Amita Manatunga, Andrew T. Taylor, and Qi Long. A Bayesian latent class model to predict kidney obstruction in the absence of gold standard. Journal of the American Statistical Association, 115 (532):1645–1663, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Claeskens:2023:DSF

[CJZ23] Gerda Claeskens, Maarten Jansen, and Jing Zhou. Discussion on: “A Scale-Free Approach for False Discovery Rate Control in Generalized Linear Models” by Dai, Lin, Zing, Liu. Journal of the American Statistical Association, 118(543):1573–1577, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [DLXL23c].

Carlan:2024:BCT

[CKK24] Manuel Carlan, Thomas Kneib, and Nadja Klein. Bayesian conditional transformation models. Journal of the American Statistical Association, 119(546):1360–1373, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chiang:2023:IHD

[CKS23] Harold D. Chiang, Kengo Kato, and Yuya Sasaki. Inference for high-dimensional exchangeable arrays. Journal of the American Statistical Association, 118(543):1595–1605, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cattaneo:2021:ETE

[CKTVB21] Matias D. Cattaneo, Luke Keele, Rocío Titiunik, and Gonzalo Vazquez-Bare. Extrapolating treatment effects in multi-cutoff regression discontinuity designs. Journal of the American Statistical Association, 116(536):1941–1952, 2021. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chang:2022:ESD

[CKY22] Jinyuan Chang, Eric D. Kolaczyk, and Qiwei Yao. Estimation of subgraph densities in noisy networks. Journal of the American Statistical Association, 117(537):361–374, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2022:DEP

[CLLY22] Xi Chen, Jason D. Lee, He Li, and Yun Yang. Distributed estimation for principal component analysis: an enlarged eigenspace analysis. Journal of the American Statistical Association, 117(540):1775–1786, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2023:WR

[CLM23] Yaqing Chen, Zhenhua Lin, and Hans-Georg Müller. Wasserstein regression. Journal of the American Statistical Association, 118(542):869–882, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2020:DFP

[CLS20] Xi Chen, Qihang Lin, and Bodhisattva Sen. On degrees of freedom of projection estimators with applications to multivariate nonparametric regression. Journal of the American Statistical Association, 115(529):173–186, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2021:SIOa

[CLS21a] Haoyu Chen, Wenbin Lu, and Rui Song. Statistical inference for online decision making: In a contextual bandit setting. Journal of the American Statistical Association, 116(533): 240–255, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2021:SIOb

[CLS21b] Haoyu Chen, Wenbin Lu, and Rui Song. Statistical inference for online decision making via stochastic gradient descent. Journal of the American Statistical Association, 116 (534):708–719, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2024:LTC

[CLSG24] Haoyu Chen, Wenbin Lu, Rui Song, and Pulak Ghosh. On learning and testing of counterfactual fairness through data preprocessing. Journal of the American Statistical Association, 119(546):1286–1296, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2020:DRI

[CLW20] Yilin Chen, Pengfei Li, and Changbao Wu. Doubly robust inference with nonprobability survey samples. Journal of the American Statistical Association, 115(532):2011–2021, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cai:2022:IDP

[CLX22] Tianxi Cai, Molei Liu, and Yin Xia. Individual data protected integrative regression analysis of high-dimensional heterogeneous data. Journal of the American Statistical Association, 117(540):2105–2119, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cai:2023:GLR

[CLX23] Jian-Feng Cai, Jingyang Li, and Dong Xia. Generalized lowrank plus sparse tensor estimation by fast Riemannian optimization. Journal of the American Statistical Association, 118 (544):2588–2604, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2020:SLF

[CLZ20] Yunxiao Chen, Xiaoou Li, and Siliang Zhang. Structured latent factor analysis for large-scale data: Identifiability, estimability, and their implications. Journal of the American Statistical Association, 115(532):1756–1770, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2022:FON

[CLZ22] Xi Chen, Weidong Liu, and Yichen Zhang. First-order Newton-type estimator for distributed estimation and inference. Journal of the American Statistical Association, 117 (540):1858–1874, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Crawford:2020:PCO

[CMC^{+}20] Lorin Crawford, Anthea Monod, Andrew X. Chen, Sayan Mukherjee, and Raúl Rabadán. Predicting clinical outcomes in glioblastoma: an application of topological and functional data analysis. Journal of the American Statistical Association, 115(531):1139–1150, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2024:HEI

[CMW24] Zhehui Chen, Simon Mak, and C. F. Jeff Wu. A hierarchical expected improvement method for Bayesian optimization. Journal of the American Statistical Association, 119 (546):1619–1632, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2020:FTU

[CMZ20] Dachuan Chen, Per A. Mykland, and Lan Zhang. The five trolls under the bridge: Principal component analysis with asynchronous and noisy high frequency data. Journal of the American Statistical Association, 115(532):1960–1977, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cook:2020:SAM

[Coo20] Richard J. Cook. The statistical analysis of multivariate failure time data: a marginal modeling approach. Journal of the American Statistical Association, 115(532):2102–2104, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cox:2020:DPB

[Cox20] D. R. Cox. Discussion of paper by Brad Efron. Journal of the American Statistical Association, 115(530):659, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Efr20a].

Castelletti:2023:NSL

[CP23] Federico Castelletti and Stefano Peluso. Network structure learning under uncertain interventions. Journal of the American Statistical Association, 118(543):2117–2128, 2023. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ciarleglio:2022:EAS

[CPH22] Adam Ciarleglio, Eva Petkova, and Ofer Harel. Elucidating age and sex-dependent association between frontal EEG asymmetry and depression: an application of multiple imputation in functional regression. Journal of the American Statistical Association, 117(537):12–26, 2022. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cui:2024:SPC

[CPS^{+}24] Yifan Cui, Hongming Pu, Xu Shi, Wang Miao, and Eric Tchetgen Tchetgen. Semiparametric proximal causal inference. Journal of the American Statistical Association, 119(546): 1348–1359, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cressie:2020:CWI

[Cre20] Noel Cressie. Comment: When is it data science and when is it data engineering? Journal of the American Statistical Association, 115(530):660–662, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cressie:2022:NEB

[Cre22] Noel Cressie. Nonparametric empirical Bayes prediction. Journal of the American Statistical Association, 117(539): 1167–1170, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2023:DDS

[CRYZ23] Hui Chen, Haojie Ren, Fang Yao, and Changliang Zou. Datadriven selection of the number of change-points via error rate control. Journal of the American Statistical Association, 118 (542):1415–1428, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Candes:2020:DPP

[CS20] Emmanuel Candès and Chiara Sabatti. Discussion of the paper “Prediction, Estimation, and Attribution” by B. Efron. Journal of the American Statistical Association, 115(530): 656–658, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Efr20a].

Chen:2023:LCC

[CS23] Jie Chen and Michael L. Stein. Linear-cost covariance functions for Gaussian random fields. Journal of the American Statistical Association, 118(541):147–164, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chu:2024:CIR

[CS24] Chi Wing Chu and Tony Sit. Censored interquantile regression model with time-dependent covariates. Journal of the American Statistical Association, 119(546):1592–1603, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chandra:2023:BNC

[CSdG^{+}23] Noirrit Kiran Chandra, Abhra Sarkar, John F. de Groot, Ying Yuan, and Peter Müller. Bayesian nonparametric common atoms regression for generating synthetic controls in clinical trials. Journal of the American Statistical Association, 118 (544):2301–2314, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chan:2021:MSE

[CSV21] Lax Chan, Bernard W. Silverman, and Kyle Vincent. Multiple systems estimation for sparse capture data: Inferential challenges when there are nonoverlapping lists. Journal of the American Statistical Association, 116(535):1297–1306, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chu:2021:TDQ

[CSX21] Chi Wing Chu, Tony Sit, and Gongjun Xu. Transformed dynamic quantile regression on censored data. Journal of the American Statistical Association, 116(534):874–886, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cai:2022:LLA

[CSX22] T. Tony Cai, Wenguang Sun, and Yin Xia. LAWS: a locally adaptive weighting and screening approach to spatial multiple testing. Journal of the American Statistical Association, 117 (539):1370–1383, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cui:2021:MII

[CT21a] Yifan Cui and Eric Tchetgen Tchetgen. Machine intelligence for individualized decision making under a counterfactual world: a rejoinder. Journal of the American Statistical Association, 116(533):200–206, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See rejoinder [QCS^{+}21b].

Cui:2021:SIV

[CT21b] Yifan Cui and Eric Tchetgen Tchetgen. A semiparametric instrumental variable approach to optimal treatment regimes under endogeneity. Journal of the American Statistical Association, 116(533):162–173, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2020:CFM

[CTC20] Elynn Y. Chen, Ruey S. Tsay, and Rong Chen. Constrained factor models for high-dimensional matrix-variate time series. Journal of the American Statistical Association, 115(530): 775–793, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chau:2021:IWR

[CvS21] Joris Chau and Rainer von Sachs. Intrinsic wavelet regression for curves of Hermitian positive definite matrices. Journal of the American Statistical Association, 116(534):819–832, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2024:NDI

[CWF^{+}24] Yi Chen, Yining Wang, Ethan X. Fang, Zhaoran Wang, and Runze Li. Nearly dimension-independent sparse linear bandit over small action spaces via best subset selection. Journal of the American Statistical Association, 119(545):246–258, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2024:IHD

[CWS24] Yudong Chen, Tengyao Wang, and Richard J. Samworth. Inference in high-dimensional online changepoint detection. Journal of the American Statistical Association, 119(546): 1461–1472, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2022:IBH

[CWW22] Likai Chen, Weining Wang, and Wei Biao Wu. Inference of breakpoints in high-dimensional time series. Journal of the American Statistical Association, 117(540):1951–1963, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chernozhukov:2021:ERC

[CWZ21] Victor Chernozhukov, Kaspar Wüthrich, and Yinchu Zhu. An exact and robust conformal inference method for counterfactual and synthetic controls. Journal of the American Statistical Association, 116(536):1849–1864, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2023:NTH

[CX23] Hao Chen and Yin Xia. A normality test for high-dimensional data based on the nearest neighbor approach. Journal of the American Statistical Association, 118(541):719–731, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cai:2023:OPG

[CXH23] Zhibo Cai, Yingcun Xia, and Weiqiang Hang. An outerproduct-of-gradient approach to dimension reduction and its application to classification in high dimensional space. Journal of the American Statistical Association, 118(543):1671– 1681, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2022:FMH

[CYZ22a] Rong Chen, Dan Yang, and Cun-Hui Zhang. Factor models for high-dimensional tensor time series. Journal of the American Statistical Association, 117(537):94–116, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2022:R

[CYZ22b] Rong Chen, Dan Yang, and Cun-Hui Zhang. Rejoinder. Journal of the American Statistical Association, 117(537):128– 132, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cai:2024:LNS

[CZG24] Biao Cai, Jingfei Zhang, and Yongtao Guan. Latent network structure learning from high-dimensional multivariate point processes. Journal of the American Statistical Association, 119(545):95–108, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2021:LIT

[CZW21] Yuan Chen, Donglin Zeng, and Yuanjia Wang. Learning individualized treatment rules for multiple-domain latent outcomes. Journal of the American Statistical Association, 116 (533):269–282, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Duarte:2022:ODE

[DAGO22] Belmiro P. M. Duarte, Anthony C. Atkinson, José F. O. Granjo, and Nuno M. C. Oliveira. Optimal design of experiments for implicit models. Journal of the American Statistical Association, 117(539):1424–1437, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Davison:2020:D

[Dav20] A. C. Davison. Discussion. Journal of the American Statistical Association, 115(530):663–664, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Denti:2023:CAM

[DCGM23] Francesco Denti, Federico Camerlenghi, Michele Guindani, and Antonietta Mira. A common atoms model for the Bayesian nonparametric analysis of nested data. Journal of the American Statistical Association, 118(541):405–416, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

DeBoeck:2022:PFA

[DDX22] Paul De Boeck, Michael L. DeKay, and Menglin Xu. The potential of factor analysis for replication, generalization, and integration. Journal of the American Statistical Association, 117(540):1622–1626, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Demirkaya:2024:ONI

[DFG^{+}24] Emre Demirkaya, Yingying Fan, Lan Gao, Jinchi Lv, Patrick Vossler, and Jingbo Wang. Optimal nonparametric inference with two-scale distributional nearest neighbors. Journal of the American Statistical Association, 119(545):297–307, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Davis:2021:CTS

[DFH^{+}21] Richard A. Davis, Konstantinos Fokianos, Scott H. Holan, Harry Joe, James Livsey, Robert Lund, Vladas Pipiras, and Nalini Ravishanker. Count time series: a methodological review. Journal of the American Statistical Association, 116 (535):1533–1547, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dette:2020:LRA

[DG20] Holger Dette and Josua Gösmann. A likelihood ratio approach to sequential change point detection for a general class of parameters. Journal of the American Statistical Association, 115(531):1361–1377, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dorn:2023:SSA

[DG23] Jacob Dorn and Kevin Guo. Sharp sensitivity analysis for inverse propensity weighting via quantile balancing. Journal of the American Statistical Association, 118(544):2645–2657, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Das:2021:MMT

[DGD21] Kiranmoy Das, Pulak Ghosh, and Michael J. Daniels. Modeling multiple time-varying related groups: a dynamic hierarchical Bayesian approach with an application to the health and retirement study. Journal of the American Statistical Association, 116(534):558–568, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dagdoug:2023:MAE

[DGH23] Mehdi Dagdoug, Camelia Goga, and David Haziza. Modelassisted estimation through random forests in finite population sampling. Journal of the American Statistical Association, 118(542):1234–1251, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Daouia:2022:ER

[DGS22] Abdelaati Daouia, Irène Gijbels, and Gilles Stupfler. Extremile regression. Journal of the American Statistical Association, 117(539):1579–1586, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Du:2023:FDR

[DGSZ23] Lilun Du, Xu Guo, Wenguang Sun, and Changliang Zou. False discovery rate control under general dependence by symmetrized data aggregation. Journal of the American Statistical Association, 118(541):607–621, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Du:2023:HDP

[DGW23] Jin-Hong Du, Yifeng Guo, and Xueqin Wang. Highdimensional portfolio selection with cardinality constraints. Journal of the American Statistical Association, 118(542): 779–791, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Du:2022:EBM

[DH22] Lilun Du and Inchi Hu. An empirical Bayes method for chisquared data. Journal of the American Statistical Association, 117(537):334–347, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Delaigle:2021:ECF

[DHHK21] Aurore Delaigle, Peter Hall, Wei Huang, and Alois Kneip. Estimating the covariance of fragmented and other related types of functional data. Journal of the American Statistical Association, 116(535):1383–1401, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dai:2022:ISM

[DHJW22] Chenguang Dai, Jeremy Heng, Pierre E. Jacob, and Nick Whiteley. An invitation to sequential Monte Carlo samplers. Journal of the American Statistical Association, 117 (539):1587–1600, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Delaigle:2020:ECP

[DHL20] Aurore Delaigle, Wei Huang, and Shaoke Lei. Estimation of conditional prevalence from group testing data with missing covariates. Journal of the American Statistical Association, 115(529):467–480, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Deng:2023:ILP

[DHS23] Hang Deng, Qiyang Han, and Bodhisattva Sen. Inference for local parameters in convexity constrained models. Journal of the American Statistical Association, 118(544):2721–2735, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

deHaan:2021:TEV

[dHZ21] Laurens de Haan and Chen Zhou. Trends in extreme value indices. Journal of the American Statistical Association, 116 (535):1265–1279, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

deHaan:2024:BEV

[dHZ24] Laurens de Haan and Chen Zhou. Bootstrapping extreme value estimators. Journal of the American Statistical Association, 119(545):382–393, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Deshpande:2023:ODA

[DJM23] Yash Deshpande, Adel Javanmard, and Mohammad Mehrabi. Online debiasing for adaptively collected high-dimensional data with applications to time series analysis. Journal of the American Statistical Association, 118(542):1126–1139, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Das:2024:DRC

[DKJ24] Manjari Das, Edward H. Kennedy, and Nicholas P. Jewell. Doubly robust capture-recapture methods for estimating population size. Journal of the American Statistical Association, 119(546):1309–1321, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dryden:2021:RMS

[DKPW21] Ian L. Dryden, Alfred Kume, Phillip J. Paine, and Andrew T. A. Wood. Regression modeling for size-and-shape data based on a Gaussian model for landmarks. Journal of the American Statistical Association, 116(534):1011–1022, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dou:2021:TNN

[DL21] Xialiang Dou and Tengyuan Liang. Training neural networks as learning data-adaptive kernels: Provable representation and approximation benefits. Journal of the American Statistical Association, 116(535):1507–1520, 2021. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dai:2022:MCA

[DL22a] Chenguang Dai and Jun S. Liu. Monte Carlo approximation of Bayes factors via mixing with surrogate distributions. Journal of the American Statistical Association, 117(538): 765–780, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dai:2022:KOD

[DL22b] Xiaowu Dai and Lexin Li. Kernel ordinary differential equations. Journal of the American Statistical Association, 117 (540):1711–1725, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dai:2023:OKD

[DL23a] Xiaowu Dai and Lexin Li. Orthogonalized kernel debiased machine learning for multimodal data analysis. Journal of the American Statistical Association, 118(543):1796–1810, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dudek:2023:SDE

[DL23b] Anna E. Dudek and Lukasz Lenart. Spectral density estimation for nonstationary data with nonzero mean function. Journal of the American Statistical Association, 118(543): 1900–1910, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dong:2023:RDD

[DLG23] Yingying Dong, Ying-Ying Lee, and Michael Gou. Regression discontinuity designs with a continuous treatment. Journal of the American Statistical Association, 118(541):208–221, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dai:2023:KKS

[DLL23] Xiaowu Dai, Xiang Lyu, and Lexin Li. Kernel knockoffs selection for nonparametric additive models. Journal of the American Statistical Association, 118(543):2158–2170, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dai:2023:TDO

[DLP23] Xiongtao Dai and Sara Lopez-Pintado. Tukey’s depth for object data. Journal of the American Statistical Association, 118(543):1760–1772, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ding:2024:SLI

[DLS24] Yi Ding, Yingying Li, and Rui Song. Statistical learning for individualized asset allocation. Journal of the American Statistical Association, 119(545):639–649, 2024. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dai:2023:FDR

[DLXL23a] Chenguang Dai, Buyu Lin, Xin Xing, and Jun S. Liu. False discovery rate control via data splitting. Journal of the American Statistical Association, 118(544):2503–2520, 2023. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dai:2023:RSF

[DLXL23b] Chenguang Dai, Buyu Lin, Xin Xing, and Jun S. Liu. Rejoinder: A Scale-free Approach for False Discovery Rate Control in Generalized Linear Models. Journal of the American Statistical Association, 118(543):1590–1594, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [DLXL23c, ZM23, XC23b, CJZ23, LB23, Jan23, LYZ23].

Dai:2023:SFA

[DLXL23c] Chenguang Dai, Buyu Lin, Xin Xing, and Jun S. Liu. A scale-free approach for false discovery rate control in generalized linear models. Journal of the American Statistical Association, 118(543):1551–1565, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See comments and discussion [ZM23, XC23b, CJZ23, LB23, Jan23, LYZ23], and rejoinder [DLXL23b].

Datta:2021:DRM

[DM21] Jyotishka Datta and Bhramar Mukherjee. Discussion on “Regression Models for Understanding COVID-19 Epidemic Dynamics With Incomplete Data”. Journal of the American Statistical Association, 116(536):1583–1586, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [QDL21a].

Dubey:2022:MTV

[DM22] Paromita Dubey and Hans-Georg Müller. Modeling timevarying random objects and dynamic networks. Journal of the American Statistical Association, 117(540):2252–2267, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See correction [Ano23b].

deMicheaux:2021:DCD

[dMMV21] Pierre Lafaye de Micheaux, Pavlo Mozharovskyi, and Myriam Vimond. Depth for curve data and applications. Journal of the American Statistical Association, 116(536):1881–1897, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dempsey:2022:HNM

[DOH22] Walter Dempsey, Brandon Oselio, and Alfred Hero. Hierarchical network models for exchangeable structured interaction processes. Journal of the American Statistical Association, 117(540):2056–2073, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Das:2021:PIL

[DP21] Srinjoy Das and Dimitris N. Politis. Predictive inference for locally stationary time series with an application to climate data. Journal of the American Statistical Association, 116 (534):919–934, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dette:2022:ECP

[DPY22] Holger Dette, Guangming Pan, and Qing Yang. Estimating a change point in a sequence of very high-dimensional covariance matrices. Journal of the American Statistical Association, 117(537):444–454, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Drechsler:2023:DPG

[Dre23] Jörg Drechsler. Differential privacy for government agenciesare we there yet? Journal of the American Statistical Association, 118(541):761–773, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Diaz-Rodriguez:2021:NEG

[DREM^{+}21] Jairo Diaz-Rodriguez, Dominique Eckert, Hatef Monajemi, Stéphane Paltani, and Sylvain Sardy. Nonparametric estimation of galaxy cluster emissivity and detection of point sources in astrophysics with two lasso penalties. Journal of the American Statistical Association, 116(535):1088–1099, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Deb:2023:MRB

[DS23] Nabarun Deb and Bodhisattva Sen. Multivariate rank-based distribution-free nonparametric testing using measure transportation. Journal of the American Statistical Association, 118(541):192–207, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Deb:2022:TCM

[DSGS22] Nabarun Deb, Sujayam Saha, Adityanand Guntuboyina, and Bodhisattva Sen. Two-component mixture model in the presence of covariates. Journal of the American Statistical Association, 117(540):1820–1834, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dai:2022:MTP

[DSL22] James Y. Dai, Janet L. Stanford, and Michael LeBlanc. A multiple-testing procedure for high-dimensional mediation hypotheses. Journal of the American Statistical Association, 117(537):198–213, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dai:2022:EL

[DSW22a] Ben Dai, Xiaotong Shen, and Junhui Wang. Embedding learning. Journal of the American Statistical Association, 117(537):307–319, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dai:2022:CG

[DSW22b] Ben Dai, Xiaotong Shen, and Wing Wong. Coupled generation. Journal of the American Statistical Association, 117 (539):1243–1253, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dai:2021:SCR

[DSWQ21] Ben Dai, Xiaotong Shen, Junhui Wang, and Annie Qu. Scalable collaborative ranking for personalized prediction. Journal of the American Statistical Association, 116(535):1215– 1223, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Deng:2023:CTD

[DTQ23] Yujia Deng, Xiwei Tang, and Annie Qu. Correlation tensor decomposition and its application in spatial imaging data. Journal of the American Statistical Association, 118(541): 440–456, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Duan:2023:TMC

[Dua23] Leo L. Duan. Transport Monte Carlo: High-accuracy posterior approximation via random transport. Journal of the American Statistical Association, 118(543):1659–1670, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Deresa:2024:CBC

[DV24] Negera Wakgari Deresa and Ingrid Van Keilegom. Copula based Cox proportional hazards models for dependent censoring. Journal of the American Statistical Association, 119 (546):1044–1054, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Diaconis:2021:DGS

[DW21] Persi Diaconis and Guanyang Wang. Discussion of “A Gibbs Sampler for a Class of Random Convex Polytopes”. Journal of the American Statistical Association, 116(535):1193–1195, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [JGED21a, JGED21b].

Diaz:2023:NCE

[DWHS23] Iván Díaz, Nicholas Williams, Katherine L. Hoffman, and Edward J. Schenck. Nonparametric causal effects based on longitudinal modified treatment policies. Journal of the American Statistical Association, 118(542):846–857, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dunn:2023:DFP

[DWR23] Robin Dunn, Larry Wasserman, and Aaditya Ramdas. Distribution-free prediction sets for two-layer hierarchical models. Journal of the American Statistical Association, 118 (544):2491–2502, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dean:2021:DRM

[DY21] Natalie Dean and Yang Yang. Discussion of “Regression Models for Understanding COVID-19 Epidemic Dynamics With Incomplete Data”. Journal of the American Statistical Association, 116(536):1587–1590, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [QDL21a].

Deng:2023:QAA

[DYFQ23] Yujia Deng, Yubai Yuan, Haoda Fu, and Annie Qu. Queryaugmented active metric learning. Journal of the American Statistical Association, 118(543):1862–1875, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Eckles:2021:BHD

[EB21] Dean Eckles and Eytan Bakshy. Bias and high-dimensional adjustment in observational studies of peer effects. Journal of the American Statistical Association, 116(534):507–517, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Efron:2020:PEA

[Efr20a] Bradley Efron. Prediction, estimation, and attribution. Journal of the American Statistical Association, 115(530):636– 655, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See discussion [CS20, Cox20, FHT20, gXZ20] and rejoinder [Efr20b].

Efron:2020:R

[Efr20b] Bradley Efron. Rejoinder. Journal of the American Statistical Association, 115(530):675–677, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Efr20a].

Efron:2022:DCI

[Efr22] Bradley Efron. Discussion of “Confidence Intervals for Nonparametric Empirical Bayes Analysis” by Nikolaos Ignatiadis and Stefan Wager. Journal of the American Statistical Association, 117(539):1179–1180, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [IW22a] and rejoinder [IW22b].

Ellenberg:2023:BRS

[Ell23] Susan S. Ellenberg. Book review: Statistical Thinking in Clinical Trials Michael A. Proschan, Boca Raton, FL: Chapman & Hall/CRC Press, 2022, ix + 253 pp., $99.95(H), ISBN 978-1-138-05859-0. Journal of the American Statistical Association, 118(542):1448–1449, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Edwards:2023:SNT

[EM23] David J. Edwards and Robert W. Mee. Structure of nonregular two-level designs. Journal of the American Statistical Association, 118(542):1222–1233, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ertefaie:2021:RQL

[EMOS21] Ashkan Ertefaie, James R. McKay, David Oslin, and Robert L. Strawderman. Robust Q-learning. Journal of the American Statistical Association, 116(533):368–381, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Forastiere:2021:IET

[FAM21] Laura Forastiere, Edoardo M. Airoldi, and Fabrizia Mealli. Identification and estimation of treatment and interference effects in observational studies on networks. Journal of the American Statistical Association, 116(534):901–918, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Feng:2022:TBS

[FAS^{+}22] Derek Feng, Randolf Altmeyer, Derek Stafford, Nicholas A. Christakis, and Harrison H. Zhou. Testing for balance in social networks. Journal of the American Statistical Association, 117(537):156–174, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Frumento:2021:PMQ

[FBFV21] Paolo Frumento, Matteo Bottai, and Iván Fernández-Val. Parametric modeling of quantile regression coefficient functions with longitudinal data. Journal of the American Statistical Association, 116(534):783–797, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Feng:2021:BRI

[FBZ21] Long Feng, Xuan Bi, and Heping Zhang. Brain regions identified as being associated with verbal reasoning through the use of imaging regression via internal variation. Journal of the American Statistical Association, 116(533):144–158, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ferrari:2021:BFA

[FD21] Federico Ferrari and David B. Dunson. Bayesian factor analysis for inference on interactions. Journal of the American Statistical Association, 116(535):1521–1532, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fiksel:2022:GBQ

[FDAZ22] Jacob Fiksel, Abhirup Datta, Agbessi Amouzou, and Scott Zeger. Generalized Bayes quantification learning under dataset shift. Journal of the American Statistical Association, 117(540):2163–2181, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Franks:2020:FSA

[FDF20] Alexander M. Franks, Alexander D’Amour, and Avi Feller. Flexible sensitivity analysis for observational studies without observable implications. Journal of the American Statistical Association, 115(532):1730–1746, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2020:RLS

[FDLL20] Yingying Fan, Emre Demirkaya, Gaorong Li, and Jinchi Lv. RANK: Large-scale inference with graphical nonlinear knockoffs. Journal of the American Statistical Association, 115 (529):362–379, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Feng:2024:TSH

[FDNZ24] Huijie Feng, Jingyi Duan, Yang Ning, and Jiwei Zhao. Test of significance for high-dimensional thresholds with application to individualized minimal clinically important difference. Journal of the American Statistical Association, 119(546): 1396–1408, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ferman:2021:PSC

[Fer21] Bruno Ferman. On the properties of the synthetic control estimator with many periods and many controls. Journal of the American Statistical Association, 116(536):1764–1772, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2022:ATE

[FFHL22] Jianqing Fan, Yingying Fan, Xiao Han, and Jinchi Lv. Asymptotic theory of eigenvectors for random matrices with diverging spikes. Journal of the American Statistical Association, 117(538):996–1009, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fu:2022:HAR

[FGJS22] Luella Fu, Bowen Gang, Gareth M. James, and Wenguang Sun. Heteroscedasticity-adjusted ranking and thresholding for large-scale multiple testing. Journal of the American Statistical Association, 117(538):1028–1040, 2022. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fang:2024:AFT

[FGQ24] Qin Fang, Shaojun Guo, and Xinghao Qiao. Adaptive functional thresholding for sparse covariance function estimation in high dimensions. Journal of the American Statistical Association, 119(546):1473–1485, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fernandez:2023:KLR

[FGRS23] Tamara Fernández, Arthur Gretton, David Rindt, and Dino Sejdinovic. A kernel log-rank test of independence for rightcensored data. Journal of the American Statistical Association, 118(542):925–936, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2023:CEA

[FGW23] Jianqing Fan, Yongyi Guo, and Kaizheng Wang. Communicationefficient accurate statistical estimation. Journal of the American Statistical Association, 118(542):1000–1010, 2023. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2024:POU

[FGY24] Jianqing Fan, Yongyi Guo, and Mengxin Yu. Policy optimization using semiparametric models for dynamic pricing. Journal of the American Statistical Association, 119(545): 552–564, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2022:ENF

[FGZ22] Jianqing Fan, Jianhua Guo, and Shurong Zheng. Estimating number of factors by adjusted eigenvalues thresholding. Journal of the American Statistical Association, 117(538): 852–861, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Friedman:2020:DPE

[FHT20] Jerome Friedman, Trevor Hastie, and Robert Tibshirani. Discussion of “Prediction, Estimation, and Attribution” by Bradley Efron. Journal of the American Statistical Association, 115(530):665–666, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Efr20a].

Fan:2022:LLF

[FL22] Jianqing Fan and Yuan Liao. Learning latent factors from diversified projections and its applications to over-estimated and weak factors. Journal of the American Statistical Association, 117(538):909–924, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fogarty:2021:BEH

[FLKK21] Colin B. Fogarty, Kwonsang Lee, Rachel R. Kelz, and Luke J. Keele. Biased encouragements and heterogeneous effects in an instrumental variable study of emergency general surgical outcomes. Journal of the American Statistical Association, 116(536):1625–1636, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2022:ABH

[FLLZ22] Yiwei Fan, Xiaoling Lu, Yufeng Liu, and Junlong Zhao. Angle-based hierarchical classification using exact label embedding. Journal of the American Statistical Association, 117 (538):704–717, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2020:ISI

[FLSU20] Yingying Fan, Jinchi Lv, Mahrad Sharifvaghefi, and Yoshimasa Uematsu. IPAD: Stable interpretable forecasting with knockoffs inference. Journal of the American Statistical Association, 115(532):1822–1834, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Feng:2020:BSI

[FLSZ20] Xiangnan Feng, Tengfei Li, Xinyuan Song, and Hongtu Zhu. Bayesian scalar on image regression with nonignorable nonresponse. Journal of the American Statistical Association, 115 (532):1574–1597, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2024:LFR

[FLY24] Jianqing Fan, Zhipeng Lou, and Mengxin Yu. Are latent factor regression and sparse regression adequate? Journal of the American Statistical Association, 119(546):1076–1088, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2022:DWE

[FMM22] Jianqing Fan, Ricardo Masini, and Marcelo C. Medeiros. Do we exploit all information for counterfactual analysis? benefits of factor models and idiosyncratic correction. Journal of the American Statistical Association, 117(538):574–590, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2020:CTF

[FMW20] Jianqing Fan, Cong Ma, and Kaizheng Wang. Comment on “A tuning-free robust and efficient approach to highdimensional regression”. Journal of the American Statistical Association, 115(532):1720–1725, 2020. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [WPB^{+}20b].

Frazier:2023:BIU

[FNDK23] David T. Frazier, David J. Nott, Christopher Drovandi, and Robert Kohn. Bayesian inference using synthetic likelihood: Asymptotics and adjustments. Journal of the American Statistical Association, 118(544):2821–2832, 2023. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fogarty:2020:SSA

[Fog20] Colin B. Fogarty. Studentized sensitivity analysis for the sample average treatment effect in paired observational studies. Journal of the American Statistical Association, 115(531): 1518–1530, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Franks:2020:HAB

[Fra20] Jordan J. Franks. Handbook of approximate Bayesian computation. Journal of the American Statistical Association, 115 (532):2100–2101, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fryzlewicz:2024:NSP

[Fry24] Piotr Fryzlewicz. Narrowest significance pursuit: Inference for multiple change-points in linear models. Journal of the American Statistical Association, 119(546):1633–1646, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Florens:2020:EBV

[FSV20] Jean-Pierre Florens, Léopold Simar, and Ingrid Van Keilegom. Estimation of the boundary of a variable observed with symmetric error. Journal of the American Statistical Association, 115(529):425–441, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Feng:2020:EHI

[FVX20] Qian Feng, Quang Vuong, and Haiqing Xu. Estimation of heterogeneous individual treatment effects with endogenous treatments. Journal of the American Statistical Association, 115(529):231–240, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Frigau:2022:OJP

[FWB22] Luca Frigau, Qiuyi Wu, and David Banks. Optimizing the JSM program. Journal of the American Statistical Association, 117(538):617–626, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fang:2023:FOL

[FWW23] Ethan X. Fang, Zhaoran Wang, and Lan Wang. Fairnessoriented learning for optimal individualized treatment rules. Journal of the American Statistical Association, 118(543): 1733–1746, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fasiolo:2021:FCA

[FWZ^{+}21] Matteo Fasiolo, Simon N. Wood, Margaux Zaffran, Raphaël Nedellec, and Yannig Goude. Fast calibrated additive quantile regression. Journal of the American Statistical Association, 116(535):1402–1412, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2023:UIR

[FYY23] Jianqing Fan, Zhuoran Yang, and Mengxin Yu. Understanding implicit regularization in over-parameterized single index model. Journal of the American Statistical Association, 118 (544):2315–2328, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fei:2023:IHD

[FZHL23] Zhe Fei, Qi Zheng, Hyokyoung G. Hong, and Yi Li. Inference for high-dimensional censored quantile regression. Journal of the American Statistical Association, 118(542):898–912, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guo:2023:GOB

[GB23] Kevin Guo and Guillaume Basse. The generalized oaxacablinder estimator. Journal of the American Statistical Association, 118(541):524–536, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Glennie:2021:IAM

[GBL^{+}21] R. Glennie, S. T. Buckland, R. Langrock, T. Gerrodette, L. T. Ballance, S. J. Chivers, and M. D. Scott. Incorporating animal movement into distance sampling. Journal of the American Statistical Association, 116(533):107–115, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gao:2024:SIH

[GBW24] Lucy L. Gao, Jacob Bien, and Daniela Witten. Selective inference for hierarchical clustering. Journal of the American Statistical Association, 119(545):332–342, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guo:2022:MDI

[GC22] Xiao Guo and Guang Cheng. Moderate-dimensional inferences on quadratic functionals in ordinary least squares. Journal of the American Statistical Association, 117(540):1931– 1950, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Geng:2024:LST

[GCRZ24] Haoyu Geng, Xiaolong Cui, Haojie Ren, and Changliang Zou. Large-scale two-sample comparison of support sets. Journal of the American Statistical Association, 119(546):1604–1618, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guan:2020:DSS

[GCW20] Leying Guan, Xi Chen, and Wing Hung Wong. Detecting strong signals in gene perturbation experiments: an adaptive approach with power guarantee and FDR control. Journal of the American Statistical Association, 115(532):1747–1755, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Geenens:2022:HC

[GdM22] Gery Geenens and Pierre Lafaye de Micheaux. The Hellinger correlation. Journal of the American Statistical Association, 117(538):639–653, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Garcia-Donato:2022:VSP

[GDP22] Gonzalo García-Donato and Rui Paulo. Variable selection in the presence of factors: a model selection perspective. Journal of the American Statistical Association, 117(540):1847–1857, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Grantham:2020:MBM

[GGR^{+}20] Neal S. Grantham, Yawen Guan, Brian J. Reich, Elizabeth T. Borer, and Kevin Gross. MIMIX: a Bayesian mixed-effects model for microbiome data from designed experiments. Journal of the American Statistical Association, 115(530):599– 609, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guo:2021:ISS

[GH21] Xinzhou Guo and Xuming He. Inference on selected subgroups in clinical trials. Journal of the American Statistical Association, 116(535):1498–1506, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Grunwald:2024:AVT

[GHL24] Peter Grünwald, Alexander Henzi, and Tyron Lardy. Anytime-valid tests of conditional independence under model-X. Journal of the American Statistical Association, 119(546): 1554–1565, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ghosal:2022:DCI

[Gho22] Subhashis Ghosal. Discussion of “Confidence Intervals for Nonparametric Empirical Bayes Analysis” by Ignatiadis and Wager. Journal of the American Statistical Association, 117 (539):1171–1174, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [IW22a] and rejoinder [IW22b].

Giordano:2023:BRD

[Gio23] Sabrina Giordano. Book review: Data Science Ethics: Concepts, Techniques and Cautionary Tales, David Martens, Oxford, UK: Oxford University Press, 2022, xii + 255 pp., $80.00(H), ISBN 978-0-19-284726-3. Journal of the American Statistical Association, 118(541):774–776, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guan:2020:FSS

[GJK^{+}20] Yawen Guan, Margaret C. Johnson, Matthias Katzfuss, Elizabeth Mannshardt, Kyle P. Messier, Brian J. Reich, and Joon J. Song. Fine-scale spatiotemporal air pollution analysis using mobile monitors on Google Street View vehicles. Journal of the American Statistical Association, 115(531):1111– 1124, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gu:2022:NML

[GK22] Jiaying Gu and Roger Koenker. Nonparametric maximum likelihood methods for binary response models with random coefficients. Journal of the American Statistical Association, 117(538):732–751, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gorfine:2021:MFB

[GKB^{+}21] Malka Gorfine, Nir Keret, Asaf Ben Arie, David Zucker, and Li Hsu. Marginalized frailty-based illness–death model: Application to the UK-Biobank survival data. Journal of the American Statistical Association, 116(535):1155–1167, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gerber:2021:ICC

[GLLT21] Guillaume Gerber, Yohann Le Faou, Olivier Lopez, and Michael Trupin. The impact of churn on client value in health insurance, evaluation using a random forest under various censoring mechanisms. Journal of the American Statistical Association, 116(536):2053–2064, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guo:2022:HDM

[GLLZ22] Xu Guo, Runze Li, Jingyuan Liu, and Mudong Zeng. Highdimensional mediation analysis for selecting DNA methylation loci mediating childhood trauma and cortisol stress reactivity. Journal of the American Statistical Association, 117 (539):1110–1121, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gil-Leyva:2023:SBP

[GLM23] María F. Gil-Leyva and Ramsés H. Mena. Stick-breaking processes with exchangeable length variables. Journal of the American Statistical Association, 118(541):537–550, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gasperoni:2023:SDM

[GLPD23] Francesca Gasperoni, Alessandra Luati, Lucia Paci, and Enzo D’Innocenzo. Score-driven modeling of spatio-temporal data. Journal of the American Statistical Association, 118(542): 1066–1077, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guerrier:2022:RTS

[GMVFX22] Stéphane Guerrier, Roberto Molinari, Maria-Pia Victoria-Feser, and Haotian Xu. Robust two-step wavelet-based inference for time series models. Journal of the American Statistical Association, 117(540):1996–2013, 2022. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Garxes:2022:ROT

[GO22] Valérie Garès and Jérémy Omer. Regularized optimal transport of covariates and outcomes in data recoding. Journal of the American Statistical Association, 117(537):320–333, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Garcia-Portugues:2020:OTR

[GPPV20] Eduardo García-Portugués, Davy Paindaveine, and Thomas Verdebout. On optimal tests for rotational symmetry against new classes of hyperspherical distributions. Journal of the American Statistical Association, 115(532):1873–1887, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guha:2021:BRU

[GR21] Sharmistha Guha and Abel Rodriguez. Bayesian regression with undirected network predictors with an application to brain connectome data. Journal of the American Statistical Association, 116(534):581–593, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guan:2020:BNP

[GRLB20] Qian Guan, Brian J. Reich, Eric B. Laber, and Dipankar Bandyopadhyay. Bayesian nonparametric policy search with application to periodontal recall intervals. Journal of the American Statistical Association, 115(531):1066–1078, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guo:2023:TSF

[GRZL23] Xu Guo, Haojie Ren, Changliang Zou, and Runze Li. Threshold selection in feature screening for error rate control. Journal of the American Statistical Association, 118(543):1773– 1785, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gunsilius:2023:INC

[GS23] Florian Gunsilius and Susanne Schennach. Independent nonlinear component analysis. Journal of the American Statistical Association, 118(542):1305–1318, 2023. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gabriel:2022:CBO

[GSS22a] Erin E. Gabriel, Michael C. Sachs, and Arvid Sjölander. Causal bounds for outcome-dependent sampling in observational studies. Journal of the American Statistical Association, 117(538):939–950, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Goldman:2022:GBM

[GSS22b] Jacob Vorstrup Goldman, Torben Sell, and Sumeetpal Sidhu Singh. Gradient-based Markov chain Monte Carlo for Bayesian inference with non-differentiable priors. Journal of the American Statistical Association, 117(540):2182–2193, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gabriel:2023:SNB

[GSS23a] Erin E. Gabriel, Michael C. Sachs, and Arvid Sjölander. Sharp nonparametric bounds for decomposition effects with two binary mediators. Journal of the American Statistical Association, 118(544):2446–2453, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gabriel:2023:NBC

[GSS23b] Erin E. Gabriel, Arvid Sjölander, and Michael C. Sachs. Nonparametric bounds for causal effects in imperfect randomized experiments. Journal of the American Statistical Association, 118(541):684–692, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gang:2023:SAS

[GSW23] Bowen Gang, Wenguang Sun, and Weinan Wang. Structureadaptive sequential testing for online false discovery rate control. Journal of the American Statistical Association, 118 (541):732–745, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gao:2022:MHD

[GT22] Zhaoxing Gao and Ruey S. Tsay. Modeling high-dimensional time series: a factor model with dynamically dependent factors and diverging eigenvalues. Journal of the American Statistical Association, 117(539):1398–1414, 2022. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gao:2023:DCD

[GT23] Zhaoxing Gao and Ruey S. Tsay. Divide-and-conquer: a distributed hierarchical factor approach to modeling large-scale time series data. Journal of the American Statistical Association, 118(544):2698–2711, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guggisberg:2023:BAM

[Gug23] Michael Guggisberg. A Bayesian approach to multiple-output quantile regression. Journal of the American Statistical Association, 118(544):2736–2745, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gelman:2021:WMI

[GV21] Andrew Gelman and Aki Vehtari. What are the most important statistical ideas of the past 50 years? Journal of the American Statistical Association, 116(536):2087–2097, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gerstenberger:2020:TSC

[GVW20] Carina Gerstenberger, Daniel Vogel, and Martin Wendler. Tests for scale changes based on pairwise differences. Journal of the American Statistical Association, 115(531):1336–1348, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guo:2023:AMV

[GWL^{+}23] Xinzhou Guo, Waverly Wei, Molei Liu, Tianxi Cai, Chong Wu, and Jingshen Wang. Assessing the most vulnerable subgroup to type II diabetes associated with statin usage: Evidence from electronic health record data. Journal of the American Statistical Association, 118(543):1488–1499, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gu:2023:JMA

[GX23] Yuqi Gu and Gongjun Xu. A joint MLE approach to large-scale structured latent attribute analysis. Journal of the American Statistical Association, 118(541):746–760, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xie:2020:DPB

[gXZ20] Min ge Xie and Zheshi Zheng. Discussion of Professor Bradley Efron’s article on “Prediction, Estimation, and Attribution”. Journal of the American Statistical Association, 115(530): 667–671, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Efr20a].

Guo:2022:FME

[GYY^{+}22] Wensheng Guo, Mengying You, Jialin Yi, Michel A. Pontari, and J. Richard Landis. Functional mixed effects clustering with application to longitudinal urologic chronic pelvic pain syndrome symptom data. Journal of the American Statistical Association, 117(540):1631–1641, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guo:2021:EOI

[GZM21] Wenchuan Guo, Xiao-Hua Zhou, and Shujie Ma. Estimation of optimal individualized treatment rules using a covariatespecific treatment effect curve with high-dimensional covariates. Journal of the American Statistical Association, 116 (533):309–321, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guo:2024:SIR

[GZW24] Xingche Guo, Donglin Zeng, and Yuanjia Wang. A semiparametric inverse reinforcement learning approach to characterize decision making for mental disorders. Journal of the American Statistical Association, 119(545):27–38, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Han:2021:CIT

[Han21] Sukjin Han. Comment: Individualized treatment rules under endogeneity. Journal of the American Statistical Association, 116(533):192–195, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Harris:2023:BRM

[Har23] Mark N. Harris. Book review: Modern Applied Regressions: Bayesian and Frequentist Analysis of Categorical and Limited Response variables with R and Stan, Jun Xu, Florida, USA: CRC Press (Taylor & Francis Group), 2023, xv + 281 pp., ISBN 978-0-367-17387-6. Journal of the American Statistical Association, 118(543):2209–2211, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hadj-Amar:2020:BMS

[HARF^{+}20] Beniamino Hadj-Amar, Bärbel Finkenstädt Rand, Mark Fiecas, Francis Lévi, and Robert Huckstepp. Bayesian model search for nonstationary periodic time series. Journal of the American Statistical Association, 115(531):1320–1335, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Halder:2024:BMS

[HBD24] Aritra Halder, Sudipto Banerjee, and Dipak K. Dey. Bayesian modeling with spatial curvature processes. Journal of the American Statistical Association, 119(546):1155–1167, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Heaton:2020:MBI

[HBP^{+}20] Matthew J. Heaton, Candace Berrett, Sierra Pugh, Amber Evans, and Chantel Sloan. Modeling bronchiolitis incidence proportions in the presence of spatio-temporal uncertainty. Journal of the American Statistical Association, 115(529):66– 78, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hou:2023:TEE

[HBX23] Jue Hou, Jelena Bradic, and Ronghui Xu. Treatment effect estimation under additive hazards models with highdimensional confounding. Journal of the American Statistical Association, 118(541):327–342, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Han:2024:SDM

[HCZY24] Yuefeng Han, Rong Chen, Cun-Hui Zhang, and Qiwei Yao. Simultaneous decorrelation of matrix time series. Journal of the American Statistical Association, 119(546):957–969, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

He:2024:RRU

[He24] Yi He. Ridge regression under dense factor augmented models. Journal of the American Statistical Association, 119(546): 1566–1578, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Harman:2020:REA

[HFR20] Radoslav Harman, Lenka Filová, and Peter Richtárik. A randomized exchange algorithm for computing optimal approximate designs of experiments. Journal of the American Statistical Association, 115(529):348–361, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Huling:2024:IWC

[HGC24] Jared D. Huling, Noah Greifer, and Guanhua Chen. Independence weights for causal inference with continuous treatments. Journal of the American Statistical Association, 119 (546):1657–1670, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

He:2023:STE

[HH23] Jingyu He and P. Richard Hahn. Stochastic tree ensembles for regularized nonlinear regression. Journal of the American Statistical Association, 118(541):551–570, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hallin:2023:EFD

[HHH23] Marc Hallin, Daniel Hlubinka, and Sárka Hudecová. Efficient fully distribution-free center-outward rank tests for multipleoutput regression and MANOVA. Journal of the American Statistical Association, 118(543):1923–1939, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Heinrich:2021:MPM

[HHLT21] Claudio Heinrich, Kristoffer H. Hellton, Alex Lenkoski, and Thordis L. Thorarinsdottir. Multivariate postprocessing methods for high-dimensional seasonal weather forecasts. Journal of the American Statistical Association, 116(535): 1048–1059, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hoffman:2021:CGS

[HHZ21] Kentaro Hoffman, Jan Hannig, and Kai Zhang. Comments on “A Gibbs Sampler for a Class of Random Convex Polytopes”. Journal of the American Statistical Association, 116 (535):1206–1210, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [JGED21a, JGED21b].

Hall:2020:FAB

[HJO^{+}20] P. Hall, I. M. Johnstone, J. T. Ormerod, M. P. Wand, and J. C. F. Yu. Fast and accurate binary response mixed model analysis via expectation propagation. Journal of the American Statistical Association, 115(532):1902–1916, 2020. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Henzi:2023:DSI

[HKZ23] Alexander Henzi, Gian-Reto Kleger, and Johanna F. Ziegel. Distributional (single) index models. Journal of the American Statistical Association, 118(541):489–503, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hirose:2021:MGP

[HL21a] Masayo Y. Hirose and Partha Lahiri. Multi-goal prior selection: a way to reconcile Bayesian and classical approaches for random effects models. Journal of the American Statistical Association, 116(535):1487–1497, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hu:2021:DEF

[HL21b] Jian Hu and Mingyao Li. Discussion of “Exponential-Family Embedding With Application to Cell Developmental Trajectories for Single-Cell RNA-Seq Data”. Journal of the American Statistical Association, 116(534):475–477, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [LLR21a].

Hu:2024:TSC

[HL24] Xiaoyu Hu and Jing Lei. A two-sample conditional distribution test using conformal prediction and weighted rank sum. Journal of the American Statistical Association, 119 (546):1136–1154, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hallin:2022:COE

[HLL22] M. Hallin, D. La Vecchia, and H. Liu. Center-outward Restimation for semiparametric VARMA models. Journal of the American Statistical Association, 117(538):925–938, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Harris:2021:EPI

[HLS^{+}21] Trevor Harris, Bo Li, Nathan J. Steiger, Jason E. Smerdon, Naveen Narisetty, and J. Derek Tucker. Evaluating proxy influence in assimilated paleoclimate reconstructionstesting the exchangeability of two ensembles of spatial processes. Journal of the American Statistical Association, 116 (535):1100–1113, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hung:2023:OSS

[HLW23] Ying Hung, Li-Hsiang Lin, and C. F. Jeff Wu. Optimal simulator selection. Journal of the American Statistical Association, 118(542):1264–1271, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Holub:2023:BRS

[HM23] Martin Holub and Patrícia Martinková. Book review: Supervised Machine Learning for Text Analysis in R, Emil Hvitfeldt, Julia Silge, Boca Raton, FL, CRC Press, 2022, xix + 381 pp., $64.95(P), ISBN 978-0-367-55419-4. Journal of the American Statistical Association, 118(543):2207–2209, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Henderson:2020:STR

[HMB^{+}20] Robin Henderson, Irina Makarenko, Paul Bushby, Andrew Fletcher, and Anvar Shukurov. Statistical topology and the random interstellar medium. Journal of the American Statistical Association, 115(530):625–635, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hiabu:2021:SBP

[HMMMN21] Munir Hiabu, Enno Mammen, M. Dolores Martínez-Miranda, and Jens P. Nielsen. Smooth backfitting of proportional hazards with multiplicative components. Journal of the American Statistical Association, 116(536):1983–1993, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Han:2020:AFR

[HMP20] Kyunghee Han, Hans-Georg Müller, and Byeong U. Park. Additive functional regression for densities as responses. Journal of the American Statistical Association, 115(530):997–1010, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hui:2023:GAV

[HMW23] Francis K. C. Hui, Samuel Müller, and A. H. Welsh. GEEassisted variable selection for latent variable models with multivariate binary data. Journal of the American Statistical Association, 118(542):1252–1263, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

He:2023:RBG

[HMZZ23] Baihua He, Shuangge Ma, Xinyu Zhang, and Li-Xing Zhu. Rank-based greedy model averaging for high-dimensional survival data. Journal of the American Statistical Association, 118(544):2658–2670, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hoff:2022:CPE

[Hof22a] Peter Hoff. Coverage properties of empirical Bayes intervals: a discussion of “Confidence intervals for nonparametric empirical Bayes analysis” by ignatiadis and wager. Journal of the American Statistical Association, 117(539):1175–1178, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [IW22a] and rejoinder [IW22b].

Hoff:2022:SVI

[Hof22b] Peter Hoff. Smaller p-values via indirect information. Journal of the American Statistical Association, 117(539):1254–1269, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Horton:2022:BRF

[Hor22] Nicholas J. Horton. Book review: Foundations of Statistics for Data Scientists: With R and Python Alan Agresti, and Maria Kateri, Boca Raton. FL: CRC Press, 2022, 446 pp., $99.95 (textbook), ISBN 978-0-367-74845-6. Journal of the American Statistical Association, 117(539):1603–1604, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hu:2021:PPL

[HQCC21] Xinyu Hu, Min Qian, Bin Cheng, and Ying Kuen Cheung. Personalized policy learning using longitudinal mobile health data. Journal of the American Statistical Association, 116 (533):410–420, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hu:2020:CBI

[HQYZ20] Jianwei Hu, Hong Qin, Ting Yan, and Yunpeng Zhao. Corrected Bayesian information criterion for stochastic block models. Journal of the American Statistical Association, 115 (532):1771–1783, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hector:2024:DIS

[HR24] Emily C. Hector and Brian J. Reich. Distributed inference for spatial extremes modeling in high dimensions. Journal of the American Statistical Association, 119(546):1297–1308, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hector:2021:DIM

[HS21] Emily C. Hector and Peter X.-K. Song. A distributed and integrated method of moments for high-dimensional correlated data analysis. Journal of the American Statistical Association, 116(534):805–818, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hasegawa:2022:EMV

[HS22] Raiden B. Hasegawa and Dylan S. Small. Estimating malaria vaccine efficacy in the absence of a gold standard case definition: Mendelian factorial design. Journal of the American Statistical Association, 117(539):1466–1481, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ha:2021:BSL

[HSB21] Min Jin Ha, Francesco Claudio Stingo, and Veerabhadran Baladandayuthapani. Bayesian structure learning in multilayered genomic networks. Journal of the American Statistical Association, 116(534):605–618, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Huling:2021:TPF

[HSC21] Jared D. Huling, Maureen A. Smith, and Guanhua Chen. A two-part framework for estimating individualized treatment rules from semicontinuous outcomes. Journal of the American Statistical Association, 116(533):210–223, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Han:2024:GFT

[HSZ24] Rungang Han, Pixu Shi, and Anru R. Zhang. Guaranteed functional tensor singular value decomposition. Journal of the American Statistical Association, 119(546):995–1007, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hassler:2022:IPT

[HTA^{+}22] Gabriel Hassler, Max R. Tolkoff, William L. Allen, Lam Si Tung Ho, Philippe Lemey, and Marc A. Suchard. Inferring phenotypic trait evolution on large trees with many incomplete measurements. Journal of the American Statistical Association, 117(538):678–692, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Han:2023:ESS

[HTF23] Xiao Han, Xin Tong, and Yingying Fan. Eigen selection in spectral clustering: a theory-guided practice. Journal of the American Statistical Association, 118(541):109–121, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Henderson:2023:ISD

[HVL23] Nicholas C. Henderson, Ravi Varadhan, and Thomas A. Louis. Improved small domain estimation via compromise regression weights. Journal of the American Statistical Association, 118(544):2793–2809, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hou:2021:EIT

[HW21] Yanxi Hou and Xing Wang. Extreme and inference for tail Gini functionals with applications in tail risk measurement. Journal of the American Statistical Association, 116(535): 1428–1443, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Han:2023:GPC

[HXC23] Ruijian Han, Yiming Xu, and Kani Chen. A general pairwise comparison model for extremely sparse networks. Journal of the American Statistical Association, 118(544):2422–2432, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hessellund:2022:SML

[HXGW22] Kristian Bjørn Hessellund, Ganggang Xu, Yongtao Guan, and Rasmus Waagepetersen. Semiparametric multinomial logistic regression for multivariate point pattern data. Journal of the American Statistical Association, 117(539):1500–1515, 2022.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hedayat:2020:ODT

[HXZ20] A. S. Hedayat, Heng Xu, and Wei Zheng. Optimal designs for the two-dimensional interference model. Journal of the American Statistical Association, 115(532):1812–1821, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hu:2024:DPC

[HY24] Xiaoyu Hu and Fang Yao. Dynamic principal component analysis in high dimensions. Journal of the American Statistical Association, 119(545):308–319, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hao:2021:SIF

[HyLXZ21] Meiling Hao, Kin yat Liu, Wei Xu, and Xingqiu Zhao. Semiparametric inference for the functional Cox model. Journal of the American Statistical Association, 116(535):1319–1329, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

He:2024:NIU

[HZB^{+}24] Zhibing He, Yunpeng Zhao, Peter Bickel, Charles Weko, Dan Cheng, and Jirui Wang. Network inference using the hub model and variants. Journal of the American Statistical Association, 119(546):1264–1273, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hu:2022:MMN

[HZLG22] Yingtian Hu, Mahmoud Zeydabadinezhad, Longchuan Li, and Ying Guo. A multimodal multilevel neuroimaging model for investigating brain connectome development. Journal of the American Statistical Association, 117(539):1134–1148, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hu:2021:UME

[HZQ^{+}21] Jianwei Hu, Jingfei Zhang, Hong Qin, Ting Yan, and Ji Zhu. Using maximum entry-wise deviation to test the goodness of fit for stochastic block models. Journal of the American Statistical Association, 116(535):1373–1382, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ionides:2023:BFP

[IAPK23] Edward L. Ionides, Kidus Asfaw, Joonha Park, and Aaron A. King. Bagged filters for partially observed interacting systems. Journal of the American Statistical Association, 118 (542):1078–1089, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Imai:2021:CII

[IJM21] Kosuke Imai, Zhichao Jiang, and Anup Malani. Causal inference with interference and noncompliance in two-stage randomized experiments. Journal of the American Statistical Association, 116(534):632–644, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Imai:2023:EEI

[IL23] Kosuke Imai and Michael Lingzhi Li. Experimental evaluation of individualized treatment rules. Journal of the American Statistical Association, 118(541):242–256, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ikefuji:2022:ERE

[ILMY22] Masako Ikefuji, Roger J. A. Laeven, Jan R. Magnus, and Yuan Yue. Earthquake risk embedded in property prices: Evidence from five Japanese cities. Journal of the American Statistical Association, 117(537):82–93, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Imbens:2022:CCI

[Imb22] Guido Imbens. Comment on: “Confidence Intervals for Nonparametric Empirical Bayes Analysis” by Ignatiadis and Wager. Journal of the American Statistical Association, 117 (539):1181–1182, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [IW22a] and rejoinder [IW22b].

Insua:2023:AML

[INGP23] David Rios Insua, Roi Naveiro, Víctor Gallego, and Jason Poulos. Adversarial machine learning: Bayesian perspectives. Journal of the American Statistical Association, 118 (543):2195–2206, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ibriga:2023:CAS

[IS23a] Hilda S. Ibriga and Will Wei Sun. Covariate-assisted sparse tensor completion. Journal of the American Statistical Association, 118(544):2605–2619, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Iii:2023:ECP

[IS23b] Edward Mcfowland Iii and Cosma Rohilla Shalizi. Estimating causal peer influence in homophilous social networks by inferring latent locations. Journal of the American Statistical Association, 118(541):707–718, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ismail:2020:BRM

[Ism20] Noor Azina Ismail. Book review: Measuring Agreement: Models, Methods, and Applications, by Pankaj K. Choudhary and Haikady N. Nagaraja. Hoboken, NJ: John Wiley & Sons, Inc., 2017, xvii + 336 pp., $119.50 (H), ISBN: 978-1-11-807858-7. Journal of the American Statistical Association, 115(529):485–486, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ignatiadis:2023:EBM

[ISSM23] Nikolaos Ignatiadis, Sujayam Saha, Dennis L. Sun, and Omkar Muralidharan. Empirical Bayes mean estimation with nonparametric errors via order statistic regression on replicated data. Journal of the American Statistical Association, 118(542):987–999, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ignatiadis:2022:CIN

[IW22a] Nikolaos Ignatiadis and Stefan Wager. Confidence intervals for nonparametric empirical Bayes analysis. Journal of the American Statistical Association, 117(539):1149–1166, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See discussion [Gho22, Hof22a, Efr22, Imb22, Pen22b, XS22] and rejoinder [IW22b].

Ignatiadis:2022:RCI

[IW22b] Nikolaos Ignatiadis and Stefan Wager. Rejoinder: Confidence intervals for nonparametric empirical Bayes analysis. Journal of the American Statistical Association, 117(539):1192–1199, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [IW22a] and rejoinder [IW22b].

Izenman:2021:SAD

[Ize21] Alan Julian Izenman. Sampling algorithms for discrete Markov random fields and related graphical models. Journal of the American Statistical Association, 116(536):2065–2086, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Janson:2023:DSF

[Jan23] Lucas Janson. Discussion of “A Scale-Free Approach for False Discovery Rate Control in Generalized Linear Models” by Chenguang Dai, Buyu Lin, Xin Xing, and Jun S. Liu. Journal of the American Statistical Association, 118(543):1584–1585, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [DLXL23c].

Jiao:2023:FTS

[JAO23] Shuhao Jiao, Alexander Aue, and Hernando Ombao. Functional time series prediction under partial observation of the future curve. Journal of the American Statistical Association, 118(541):315–326, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jiang:2021:GBA

[JCY^{+}21] Yingda Jiang, Chi-Yang Chiu, Qi Yan, Wei Chen, Michael B. Gorin, Yvette P. Conley, M’Hamed Lajmi Lakhal-Chaieb, Richard J. Cook, Christopher I. Amos, Alexander F. Wilson, Joan E. Bailey-Wilson, Francis J. McMahon, Ana I. Vazquez, Ao Yuan, Xiaogang Zhong, Momiao Xiong, Daniel E. Weeks, and Ruzong Fan. Gene-based association testing of dichotomous traits with generalized functional linear mixed models using extended pedigrees: Applications to age-related macular degeneration. Journal of the American Statistical Association, 116(534):531–545, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jiang:2020:FCQ

[JCYS20] Fei Jiang, Qing Cheng, Guosheng Yin, and Haipeng Shen. Functional censored quantile regression. Journal of the American Statistical Association, 115(530):931–944, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jewell:2021:SMC

[Jew21] Nicholas P. Jewell. Statistical models for COVID-19 incidence, cumulative prevalence, and R t. Journal of the American Statistical Association, 116(536):1578–1582, 2021. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jacob:2021:GSC

[JGED21a] Pierre E. Jacob, Ruobin Gong, Paul T. Edlefsen, and Arthur P. Dempster. A Gibbs sampler for a class of random convex polytopes. Journal of the American Statistical Association, 116(535):1181–1192, 2021. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). See discussion [DW21, Wil21, LV21, Rug21, HHZ21] and rejoinder [JGED21b].

Jacob:2021:RGS

[JGED21b] Pierre E. Jacob, Ruobin Gong, Paul T. Edlefsen, and Arthur P. Dempster. Rejoinder — A Gibbs Sampler for a Class of Random Convex Polytopes. Journal of the American Statistical Association, 116(535):1211–1214, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Johnson:2020:RSS

[JHB20] S. R. Johnson, D. A. Henderson, and R. J. Boys. Revealing subgroup structure in ranked data using a Bayesian WAND. Journal of the American Statistical Association, 115(532): 1888–1901, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ji:2021:DEF

[JJ21] Zhicheng Ji and Hongkai Ji. Discussion of “Exponential-Family Embedding With Application to Cell Developmental Trajectories for Single-Cell RNA-seq Data”. Journal of the American Statistical Association, 116(534):471–474, 2021.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [LLR21a].

Jia:2023:LGC

[JKL^{+}23] Yisu Jia, Stefanos Kechagias, James Livsey, Robert Lund, and Vladas Pipiras. Latent Gaussian count time series. Journal of the American Statistical Association, 118(541):596– 606, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jin:2023:OEN

[JKLW23] Jiashun Jin, Zheng Tracy Ke, Shengming Luo, and Minzhe Wang. Optimal estimation of the number of network communities. Journal of the American Statistical Association, 118 (543):2101–2116, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jacob:2020:RAS

[JLS20a] P. E. Jacob, F. Lindsten, and T. B. Schön. Retracted Article: Smoothing with couplings of conditional particle filters. Journal of the American Statistical Association, 115 (529):489, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jacob:2020:SCC

[JLS20b] Pierre E. Jacob, Fredrik Lindsten, and Thomas B. Schön. Smoothing with couplings of conditional particle filters. Journal of the American Statistical Association, 115(530):721– 729, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Javadi:2020:NMF

[JM20] Hamid Javadi and Andrea Montanari. Nonnegative matrix factorization via archetypal analysis. Journal of the American Statistical Association, 115(530):896–907, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jochmans:2022:HRI

[Joc22] Koen Jochmans. Heteroscedasticity-robust inference in linear regression models with many covariates. Journal of the American Statistical Association, 117(538):887–896, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

James:2020:PCO

[JPR20] Gareth M. James, Courtney Paulson, and Paat Rusmevichientong. Penalized and constrained optimization: an application to high-dimensional Website advertising. Journal of the American Statistical Association, 115(529):107–122, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

James:2022:IEC

[JRR22] Gareth M. James, Peter Radchenko, and Bradley Rava. Irrational exuberance: Correcting bias in probability estimates. Journal of the American Statistical Association, 117(537): 455–468, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jiang:2022:BII

[JRSW22] Bei Jiang, Adrian E. Raftery, Russell J. Steele, and Naisyin Wang. Balancing inferential integrity and disclosure risk via model targeted masking and multiple imputation. Journal of the American Statistical Association, 117(537):52–66, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jana:2020:SFR

[JSK^{+}20] Kaushik Jana, Debasis Sengupta, Subrata Kundu, Arindam Chakraborty, and Purnima Shaw. The statistical face of a region under monsoon rainfall in Eastern India. Journal of the American Statistical Association, 115(532):1559–1573, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jordanger:2022:NSA

[JT22] Lars Arne Jordanger and Dag Tjøstheim. Nonlinear spectral analysis: a local Gaussian approach. Journal of the American Statistical Association, 117(538):1010–1027, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jiang:2023:SAS

[JVRS23] Chaonan Jiang, Davide La Vecchia, Elvezio Ronchetti, and Olivier Scaillet. Saddlepoint approximations for spatial panel data models. Journal of the American Statistical Association, 118(542):1164–1175, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jiang:2024:NTS

[JWW^{+}24] Yunlu Jiang, Xueqin Wang, Canhong Wen, Yukang Jiang, and Heping Zhang. Nonparametric two-sample tests of high dimensional mean vectors via random integration. Journal of the American Statistical Association, 119(545):701–714, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jiang:2023:FZI

[JZW23] Roulan Jiang, Xiang Zhan, and Tianying Wang. A flexible zero-inflated Poisson-gamma model with application to microbiome sequence count data. Journal of the American Statistical Association, 118(542):792–804, 2023. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kafadar:2020:RIS

[Kaf20] Karen Kafadar. Reinforcing the impact of statistics on society. Journal of the American Statistical Association, 115 (530):491–500, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kallus:2021:MEP

[Kal21a] Nathan Kallus. More efficient policy learning via optimal retargeting. Journal of the American Statistical Association, 116(534):646–658, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See discussion [VD21, LLL21, LZ21b] and rejoinder [Kal21b].

Kallus:2021:RNO

[Kal21b] Nathan Kallus. Rejoinder: New objectives for policy learning. Journal of the American Statistical Association, 116(534): 694–698, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Kal21a, VD21, LLL21, LZ21b].

Kang:2021:ASM

[Kan21] Sangwook Kang. Advanced survival models. Journal of the American Statistical Association, 116(536):2098–2099, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kong:2020:LLR

[KAZZ20] Dehan Kong, Baiguo An, Jingwen Zhang, and Hongtu Zhu. L2RM: Low-rank linear regression models for highdimensional matrix responses. Journal of the American Statistical Association, 115(529):403–424, 2020. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Khan:2022:RSR

[KC22] Kori Khan and Catherine A. Calder. Restricted spatial regression methods: Implications for inference. Journal of the American Statistical Association, 117(537):482–494, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Krivitsky:2023:RDT

[KCH23a] Pavel N. Krivitsky, Pietro Coletti, and Niel Hens. Rejoinder to discussion of “A Tale of Two Datasets: Representativeness and Generalisability of Inference for Samples of Networks”. Journal of the American Statistical Association, 118 (544):2235–2238, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [KCH23b, SF23, Yon23, Nie23, KCH23a, KCH23a].

Krivitsky:2023:TTD

[KCH23b] Pavel N. Krivitsky, Pietro Coletti, and Niel Hens. A tale of two datasets: Representativeness and generalisability of inference for samples of networks. Journal of the American Statistical Association, 118(544):2213–2224, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See correction [KCH24], discussion [SF23, Yon23, Nie23], and rejoinder [KCH23a].

Krivitsky:2024:C

[KCH24] Pavel N. Krivitsky, Pietro Coletti, and Niel Hens. Correction. Journal of the American Statistical Association, 119 (546):1694–1695, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [KCH23b].

Khan:2021:CFR

[KCPW21] Md Kamrul Hasan Khan, Avishek Chakraborty, Giovanni Petris, and Barry T. Wilson. Constrained functional regression of national forest inventory data over time using remote sensing observations. Journal of the American Statistical Association, 116(535):1168–1180, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Keret:2023:ABE

[KG23] Nir Keret and Malka Gorfine. Analyzing big EHR dataoptimal Cox regression subsampling procedure with rare events. Journal of the American Statistical Association, 118 (544):2262–2275, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kuenzer:2021:PCA

[KHK21] Thomas Kuenzer, Siegfried Hörmann, and Piotr Kokoszka. Principal component analysis of spatially indexed functions. Journal of the American Statistical Association, 116(535): 1444–1456, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kramlinger:2023:MCM

[KKS23] Peter Kramlinger, Tatyana Krivobokova, and Stefan Sperlich. Marginal and conditional multiple inference for linear mixed model predictors. Journal of the American Statistical Association, 118(544):2344–2355, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Khim:2021:PTI

[KL21a] Justin Khim and Po-Ling Loh. Permutation tests for infection graphs. Journal of the American Statistical Association, 116(534):770–782, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kivaranovic:2021:LPM

[KL21b] Danijel Kivaranovic and Hannes Leeb. On the length of postmodel-selection confidence intervals conditional on polyhedral constraints. Journal of the American Statistical Association, 116(534):845–857, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kong:2023:DBG

[KLLL23] Xin-Bing Kong, Jin-Guan Lin, Cheng Liu, and Guang-Ying Liu. Discrepancy between global and local principal component analysis on large-panel high-frequency data. Journal of the American Statistical Association, 118(542):1333–1344, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kosorok:2021:ITM

[KLSZ21] Michael R. Kosorok, Eric B. Laber, Dylan S. Small, and Donglin Zeng. Introduction to the theory and methods special issue on precision medicine and individualized policy discovery. Journal of the American Statistical Association, 116 (533):159–161, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ke:2023:ENS

[KML23] Zheng Tracy Ke, Yucong Ma, and Xihong Lin. Estimation of the number of spiked eigenvalues in a covariance matrix by bulk eigenvalue matching analysis. Journal of the American Statistical Association, 118(541):374–392, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kellogg:2021:CMS

[KMPT21] Maxwell Kellogg, Magne Mogstad, Guillaume A. Pouliot, and Alexander Torgovitsky. Combining matching and synthetic control to tradeoff biases from extrapolation and interpolation. Journal of the American Statistical Association, 116 (536):1804–1816, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kowal:2022:FOT

[Kow22] Daniel R. Kowal. Fast, optimal, and targeted predictions using parameterized decision analysis. Journal of the American Statistical Association, 117(540):1875–1886, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kuchibhotla:2023:SEC

[KPS23] Arun K. Kuchibhotla, Rohit K. Patra, and Bodhisattva Sen. Semiparametric efficiency in convexity constrained singleindex model. Journal of the American Statistical Association, 118(541):272–286, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kock:2022:FST

[KPV22] Anders Bredahl Kock, David Preinerstorfer, and Bezirgen Veliyev. Functional sequential treatment allocation. Journal of the American Statistical Association, 117(539):1311–1323, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kaji:2023:MHC

[KR23] Tetsuya Kaji and Veronika Rocková. Metropolis–Hastings via classification. Journal of the American Statistical Association, 118(544):2533–2547, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kim:2024:HTC

[KRW24] Jae Kwang Kim, J. N. K. Rao, and Zhonglei Wang. Hypotheses testing from complex survey data using bootstrap weights: a unified approach. Journal of the American Statistical Association, 119(546):1229–1239, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Katzfuss:2024:SBT

[KS24] Matthias Katzfuss and Florian Schäfer. Scalable Bayesian transport maps for high-dimensional non-Gaussian spatial fields. Journal of the American Statistical Association, 119 (546):1409–1423, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Katsevich:2023:FRS

[KSB23] Eugene Katsevich, Chiara Sabatti, and Marina Bogomolov. Filtering the rejection set while preserving false discovery rate control. Journal of the American Statistical Association, 118(541):165–176, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Karmakar:2021:RDM

[KSR21] Bikram Karmakar, Dylan S. Small, and Paul R. Rosenbaum. Reinforced designs: Multiple instruments plus control groups as evidence factors in an observational study of the effectiveness of Catholic schools. Journal of the American Statistical Association, 116(533):82–92, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Katzfuss:2020:EKM

[KSW20] Matthias Katzfuss, Jonathan R. Stroud, and Christopher K. Wikle. Ensemble Kalman methods for high-dimensional hierarchical dynamic space-time models. Journal of the American Statistical Association, 115(530):866–885, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Klusowski:2024:LSP

[KT24] Jason M. Klusowski and Peter M. Tian. Large scale prediction with decision trees. Journal of the American Statistical Association, 119(545):525–537, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kamm:2020:EID

[KTDS20] Jack Kamm, Jonathan Terhorst, Richard Durbin, and Yun S. Song. Efficiently inferring the demographic history of many populations with allele count data. Journal of the American Statistical Association, 115(531):1472–1487, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ke:2024:UST

[KW24] Zheng Tracy Ke and Minzhe Wang. Using SVD for topic modeling. Journal of the American Statistical Association, 119(545):434–449, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ke:2020:ECC

[KY20] Chenlu Ke and Xiangrong Yin. Expected conditional characteristic function-based measures for testing independence. Journal of the American Statistical Association, 115(530): 985–996, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lila:2020:SAF

[LA20] Eardi Lila and John A. D. Aston. Statistical analysis of functions on surfaces, with an application to medical imaging. Journal of the American Statistical Association, 115(531): 1420–1434, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Law:2023:DSF

[LB23] Michael Law and Peter Bühlmann. Discussion of “A Scale-Free Approach for False Discovery Rate Control in Generalized Linear Models”. Journal of the American Statistical Association, 118(543):1578–1583, 2023. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [DLXL23c].

LeBlanc:2024:RSR

[LBF^{+}24] Patrick M. LeBlanc, David Banks, Linhui Fu, Mingyan Li, Zhengyu Tang, and Qiuyi Wu. Recommender systems: a review. Journal of the American Statistical Association, 119 (545):773–785, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Laga:2021:TYN

[LBN21] Ian Laga, Le Bao, and Xiaoyue Niu. Thirty years of the network scale-up method. Journal of the American Statistical Association, 116(535):1548–1559, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Laga:2023:CNS

[LBN23] Ian Laga, Le Bao, and Xiaoyue Niu. A correlated network scale-up model: Finding the connection between subpopulations. Journal of the American Statistical Association, 118 (543):1515–1524, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Luo:2020:FSC

[LC20] Shan Luo and Zehua Chen. Feature selection by canonical correlation search in high-dimensional multiresponse models with complex group structures. Journal of the American Statistical Association, 115(531):1227–1235, 2020. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lemyre:2022:SED

[LCD22] Félix Camirand Lemyre, Raymond J. Carroll, and Aurore Delaigle. Semiparametric estimation of the distribution of episodically consumed foods measured with error. Journal of the American Statistical Association, 117(537):469–481, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2022:IHD

[LCL22] Sai Li, T. Tony Cai, and Hongzhe Li. Inference for highdimensional linear mixed-effects models: a quasi-likelihood approach. Journal of the American Statistical Association, 117(540):1835–1846, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2023:TLL

[LCL23] Sai Li, T. Tony Cai, and Hongzhe Li. Transfer learning in large-scale Gaussian graphical models with false discovery rate control. Journal of the American Statistical Association, 118(543):2171–2183, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2020:CWI

[LD20] Meng Li and David B. Dunson. Comparing and weighting imperfect models using D-probabilities. Journal of the American Statistical Association, 115(531):1349–1360, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2024:ODE

[LD24] Hang Li and Enrique Del Castillo. Optimal design of experiments on Riemannian manifolds. Journal of the American Statistical Association, 119(546):875–886, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2022:CVE

[LDL^{+}22] Jeremiah Zhe Liu, Wenying Deng, Jane Lee, Pi i Debby Lin, Linda Valeri, David C. Christiani, David C. Bellinger, Robert O. Wright, Maitreyi M. Mazumdar, and Brent A. Coull. A cross-validated ensemble approach to robust hypothesis testing of continuous nonlinear interactions: Application to nutrition-environment studies. Journal of the American Statistical Association, 117(538):561–573, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lei:2020:CVC

[Lei20] Jing Lei. Cross-validation with confidence. Journal of the American Statistical Association, 115(532):1978–1997, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Leng:2020:BRS

[Len20] Ling Leng. Book review: Statistical Computing With R, (2nd ed.), by Maria L. Rizzo. Boca Raton, FL: Chapman & Hall/ CRC Press, 2019, xiv + 474 pp., $99.95(H), ISBN: 978-1-46-655332-3. Journal of the American Statistical Association, 115(531):1557–1558, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2022:BUC

[LH22] Yang Liu and Feifang Hu. Balancing unobserved covariates with covariate-adaptive randomized experiments. Journal of the American Statistical Association, 117(538):875–886, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liang:2023:TWS

[LHGY23] Decai Liang, Hui Huang, Yongtao Guan, and Fang Yao. Test of weak separability for spatially stationary functional field. Journal of the American Statistical Association, 118(543): 1606–1619, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2022:BAT

[LHS22] Yunxiao Li, Yi-Juan Hu, and Glen A. Satten. A bottom-up approach to testing hypotheses that have a branching tree dependence structure, with error rate control. Journal of the American Statistical Association, 117(538):664–677, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2023:EAB

[LHW23] Dongdong Li, X. Joan Hu, and Rui Wang. Evaluating association between two event times with observations subject to informative censoring. Journal of the American Statistical Association, 118(542):1282–1294, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2020:SIA

[Li20] Kathleen T. Li. Statistical inference for average treatment effects estimated by synthetic control methods. Journal of the American Statistical Association, 115(532):2068–2083, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2022:LHT

[Li22] Changcheng Li Runze Li. Linear hypothesis testing in linear models with high-dimensional responses. Journal of the American Statistical Association, 117(540):1738–1750, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:2023:CFG

[LJL^{+}23a] Kuang-Yao Lee, Dingjue Ji, Lexin Li, Todd Constable, and Hongyu Zhao. Conditional functional graphical models. Journal of the American Statistical Association, 118(541):257– 271, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2023:KBP

[LJL23b] Xinran Li, Bo Jiang, and Jun S. Liu. Kernel-based partial permutation test for detecting heterogeneous functional relationship. Journal of the American Statistical Association, 118 (542):1429–1447, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lu:2020:KMS

[LKL20] Junwei Lu, Mladen Kolar, and Han Liu. Kernel meets sieve: Post-regularization confidence bands for sparse additive model. Journal of the American Statistical Association, 115(532):2084–2099, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2022:MFF

[LKLL22] Wanjun Liu, Yuan Ke, Jingyuan Liu, and Runze Li. Modelfree feature screening and FDR control with knockoff features. Journal of the American Statistical Association, 117 (537):428–443, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liao:2021:PEL

[LKM21] Peng Liao, Predrag Klasnja, and Susan Murphy. Off-policy estimation of long-term average outcomes with applications to mobile health. Journal of the American Statistical Association, 116(533):382–391, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lukemire:2021:BJM

[LKPG21] Joshua Lukemire, Suprateek Kundu, Giuseppe Pagnoni, and Ying Guo. Bayesian joint modeling of multiple brain functional networks. Journal of the American Statistical Association, 116(534):518–530, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2022:IFR

[LL22] Quefeng Li and Lexin Li. Integrative factor regression and its inference for multimodal data analysis. Journal of the American Statistical Association, 117(540):2207–2221, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lei:2023:BAS

[LL23] Jing Lei and Kevin Z. Lin. Bias-adjusted spectral clustering in multi-layer stochastic block models. Journal of the American Statistical Association, 118(544):2433–2445, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2022:HCD

[LLB^{+}22] Tianxi Li, Lihua Lei, Sharmodeep Bhattacharyya, Koen Van den Berge, Purnamrita Sarkar, Peter J. Bickel, and Elizaveta Levina. Hierarchical community detection by recursive partitioning. Journal of the American Statistical Association, 117(538):951–968, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2022:NFL

[LLgX22] Dungang Liu, Regina Y. Liu, and Min ge Xie. Nonparametric fusion learning for multiparameters: Synthesize inferences from diverse sources using data depth and confidence distribution. Journal of the American Statistical Association, 117 (540):2086–2104, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Luckett:2020:EDT

[LLK^{+}20] Daniel J. Luckett, Eric B. Laber, Anna R. Kahkoska, David M. Maahs, Elizabeth Mayer-Davis, and Michael R. Kosorok. Estimating dynamic treatment regimes in mobile health using V-learning. Journal of the American Statistical Association, 115(530):692–706, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2021:DKM

[LLL21] Sijia Li, Xiudi Li, and Alex Luedtke. Discussion of Kallus (2020) and Mo, Qi, and Liu (2020): New objectives for policy learning. Journal of the American Statistical Association, 116(534):680–689, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Kal21a, MQL21a].

Li:2022:SDS

[LLL22a] Zilin Li, Yaowu Liu, and Xihong Lin. Simultaneous detection of signal regions using quadratic scan statistics with applications to whole genome association studies. Journal of the American Statistical Association, 117(538):823–834, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2022:MOR

[LLL22b] Yaowu Liu, Zilin Li, and Xihong Lin. A minimax optimal ridge-type set test for global hypothesis with applications in whole genome sequencing association studies. Journal of the American Statistical Association, 117(538):897–908, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:2023:NFG

[LLLZ23] Kuang-Yao Lee, Lexin Li, Bing Li, and Hongyu Zhao. Nonparametric functional graphical modeling through functional additive regression operator. Journal of the American Statistical Association, 118(543):1718–1732, 2023. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2023:ACD

[LLM23a] Yinpu Li, Antonio R. Linero, and Jared Murray. Adaptive conditional distribution estimation with Bayesian decision tree ensembles. Journal of the American Statistical Association, 118(543):2129–2142, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lin:2023:HDM

[LLM23b] Zhenhua Lin, Miles E. Lopes, and Hans-Georg Müller. Highdimensional MANOVA via bootstrapping and its application to functional and sparse count data. Journal of the American Statistical Association, 118(541):177–191, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2024:UIP

[LLPY24] Xiaohui Liu, Wei Long, Liang Peng, and Bingduo Yang. A unified inference for predictive quantile regression. Journal of the American Statistical Association, 119(546):1526–1540, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lin:2021:EFE

[LLR21a] Kevin Z. Lin, Jing Lei, and Kathryn Roeder. Exponentialfamily embedding with application to cell developmental trajectories for single-cell RNA-seq data. Journal of the American Statistical Association, 116(534):457–470, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See discussion [JJ21, HL21b, LLR21b].

Lin:2021:REF

[LLR21b] Kevin Z. Lin, Jing Lei, and Kathryn Roeder. Rejoinder for “Exponential-Family Embedding With Application to Cell Developmental Trajectories for Single-Cell RNA-Seq Data”. Journal of the American Statistical Association, 116(534): 478–480, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [LLR21a].

Lin:2021:CBS

[LLR21c] Kevin Z. Lin, Han Liu, and Kathryn Roeder. Covariancebased sample selection for heterogeneous data: Applications to gene expression and autism risk gene detection. Journal of the American Statistical Association, 116(533):54–67, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2022:ASM

[LLWL22] Jialiang Li, Jing Lv, Alan T. K. Wan, and Jun Liao. AdaBoost semiparametric model averaging prediction for multiple categories. Journal of the American Statistical Association, 117(537):495–509, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2021:APA

[LLYM21] Dungang Liu, Shaobo Li, Yan Yu, and Irini Moustaki. Assessing partial association between ordinal variables: Quantification, visualization, and hypothesis testing. Journal of the American Statistical Association, 116(534):955–968, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2023:GFM

[LLZL23] Wei Liu, Huazhen Lin, Shurong Zheng, and Jin Liu. Generalized factor model for ultra-high dimensional correlated variables with mixed types. Journal of the American Statistical Association, 118(542):1385–1401, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2022:RAA

[LLZZ22] Ting Li, Tengfei Li, Zhongyi Zhu, and Hongtu Zhu. Regression analysis of asynchronous longitudinal functional and scalar data. Journal of the American Statistical Association, 117(539):1228–1242, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2021:MRT

[LM21] Xinran Li and Xiao-Li Meng. A multi-resolution theory for approximating infinite- p -zero- n: Transitional inference, individualized predictions, and a world without bias-variance tradeoff. Journal of the American Statistical Association, 116 (533):353–367, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Laga:2020:BRM

[LN20] Ian Laga and Xiaoyue Niu. Book review: Model-Based Geostatistics for Global Public Health: Methods and Applications, by Peter J. Diggle and Emanuele Giorgi. Boca Raton, FL: Chapman & Hall/CRC Press, 2019, xxvi + 247 pp., $99.95(H), ISBN: 978-1-13-873235-3. Journal of the American Statistical Association, 115(530):1030–1032, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Laga:2022:MMP

[LNB22] Ian Laga, Xiaoyue Niu, and Le Bao. Modeling the marked presence-only data: a case study of estimating the female sex worker size in Malawi. Journal of the American Statistical Association, 117(537):27–37, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:2021:NDC

[LO21] Youjin Lee and Elizabeth L. Ogburn. Network dependence can lead to spurious associations and invalid inference. Journal of the American Statistical Association, 116(535):1060– 1074, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Loh:2020:CTF

[Loh20] Po-Ling Loh. Comment on “A Tuning-Free Robust and Efficient Approach to High-Dimensional Regression”. Journal of the American Statistical Association, 115(532):1715–1716, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [WPB^{+}20b].

Lunagomez:2021:MNP

[LOW21] Simón Lunagómez, Sofia C. Olhede, and Patrick J. Wolfe. Modeling network populations via graph distances. Journal of the American Statistical Association, 116(536):2023–2040, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lijoi:2024:FDD

[LPR24] Antonio Lijoi, Igor Prünster, and Tommaso Rigon. Finitedimensional discrete random structures and Bayesian clustering. Journal of the American Statistical Association, 119 (546):929–941, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Law:2023:IER

[LR23a] Michael Law and Ya’acov Ritov. Inference and estimation for random effects in high-dimensional linear mixed models. Journal of the American Statistical Association, 118(543): 1682–1691, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2023:VST

[LR23b] Yi Liu and Veronika Rocková. Variable selection via Thompson sampling. Journal of the American Statistical Association, 118(541):287–304, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2020:LRD

[LRS20] Degui Li, Peter M. Robinson, and Han Lin Shang. Long-range dependent curve time series. Journal of the American Statistical Association, 115(530):957–971, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2024:RBJ

[LRY24] Hanzhong Liu, Jiyang Ren, and Yuehan Yang. Randomizationbased joint central limit theorem and efficient covariate adjustment in randomized block 2^{K} factorial experiments. Journal of the American Statistical Association, 119(545):136– 150, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lu:2024:SDF

[LRZ24] Zudi Lu, Xiaohang Ren, and Rongmao Zhang. On semiparametrically dynamic functional-coefficient autoregressive spatio-temporal models with irregular location wide nonstationarity. Journal of the American Statistical Association, 119 (546):1032–1043, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2020:DTF

[LS20] Xiudi Li and Ali Shojaie. Discussion of “A tuning-free robust and efficient approach to high-dimensional regression”. Journal of the American Statistical Association, 115(532):1717– 1719, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [WPB^{+}20b].

Liu:2022:LSH

[LSB^{+}22] Zhonghua Liu, Jincheng Shen, Richard Barfield, Joel Schwartz, Andrea A. Baccarelli, and Xihong Lin. Large-scale hypothesis testing for causal mediation effects with applications in genome-wide epigenetic studies. Journal of the American Statistical Association, 117(537):67–81, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:2021:DHE

[LSD21] Kwonsang Lee, Dylan S. Small, and Francesca Dominici. Discovering heterogeneous exposure effects using randomization inference in air pollution studies. Journal of the American Statistical Association, 116(534):569–580, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2023:ATV

[LSN23] Ziyi Li, Yu Shen, and Jing Ning. Accommodating timevarying heterogeneity in risk estimation under the Cox model: a transfer learning approach. Journal of the American Statistical Association, 118(544):2276–2287, 2023. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2020:LRT

[LSP20] Chunlin Li, Xiaotong Shen, and Wei Pan. Likelihood ratio tests for a large directed acyclic graph. Journal of the American Statistical Association, 115(531):1304–1319, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2024:NCD

[LSP24] Chunlin Li, Xiaotong Shen, and Wei Pan. Nonlinear causal discovery with confounders. Journal of the American Statistical Association, 119(546):1205–1214, 2024. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2021:EEO

[LSRR21] Lin Liu, Zach Shahn, James M. Robins, and Andrea Rotnitzky. Efficient estimation of optimal regimes under a no direct effect assumption. Journal of the American Statistical Association, 116(533):224–239, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2023:BON

[LSS23] Zhu Li, Weijie J. Su, and Dino Sejdinovic. Benign overfitting and noisy features. Journal of the American Statistical Association, 118(544):2876–2888, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:2023:EEC

[LSX23] Sze Ming Lee, Tony Sit, and Gongjun Xu. Efficient estimation for censored quantile regression. Journal of the American Statistical Association, 118(544):2762–2775, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2024:FDP

[LSZ24] Cheng Li, Saifei Sun, and Yichen Zhu. Fixed-domain posterior contraction rates for spatial Gaussian process model with nugget. Journal of the American Statistical Association, 119 (546):1336–1347, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Linton:2022:CFM

[LT22] Oliver B. Linton and Haihan Tang. Comment on “Factor Models for High-Dimensional Tensor Time Series” by Rong Chen, Dan Yang, and Cun-Hui Zhang. Journal of the American Statistical Association, 117(537):117, 2022. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liang:2022:MFB

[LTB22] Tengyuan Liang and Hai Tran-Bach. Mehler’s formula, branching process, and compositional kernels of deep neural networks. Journal of the American Statistical Association, 117(539):1324–1337, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2021:IRA

[LTO^{+}21] Zhigang Li, Lu Tian, A. James O’Malley, Margaret R. Karagas, Anne G. Hoen, Brock C. Christensen, Juliette C. Madan, Quran Wu, Raad Z. Gharaibeh, Christian Jobin, and Hongzhe Li. IFAA: Robust association identification and inference for absolute abundance in microbiome analyses. Journal of the American Statistical Association, 116(536):1595– 1608, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lin:2021:BBA

[LTY21] Ruitao Lin, Peter F. Thall, and Ying Yuan. BAGS: a Bayesian adaptive group sequential trial design with subgroup-specific survival comparisons. Journal of the American Statistical Association, 116(533):322–334, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lawrence:2021:CGS

[LV21] Earl Lawrence and Scott Vander Wiel. Comment on “A Gibbs Sampler for a Class of Random Convex Polytopes”. Journal of the American Statistical Association, 116(535):1201–1203, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [JGED21a, JGED21b].

Lee:2020:HMN

[LW20] Clement Lee and Darren J. Wilkinson. A hierarchical model of nonhomogeneous Poisson processes for Twitter retweets. Journal of the American Statistical Association, 115(529):1– 15, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lin:2022:MCE

[LW22a] Zhenhua Lin and Jane-Ling Wang. Mean and covariance estimation for functional snippets. Journal of the American Statistical Association, 117(537):348–360, 2022. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lyu:2022:AEN

[LW22b] Ziyang Lyu and A. H. Welsh. Asymptotics for EBLUPs: Nested error regression models. Journal of the American Statistical Association, 117(540):2028–2042, 2022. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2023:ADA

[LWLL23] Yan Liu, Dewei Wang, Li Li, and Dingsheng Li. Assessing disparities in Americans’ exposure to PCBs and PBDEs based on NHANES pooled biomonitoring data. Journal of the American Statistical Association, 118(543):1538–1550, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2022:PAR

[LWSL22] Jianyu Liu, Haodong Wang, Wei Sun, and Yufeng Liu. Prioritizing autism risk genes using personalized graphical models estimated from single-cell RNA-seq data. Journal of the American Statistical Association, 117(537):38–51, 2022. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2021:SLS

[LWWftADNI21] Xinyi Li, Li Wang, Huixia Judy Wang, and for the Alzheimer’s Disease Neuroimaging Initiative. Sparse learning and structure identification for ultrahigh-dimensional imageon-scalar regression. Journal of the American Statistical Association, 116(536):1994–2008, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2021:BSS

[LWY^{+}21] Huazhang Li, Yaotian Wang, Guofen Yan, Yinge Sun, Seiji Tanabe, Chang-Chia Liu, Mark S. Quigg, and Tingting Zhang. A Bayesian state-space approach to mapping directional brain networks. Journal of the American Statistical Association, 116(536):1637–1647, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2020:CCT

[LX20] Yaowu Liu and Jun Xie. Cauchy combination test: a powerful test with analytic p-value calculation under arbitrary dependency structures. Journal of the American Statistical Association, 115(529):393–402, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2024:FSC

[LX24] Xingxiang Li and Chen Xu. Feature screening with conditional rank utility for big-data classification. Journal of the American Statistical Association, 119(546):1385–1395, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liang:2022:MNR

[LXJ22] Faming Liang, Jingnan Xue, and Bochao Jia. Markov neighborhood regression for high-dimensional inference. Journal of the American Statistical Association, 117(539):1200–1214, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2020:EMM

[LXLF20] Han Li, Minxuan Xu, Jun S. Liu, and Xiaodan Fan. An extended Mallows model for ranked data aggregation. Journal of the American Statistical Association, 115(530):730–746, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2023:TEH

[LXZZ23] Runze Li, Kai Xu, Yeqing Zhou, and Liping Zhu. Testing the effects of high-dimensional covariates via aggregating cumulative covariances. Journal of the American Statistical Association, 118(543):2184–2194, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liang:2022:SAM

[LY22] Muxuan Liang and Menggang Yu. A semiparametric approach to model effect modification. Journal of the American Statistical Association, 117(538):752–764, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2023:DIS

[LYC23] Hua Liu, Jinhong You, and Jiguo Cao. A dynamic interaction semiparametric function-on-scalar model. Journal of the American Statistical Association, 118(541):360–373, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2022:SMS

[LYL22] Xiao Liu, Kyongmin Yeo, and Siyuan Lu. Statistical modeling for spatio-temporal data from stochastic convection-diffusion processes. Journal of the American Statistical Association, 117(539):1482–1499, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2023:CSF

[LYZ23] Sai Li, Yisha Yao, and Cun-Hui Zhang. Comments on “A Scale-Free Approach for False Discovery Rate Control in Generalized Linear Models”. Journal of the American Statistical Association, 118(543):1586–1589, 2023. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [DLXL23c].

Liu:2024:PTM

[LYZL24] Wanjun Liu, Xiufan Yu, Wei Zhong, and Runze Li. Projection test for mean vector in high dimensions. Journal of the American Statistical Association, 119(545):744–756, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:2021:SKI

[LZ21a] DongHyuk Lee and Bin Zhu. A semiparametric kernel independence test with application to mutational signatures. Journal of the American Statistical Association, 116(536):

1648–1661, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liang:2021:DKM

[LZ21b] Muxuan Liang and Ying-Qi Zhao. Discussion of Kallus (2020) and Mo et al. (2020). Journal of the American Statistical Association, 116(534):690–693, 2021. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Kal21a, MQL21a].

Liang:2021:MRC

[LZCH21] Decai Liang, Haozhe Zhang, Xiaohui Chang, and Hui Huang. Modeling and regionalization of China’s PM_{2.5} using spatialfunctional mixture models. Journal of the American Statistical Association, 116(533):116–132, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2024:EIH

[LZCL24] Sai Li, Linjun Zhang, T. Tony Cai, and Hongzhe Li. Estimation and inference for high-dimensional generalized linear models with knowledge transfer. Journal of the American Statistical Association, 119(546):1274–1285, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2024:RPA

[LZH24] Changyu Liu, Xingqiu Zhao, and Jian Huang. A random projection approach to hypothesis tests in high-dimensional single-index models. Journal of the American Statistical Association, 119(546):1008–1018, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2021:TMT

[LZQ^{+}21] Yutong Li, Ruoqing Zhu, Annie Qu, Han Ye, and Zhankun Sun. Topic modeling on triage notes with semiorthogonal nonnegative matrix factorization. Journal of the American Statistical Association, 116(536):1609–1624, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Luo:2023:RTR

[LZS23] Lan Luo, Ling Zhou, and Peter X.-K. Song. Real-time regression analysis of streaming clustered data with possible abnormal data batches. Journal of the American Statistical Association, 118(543):2029–2044, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2024:RHD

[LZX^{+}24] Bingyuan Liu, Qi Zhang, Lingzhou Xue, Peter X.-K. Song, and Jian Kang. Robust high-dimensional regression with coefficient thresholding and its application to imaging data analysis. Journal of the American Statistical Association, 119(545): 715–729, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2023:GLA

[LZZ23a] Lexin Li, Jing Zeng, and Xin Zhang. Generalized liquid association analysis for multimodal data integration. Journal of the American Statistical Association, 118(543):1984–1996, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2023:AMH

[LZZ23b] Lingzhu Li, Xuehu Zhu, and Lixing Zhu. Adaptive-to-model hybrid of tests for regressions. Journal of the American Statistical Association, 118(541):514–523, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Matechou:2023:CRM

[MA23] Eleni Matechou and Raffaele Argiento. Capture-recapture models with heterogeneous temporary emigration. Journal of the American Statistical Association, 118(541):56–69, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

McAlinn:2020:MBP

[MANW20] Kenichiro McAlinn, Knut Are Aastveit, Jouchi Nakajima, and Mike West. Multivariate Bayesian predictive synthesis in macroeconomic forecasting. Journal of the American Statistical Association, 115(531):1092–1110, 2020. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Martinez:2021:BOF

[Mar21] Wendy L. Martinez. Back to our future: Text analytics insights. Journal of the American Statistical Association, 116 (535):1039–1047, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mukherjee:2022:DFM

[MAZB22] Somabha Mukherjee, Divyansh Agarwal, Nancy R. Zhang, and Bhaswar B. Bhattacharya. Distribution-free multisample tests based on optimal matchings with applications to single cell genomics. Journal of the American Statistical Association, 117(538):627–638, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ma:2023:BMR

[MB23] Pulong Ma and Anindya Bhadra. Beyond Matérn: On a class of interpretable confluent hypergeometric covariance functions. Journal of the American Statistical Association, 118 (543):2045–2058, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Matuk:2022:BFS

[MBCK22] James Matuk, Karthik Bharath, Oksana Chkrebtii, and Sebastian Kurtek. Bayesian framework for simultaneous registration and estimation of noisy, sparse, and fragmented functional data. Journal of the American Statistical Association, 117(540):1964–1980, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

McShane:2022:MLV

[MBH22a] Blakeley B. McShane, Ulf Böckenholt, and Karsten T. Hansen. Modeling and learning from variation and covariation. Journal of the American Statistical Association, 117 (540):1627–1630, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

McShane:2022:VCL

[MBH22b] Blakeley B. McShane, Ulf Böckenholt, and Karsten T. Hansen. Variation and covariation in large-scale replication projects: an evaluation of replicability. Journal of the American Statistical Association, 117(540):1605–1621, 2022. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Man:2022:MJA

[MC22] Albert Xingyi Man and Steven Andrew Culpepper. A modejumping algorithm for Bayesian factor analysis. Journal of the American Statistical Association, 117(537):277–290, 2022.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

McDonald:2020:BRS

[McD20] Daniel J. McDonald. Book review: Sufficient Dimension Reduction: Methods and Applications With R, by Bing Li. Boca Raton, FL: Chapman & Hall/CRC Press, 2018, xxi pp., $105.00(H), ISBN: 978-1-49-870447-2. Journal of the American Statistical Association, 115(530):1032–1033, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ma:2021:GSH

[MCL21a] Rong Ma, T. Tony Cai, and Hongzhe Li. Global and simultaneous hypothesis testing for high-dimensional logistic regression models. Journal of the American Statistical Association, 116(534):984–998, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ma:2021:OPR

[MCL21b] Rong Ma, T. Tony Cai, and Hongzhe Li. Optimal permutation recovery in permuted monotone matrix model. Journal of the American Statistical Association, 116(535):1358–1372, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Michael:2024:IVE

[MCLT24] Haben Michael, Yifan Cui, Scott A. Lorch, and Eric J. Tchetgen Tchetgen. Instrumental variable estimation of marginal structural mean models for time-varying treatment. Journal of the American Statistical Association, 119(546):1240–1251, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mao:2020:BGC

[MCM20] Jialiang Mao, Yuhan Chen, and Li Ma. Bayesian graphical compositional regression for microbiome data. Journal of the American Statistical Association, 115(530):610–624, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mukhopadhyay:2020:TRP

[MD20] Minerva Mukhopadhyay and David B. Dunson. Targeted random projection for prediction from high-dimensional features. Journal of the American Statistical Association, 115 (532):1998–2010, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Miao:2023:IEM

[MHOZ23] Wang Miao, Wenjie Hu, Elizabeth L. Ogburn, and Xiao-Hua Zhou. Identifying effects of multiple treatments in the presence of unmeasured confounding. Journal of the American Statistical Association, 118(543):1953–1967, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mai:2023:CGT

[MHZ23] Qing Mai, Di He, and Hui Zou. Coordinatewise Gaussianization: Theories and applications. Journal of the American Statistical Association, 118(544):2329–2343, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mies:2023:FEC

[Mie23] Fabian Mies. Functional estimation and change detection for nonstationary time series. Journal of the American Statistical Association, 118(542):1011–1022, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ma:2023:PBC

[MJ23] Haiqiang Ma and Jiming Jiang. Pseudo-Bayesian classified mixed model prediction. Journal of the American Statistical Association, 118(543):1747–1759, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mozharovskyi:2020:NID

[MJH20] Pavlo Mozharovskyi, Julie Josse, and François Husson. Nonparametric imputation by data depth. Journal of the American Statistical Association, 115(529):241–253, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ma:2024:MTL

[MKftADNI24] Xin Ma, Suprateek Kundu, and for the Alzheimer’s Disease Neuroimaging Initiative. Multi-task learning with highdimensional noisy images. Journal of the American Statistical Association, 119(545):650–663, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Miao:2024:FPP

[MKV^{+}24] Zhen Miao, Weihao Kong, Ramya Korlakai Vinayak, Wei Sun, and Fang Han. Fisher–Pitman permutation tests based on nonparametric Poisson mixtures with application to single cell genomics. Journal of the American Statistical Association, 119(545):394–406, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mork:2024:HDL

[MKW^{+}24] Daniel Mork, Marianthi-Anna Kioumourtzoglou, Marc Weisskopf, Brent A. Coull, and Ander Wilson. Heterogeneous distributed lag models to estimate personalized effects of maternal exposures to air pollution. Journal of the American Statistical Association, 119(545):14–26, 2024. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Moura:2021:IMR

[MKZ^{+}21] Ricardo Moura, Martin Klein, John Zylstra, Carlos A. Coelho, and Bimal Sinha. Inference for multivariate regression model based on synthetic data generated using plug-in sampling. Journal of the American Statistical Association, 116(534):720–733, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ma:2022:BIN

[MLH^{+}22] Tianwen Ma, Yang Li, Jane E. Huggins, Ji Zhu, and Jian Kang. Bayesian inferences on neural activity in EEG-based brain-computer interface. Journal of the American Statistical Association, 117(539):1122–1133, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ma:2021:ISA

[MLL21] Cong Ma, Junwei Lu, and Han Liu. Inter-subject analysis: a partial Gaussian graphical model approach. Journal of the American Statistical Association, 116(534):746–755, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ma:2024:NUF

[MLZH24] Wei Ma, Ping Li, Li-Xin Zhang, and Feifang Hu. A new and unified family of covariate adaptive randomization procedures and their properties. Journal of the American Statistical Association, 119(545):151–162, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Masini:2021:CAA

[MM21] Ricardo Masini and Marcelo C. Medeiros. Counterfactual analysis with artificial controls: Inference, high dimensions, and nonstationarity. Journal of the American Statistical Association, 116(536):1773–1788, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mohammadi:2023:ABS

[MML23] Reza Mohammadi, Hélène Massam, and Gérard Letac. Accelerating Bayesian structure learning in sparse Gaussian graphical models. Journal of the American Statistical Association, 118(542):1345–1358, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mao:2024:VMF

[MMR24] Huiying Mao, Ryan Martin, and Brian J. Reich. Valid modelfree spatial prediction. Journal of the American Statistical Association, 119(546):904–914, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

McKennan:2022:EAU

[MN22] Chris McKennan and Dan Nicolae. Estimating and accounting for unobserved covariates in high-dimensional correlated data. Journal of the American Statistical Association, 117 (537):225–236, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

McCulloch:2023:IPW

[MN23] Charles E. McCulloch and John M. Neuhaus. Improving predictions when interest focuses on extreme random effects. Journal of the American Statistical Association, 118(541): 504–513, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Morales-Navarrete:2024:MPR

[MNBCCC24] Diego Morales-Navarrete, Moreno Bevilacqua, Christian Caamaño-Carrillo, and Luis M. Castro. Modeling point referenced spatial count data: a Poisson process approach. Journal of the American Statistical Association, 119(545):664–677, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Menacher:2024:BLE

[MNHG24] Anna Menacher, Thomas E. Nichols, Chris Holmes, and Habib Ganjgahi. Bayesian lesion estimation with a structured spike-and-slab prior. Journal of the American Statistical Association, 119(545):66–80, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mejia:2020:TIC

[MNW^{+}20] Amanda F. Mejia, Mary Beth Nebel, Yikai Wang, Brian S. Caffo, and Ying Guo. Template independent component analysis: Targeted and reliable estimation of subject-level brain networks using big data population priors. Journal of the American Statistical Association, 115(531):1151–1177, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mastrototaro:2024:FNS

[MOA24] Alessandro Mastrototaro, Jimmy Olsson, and Johan Alenlöv. Fast and numerically stable particle-based online additive smoothing: The AdaSmooth algorithm. Journal of the American Statistical Association, 119(545):356–367, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Moraga:2021:BRH

[Mor21] Paula Moraga. Book review: Handbook of Spatial Epidemiology. Journal of the American Statistical Association, 116 (533):451–453, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mohan:2021:GMP

[MP21] Karthika Mohan and Judea Pearl. Graphical models for processing missing data. Journal of the American Statistical Association, 116(534):1023–1037, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Masak:2023:RSC

[MP23] Tomas Masak and Victor M. Panaretos. Random surface covariance estimation by shifted partial tracing. Journal of the American Statistical Association, 118(544):2562–2574, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

McElroy:2024:ESD

[MP24] Tucker McElroy and Dimitris N. Politis. Estimating the spectral density at frequencies near zero. Journal of the American Statistical Association, 119(545):612–624, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mo:2021:LOD

[MQL21a] Weibin Mo, Zhengling Qi, and Yufeng Liu. Learning optimal distributionally robust individualized treatment rules. Journal of the American Statistical Association, 116(534):659– 674, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See discussion [VD21, LLL21, LZ21b] and rejoinder [MQL21b].

Mo:2021:RLO

[MQL21b] Weibin Mo, Zhengling Qi, and Yufeng Liu. Rejoinder: Learning optimal distributionally robust individualized treatment rules. Journal of the American Statistical Association, 116 (534):699–707, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [MQL21a, VD21, LLL21, LZ21b].

Ma:2020:SIC

[MQLH20] Wei Ma, Yichen Qin, Yang Li, and Feifang Hu. Statistical inference for covariate-adaptive randomization procedures. Journal of the American Statistical Association, 115 (531):1488–1497, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Marra:2020:CLB

[MR20] Giampiero Marra and Rosalba Radice. Copula link-based additive models for right-censored event time data. Journal of the American Statistical Association, 115(530):886–895, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Molstad:2023:LBA

[MR23] Aaron J. Molstad and Adam J. Rothman. A likelihood-based approach for multivariate categorical response regression in high dimensions. Journal of the American Statistical Association, 118(542):1402–1414, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mao:2021:EMM

[MSC21] Xueyu Mao, Purnamrita Sarkar, and Deepayan Chakrabarti. Estimating mixed memberships with sharp eigenvector deviations. Journal of the American Statistical Association, 116 (536):1928–1940, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Malinsky:2021:SIN

[MST21] Daniel Malinsky, Ilya Shpitser, and Eric J. Tchetgen Tchetgen. Semiparametric inference for nonmonotone missing-not-at-random data: The no self-censoring model. Journal of the American Statistical Association, 0(0):1–9, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See correction [Ano22b].

Malinsky:2022:SIN

[MST22] Daniel Malinsky, Ilya Shpitser, and Eric J. Tchetgen Tchetgen. Semiparametric inference for nonmonotone missing-not-at-random data: The no self-censoring model. Journal of the American Statistical Association, 117(539):1415–1423, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mukherjee:2024:BRT

[Muk24] Somabha Mukherjee. Book review: Theory of Statistical Inference, Anthony Almudevar, Boca Raton, FL: Chapman & Hall/CRC Press, 2022, xii + 457 pp., $130.00(H), ISBN 978-0-367-48875-8. Journal of the American Statistical Association, 119(546):1689–1690, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Murray:2021:LLB

[Mur21] Jared S. Murray. Log-linear Bayesian additive regression trees for multinomial logistic and count regression models. Journal of the American Statistical Association, 116(534):756–769, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mathur:2020:SAU

[MV20] Maya B. Mathur and Tyler J. VanderWeele. Sensitivity analysis for unmeasured confounding in meta-analyses. Journal of the American Statistical Association, 115(529):163–172, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ma:2020:RIU

[MW20] Xinwei Ma and Jingshen Wang. Robust inference using inverse probability weighting. Journal of the American Statistical Association, 115(532):1851–1860, 2020. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mejia:2020:BGL

[MYB^{+}20] Amanda F. Mejia, Yu (Ryan) Yue, David Bolin, Finn Lindgren, and Martin A. Lindquist. A Bayesian general linear modeling approach to cortical surface fMRI data analysis. Journal of the American Statistical Association, 115(530): 501–520, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mai:2022:DEE

[MZPD22] Qing Mai, Xin Zhang, Yuqing Pan, and Kai Deng. A doubly enhanced EM algorithm for model-based tensor clustering. Journal of the American Statistical Association, 117(540): 2120–2134, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Miao:2023:WBI

[MZW23] Rui Miao, Xiaoke Zhang, and Raymond K. W. Wong. A wavelet-based independence test for functional data with an application to MEG functional connectivity. Journal of the American Statistical Association, 118(543):1876–1889, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nie:2021:LWT

[NBW21] Xinkun Nie, Emma Brunskill, and Stefan Wager. Learning when-to-treat policies. Journal of the American Statistical Association, 116(533):392–409, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nandy:2022:CIN

[NCL22] Debmalya Nandy, Francesca Chiaromonte, and Runze Li. Covariate information number for feature screening in ultrahighdimensional supervised problems. Journal of the American Statistical Association, 117(539):1516–1529, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nie:2020:SME

[NCMK20] Xiao Nie, Peter Chien, Dane Morgan, and Amy Kaczmarowski. A statistical method for emulation of computer models with invariance-preserving properties, with application to structural energy prediction. Journal of the American Statistical Association, 115(532):1798–1811, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Natarajan:2024:CRB

[NDH^{+}24] Abhinav Natarajan, Maria De Iorio, Andreas Heinecke, Emanuel Mayer, and Simon Glenn. Cohesion and repulsion in Bayesian distance clustering. Journal of the American Statistical Association, 119(546):1374–1384, 2024. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nemeth:2021:SGM

[NF21] Christopher Nemeth and Paul Fearnhead. Stochastic gradient Markov Chain Monte Carlo. Journal of the American Statistical Association, 116(533):433–450, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ni:2022:BRB

[Ni22] Yang Ni. Book review: Bayesian Thinking in Biostatistics Gary L. Rosner, Purushottam W. Laud, and Wesley O. Johnson. Boca Raton, FL: Chapman & Hall/CRC Press, 2021, xix + 607 pp., $120.00(H), $96.00(e-book), ISBN: 978-1-4398-0008-9(H), 978-1-4398-0010-2(e-book). Journal of the American Statistical Association, 117(538):1041–1042, 2022. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ni:2023:BRH

[Ni23] Yang Ni. Book review: Handbook of Bayesian Variable Selection, Mahlet G. Tadesse and Marina Vannucci, eds., Boca Raton, FL: Chapman & Hall/CRC Press, 2022, xxiii + 466 pp., $200.00(H), ISBN: 978-0-367-54376-1. Journal of the American Statistical Association, 118(542):1449–1450, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Niezink:2023:DTT

[Nie23] Nynke M. D. Niezink. Discussion of “A tale of two datasets: Representativeness and generalisability of inference for samples of networks”. Journal of the American Statistical Association, 118(544):2232–2234, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [KCH23b, KCH23a].

Nattino:2021:TME

[NLS^{+}21] Giovanni Nattino, Bo Lu, Junxin Shi, Stanley Lemeshow, and Henry Xiang. Triplet matching for estimating causal effects with three treatment arms: a comparative study of mortality by trauma center level. Journal of the American Statistical Association, 116(533):44–53, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ni:2020:BDF

[NMJ20] Yang Ni, Peter Müller, and Yuan Ji. Bayesian double feature allocation for phenotyping with electronic health records. Journal of the American Statistical Association, 115(532): 1620–1634, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nethery:2021:EHI

[NMSD21] Rachel C. Nethery, Fabrizia Mealli, Jason D. Sacks, and Francesca Dominici. Evaluation of the health impacts of the 1990 Clean Air Act Amendments using causal inference and machine learning. Journal of the American Statistical Association, 116(535):1128–1139, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nie:2023:BBS

[NR23] Lizhen Nie and Veronika Rocková. Bayesian bootstrap spikeand-slab LASSO. Journal of the American Statistical Association, 118(543):2013–2028, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nishimura:2023:PPC

[NS23] Akihiko Nishimura and Marc A. Suchard. Prior-preconditioned conjugate gradient method for accelerated Gibbs sampling in “large n, large p” Bayesian sparse regression. Journal of the American Statistical Association, 118(544):2468–2481, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nagler:2024:SEE

[NV24] Thomas Nagler and Thibault Vatter. Solving estimating equations with copulas. Journal of the American Statistical Association, 119(546):1168–1180, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Olivella:2022:DSB

[OPI22] Santiago Olivella, Tyler Pratt, and Kosuke Imai. Dynamic stochastic blockmodel regression for network data: Application to international militarized conflicts. Journal of the American Statistical Association, 117(539):1068–1081, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ogburn:2024:CIS

[OSDvdL24] Elizabeth L. Ogburn, Oleg Sofrygin, Iván Díaz, and Mark J. van der Laan. Causal inference for social network data. Journal of the American Statistical Association, 119(545):597– 611, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Overstall:2020:BOD

[OWP20] Antony M. Overstall, David C. Woods, and Ben M. Parker. Bayesian optimal design for ordinary differential equation models with application in biological science. Journal of the American Statistical Association, 115(530):583–598, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ouyang:2022:CFM

[OY22] Jialin Ouyang and Ming Yuan. Comments on “Factor Models for High-Dimensional Tensor Time Series”. Journal of the American Statistical Association, 117(537):124–127, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Painsky:2023:GGT

[Pai23] Amichai Painsky. Generalized Good–Turing improves missing mass estimation. Journal of the American Statistical Association, 118(543):1890–1899, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Park:2022:BRH

[Par22] Hyunwoo Park. Book review: A History of Data Visualization and Graphic Communication, Michael Friendly, Howard Wainer, Cambridge, MA: Harvard University Press, 2021, 320 pages, $49.95, £39.95, EUR45.00, ISBN 978-0-674-97523-1. Journal of the American Statistical Association, 117(539): 1601–1603, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Park:2023:BRB

[Par23] Jaewoo Park. Book review: Bayesian Filtering and Smoothing, 2nd ed. Journal of the American Statistical Association, 118(544):2943–2945, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Peruzzi:2022:HSB

[PBF22] Michele Peruzzi, Sudipto Banerjee, and Andrew O. Finley. Highly scalable Bayesian geostatistical modeling via meshed Gaussian processes on partitioned domains. Journal of the American Statistical Association, 117(538):969–982, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Papadogeorgou:2023:CIL

[PBOD23] Georgia Papadogeorgou, Carolina Bello, Otso Ovaskainen, and David B. Dunson. Covariate-informed latent interaction models: Addressing geographic & taxonomic bias in predicting bird-plant interactions. Journal of the American Statistical Association, 118(544):2250–2261, 2023. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pexfa:2022:CFM

[Peñ22a] Daniel Peña. Comment on “Factor Models for High-Dimensional Tensor Time Series”. Journal of the American Statistical Association, 117(537):118–123, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pensky:2022:DCI

[Pen22b] Marianna Pensky. Discussion of “Confidence Intervals for Nonparametric Empirical Bayes Analysis”. Journal of the American Statistical Association, 117(539):1183–1185, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [IW22a] and rejoinder [IW22b].

Pimentel:2020:OTM

[PK20] Samuel D. Pimentel and Rachel R. Kelz. Optimal tradeoffs in matched designs comparing US-trained and internationally trained surgeons. Journal of the American Statistical Association, 115(532):1675–1688, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Park:2023:ALA

[PK23] Chan Park and Hyunseung Kang. Assumption-lean analysis of cluster randomized trials in infectious diseases for intentto-treat effects and network effects. Journal of the American Statistical Association, 118(542):1195–1206, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Paulon:2021:BSL

[PLCS21] Giorgio Paulon, Fernando Llanos, Bharath Chandrasekaran, and Abhra Sarkar. Bayesian semiparametric longitudinal drift-diffusion mixed models for tone learning in adults. Journal of the American Statistical Association, 116(535):1114– 1127, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Park:2023:CYP

[PLL23] Yeonjoo Park, Bo Li, and Yehua Li. Crop yield prediction using Bayesian spatially varying coefficient models with functional predictors. Journal of the American Statistical Association, 118(541):70–83, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pan:2023:RRP

[PLSZ23] Yinghao Pan, Eric B. Laber, Maureen A. Smith, and Ying-Qi Zhao. Reinforced risk prediction with budget constraint using irregularly measured data from electronic health records. Journal of the American Statistical Association, 118(542): 1090–1101, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Park:2024:LRR

[PLZ24] Seyoung Park, Eun Ryung Lee, and Hongyu Zhao. Low-rank regression models for multiple binary responses and their applications to cancer cell-line encyclopedia data. Journal of the American Statistical Association, 119(545):202–216, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Perreault:2023:HTS

[PND23] Samuel Perreault, Johanna G. Neslehová, and Thierry Duchesne. Hypothesis tests for structured rank correlation matrices. Journal of the American Statistical Association, 118 (544):2889–2900, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Paindaveine:2022:PMT

[PRV22] Davy Paindaveine, Joséa Rasoafaraniaina, and Thomas Verdebout. Preliminary multiple-test estimation, with applications to k-sample covariance estimation. Journal of the American Statistical Association, 117(540):1904–1915, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Paparoditis:2023:BPB

[PS23] Efstathios Paparoditis and Han Lin Shang. Bootstrap prediction bands for functional time series. Journal of the American Statistical Association, 118(542):972–986, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See correction [Ano23a].

Park:2021:ICV

[PSBB21] Junhyung Park, Frederic Paik Schoenberg, Andrea L. Bertozzi, and P. Jeffrey Brantingham. Investigating clustering and violence interruption in gang-related violent crime data using spatial-temporal point processes with covariates. Journal of the American Statistical Association, 116(536): 1674–1687, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ponnet:2024:RIM

[PSVV24] Jolien Ponnet, Pieter Segaert, Stefan Van Aelst, and Tim Verdonck. Robust inference and modeling of mean and dispersion for generalized linear models. Journal of the American Statistical Association, 119(545):678–689, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Panigrahi:2023:ASI

[PT23] Snigdha Panigrahi and Jonathan Taylor. Approximate selective inference via maximum likelihood. Journal of the American Statistical Association, 118(544):2810–2820, 2023. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pan:2020:BCG

[PWZ^{+}20] Wenliang Pan, Xueqin Wang, Heping Zhang, Hongtu Zhu, and Jin Zhu. Ball covariance: a generic measure of dependence in Banach space. Journal of the American Statistical Association, 115(529):307–317, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Park:2021:IMD

[PXZ21] Seyoung Park, Hao Xu, and Hongyu Zhao. Integrating multidimensional data for clustering analysis with applications to cancer patient data. Journal of the American Statistical Association, 116(533):14–26, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pan:2021:IDR

[PZ21a] Yinghao Pan and Ying-Qi Zhao. Improved doubly robust estimation in learning optimal individualized treatment rules. Journal of the American Statistical Association, 116(533):

283–294, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Prentice:2021:RMM

[PZ21b] Ross L. Prentice and Shanshan Zhao. Regression models and multivariate life tables. Journal of the American Statistical Association, 116(535):1330–1345, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Qiu:2021:OID

[QCS^{+}21a] Hongxiang Qiu, Marco Carone, Ekaterina Sadikova, Maria Petukhova, Ronald C. Kessler, and Alex Luedtke. Optimal individualized decision rules using instrumental variable methods. Journal of the American Statistical Association, 116(533):174–191, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Qiu:2021:ROI

[QCS^{+}21b] Hongxiang Qiu, Marco Carone, Ekaterina Sadikova, Maria Petukhova, Ronald C. Kessler, and Alex Luedtke. Rejoinder: Optimal individualized decision rules using instrumental variable methods. Journal of the American Statistical Association, 116(533):207–209, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [CT21a].

Quick:2021:RMU

[QDL21a] Corbin Quick, Rounak Dey, and Xihong Lin. Regression models for understanding COVID-19 epidemic dynamics with incomplete data. Journal of the American Statistical Association, 116(536):1561–1577, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See discussion [DM21, DY21] and rejoinder [QDL21b].

Quick:2021:RRM

[QDL21b] Corbin Quick, Rounak Dey, and Xihong Lin. Rejoinder: Regression Models for Understanding COVID-19 Epidemic Dynamics With Incomplete Data. Journal of the American Statistical Association, 116(536):1591–1594, 2021. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [QDL21a].

Qiu:2024:NER

[QDZ24] Jiaming Qiu, Xiongtao Dai, and Zhengyuan Zhu. Nonparametric estimation of repeated densities with heterogeneous sample sizes. Journal of the American Statistical Association, 119(545):176–188, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Qian:2024:AAM

[QIL24] Wei Qian, Ching-Kang Ing, and Ji Liu. Adaptive algorithm for multi-armed bandit problem with high-dimensional covariates. Journal of the American Statistical Association, 119 (546):970–982, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Qiu:2024:BRD

[Qiu24] Xing Qiu. Book review: Data Science and Predictive Analytics, 2nd ed. Ivo D. Dinov, Cham, Switzerland: Springer, 2023, xxxiv + 918 pp., $119.99(H), ISBN 978-3-031-17482-7. Journal of the American Statistical Association, 119(546): 1692–1693, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Quan:2024:OOP

[QL24] Mingxue Quan and Zhenhua Lin. Optimal one-pass nonparametric estimation under memory constraint. Journal of the American Statistical Association, 119(545):285–296, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Qi:2020:MAA

[QLFL20] Zhengling Qi, Dacheng Liu, Haoda Fu, and Yufeng Liu. Multiarmed angle-based direct learning for estimating optimal individualized treatment rules with various outcomes. Journal of the American Statistical Association, 115(530):678–691, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Qi:2024:PLI

[QMZ24] Zhengling Qi, Rui Miao, and Xiaoke Zhang. Proximal learning for individualized treatment regimes under unmeasured confounding. Journal of the American Statistical Association, 119(546):915–928, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Qi:2023:RID

[QPL23] Zhengling Qi, Jong-Shi Pang, and Yufeng Liu. On robustness of individualized decision rules. Journal of the American Statistical Association, 118(543):2143–2157, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Qiu:2021:AAL

[QW21] Yixuan Qiu and Xiao Wang. ALMOND: Adaptive latent modeling and optimization via neural networks and Langevin diffusion. Journal of the American Statistical Association, 116 (535):1224–1236, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Qiu:2024:EMS

[QW24] Yixuan Qiu and Xiao Wang. Efficient multimodal sampling via tempered distribution flow. Journal of the American Statistical Association, 119(546):1446–1460, 2024. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Qiu:2022:IML

[QZ22] Yumou Qiu and Xiao-Hua Zhou. Inference on multi-level partial correlations based on multi-subject time series data. Journal of the American Statistical Association, 117(540):2268– 2282, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rahnavard:2024:BRS

[Rah24] Ali Rahnavard. Book review: Statistical Analytics for Health Data Science with SAS and R Jeffrey R. Wilson, Ding-Geng Chen, and Karl E. Peace, Boca Raton, FL: Chapman & Hall/ CRC Press, 2023, xxi + 257 pp., $99.95(H), ISBN: 978-1-032-32562-0. Journal of the American Statistical Association, 119(545):786–787, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rabhi:2020:NIC

[RB20] Yassir Rabhi and Taoufik Bouezmarni. Nonparametric inference for copulas and measures of dependence under lengthbiased sampling and informative censoring. Journal of the American Statistical Association, 115(531):1268–1278, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rischard:2021:DSD

[RBMB21] Maxime Rischard, Zach Branson, Luke Miratrix, and Luke Bornn. Do school districts affect NYC house prices? Identifying border differences using a Bayesian nonparametric approach to geographic regression discontinuity designs. Journal of the American Statistical Association, 116(534):619–631, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Robin:2020:MEI

[RKJ^{+}20] Geneviève Robin, Olga Klopp, Julie Josse, Éric Moulines, and Robert Tibshirani. Main effects and interactions in mixed and incomplete data frames. Journal of the American Statistical Association, 115(531):1292–1303, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rashid:2021:HDP

[RLC^{+}21] Naim U. Rashid, Daniel J. Luckett, Jingxiang Chen, Michael T. Lawson, Longshaokan Wang, Yunshu Zhang, Eric B. Laber, Yufeng Liu, Jen Jen Yeh, Donglin Zeng, and Michael R. Kosorok. High-dimensional precision medicine from patient-derived xenografts. Journal of the American Statistical Association, 116(535):1140–1154, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Reluga:2023:SIE

[RLS23] Katarzyna Reluga, María-José Lombardía, and Stefan Sperlich. Simultaneous inference for empirical best predictors with a poverty study in small areas. Journal of the American Statistical Association, 118(541):583–595, 2023. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ramprasad:2023:OBI

[RLY^{+}23] Pratik Ramprasad, Yuantong Li, Zhuoran Yang, Zhaoran Wang, Will Wei Sun, and Guang Cheng. Online bootstrap inference for policy evaluation in reinforcement learning. Journal of the American Statistical Association, 118(544):2901– 2914, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rashid:2020:MBS

[RLYI20] Naim U. Rashid, Quefeng Li, Jen Jen Yeh, and Joseph G. Ibrahim. Modeling between-study heterogeneity for improved replicability in gene signature selection and clinical prediction. Journal of the American Statistical Association, 115 (531):1125–1138, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ricciardi:2020:BIS

[RMM20] Federico Ricciardi, Alessandra Mattei, and Fabrizia Mealli. Bayesian inference for sequential treatments under latent sequential ignorability. Journal of the American Statistical Association, 115(531):1498–1517, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Robert:2024:BRS

[Rob24] Christian P. Robert. Book review: Statistical Modeling with R: a Dual Frequentist and Bayesian Approach for Life Scientists, Pablo Inchausti, Oxford, UK: Oxford University Press, 2023, xvi + 463 pp., $53.00(P), ISBN 978-0-192-85902-0. Journal of the American Statistical Association, 119(546): 1691–1692, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Riva-Palacio:2022:SRM

[RPLG22] Alan Riva-Palacio, Fabrizio Leisen, and Jim Griffin. Survival regression models with dependent Bayesian nonparametric priors. Journal of the American Statistical Association, 117 (539):1530–1539, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rabinowicz:2022:CVC

[RR22] Assaf Rabinowicz and Saharon Rosset. Cross-validation for correlated data. Journal of the American Statistical Association, 117(538):718–731, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Romano:2022:DAC

[RRRF22] Gaetano Romano, Guillem Rigaill, Vincent Runge, and Paul Fearnhead. Detecting abrupt changes in the presence of local fluctuations and autocorrelated noise. Journal of the American Statistical Association, 117(540):2147–2162, 2022. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ray:2022:VBH

[RS22] Kolyan Ray and Botond Szabó. Variational Bayes for highdimensional linear regression with sparse priors. Journal of the American Statistical Association, 117(539):1270–1281, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Romano:2020:DK

[RSC20] Yaniv Romano, Matteo Sesia, and Emmanuel Candès. Deep knockoffs. Journal of the American Statistical Association, 115(532):1861–1872, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rudolph:2021:CSD

[RSvdL21] Kara E. Rudolph, Oleg Sofrygin, and Mark J. van der Laan. Complier stochastic direct effects: Identification and robust estimation. Journal of the American Statistical Association, 116(535):1254–1264, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rosset:2020:FRRa

[RT20a] Saharon Rosset and Ryan J. Tibshirani. From fixed-X to random-X regression: Bias-variance decompositions, covariance penalties, and prediction error estimation. Journal of the American Statistical Association, 115(529):138–151, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See discussion [SH20] and rejoinder [RT20b].

Rosset:2020:FRRb

[RT20b] Saharon Rosset and Ryan J. Tibshirani. From fixed-X to random-X regression: Bias-variance decompositions, covariance penalties, and prediction error estimation: Rejoinder. Journal of the American Statistical Association, 115(529): 161–162, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [RT20a, SH20, Ano22c].

Ruggeri:2021:CGS

[Rug21] Fabrizio Ruggeri. Comment on “A Gibbs Sampler for a Class of Random Convex Polytopes” by P. E. Jacob, R. Gong, P .T. Edlefsen and A. P. Dempster. Journal of the American Statistical Association, 116(535):1204–1205, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [JGED21a, JGED21b].

Ren:2023:DK

[RWC23] Zhimei Ren, Yuting Wei, and Emmanuel Candès. Derandomizing knockoffs. Journal of the American Statistical Association, 118(542):948–958, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ren:2022:LSD

[RZCL22] Haojie Ren, Changliang Zou, Nan Chen, and Runze Li. Largescale datastreams surveillance via pattern-oriented-sampling. Journal of the American Statistical Association, 117(538): 794–808, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Smith:2020:DMR

[SA20] Adam N. Smith and Greg M. Allenby. Demand models with random partitions. Journal of the American Statistical Association, 115(529):47–65, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Saha:2023:RFS

[SBD23] Arkajyoti Saha, Sumanta Basu, and Abhirup Datta. Random forests for spatially dependent data. Journal of the American Statistical Association, 118(541):665–683, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shin:2020:FHP

[SBJ20] Minsuk Shin, Anirban Bhattacharya, and Valen E. Johnson. Functional horseshoe priors for subspace shrinkage. Journal of the American Statistical Association, 115(532):1784–1797, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Saha:2020:GVA

[SBK20] Abhijoy Saha, Karthik Bharath, and Sebastian Kurtek. A geometric variational approach to Bayesian inference. Journal of the American Statistical Association, 115(530):822–835, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schultheiss:2024:HOL

[SBY24] Christoph Schultheiss, Peter Bühlmann, and Ming Yuan. Higher-order least squares: Assessing partial goodness of fit of linear causal models. Journal of the American Statistical Association, 119(546):1019–1031, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stoepker:2024:ADL

[SCACvdH24] Ivo V. Stoepker, Rui M. Castro, Ery Arias-Castro, and Edwin van den Heuvel. Anomaly detection for a large number of streams: a permutation-based higher criticism approach. Journal of the American Statistical Association, 119(545): 461–474, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schoenberg:2020:BRT

[Sch20] Frederic P. Schoenberg. Book review: Theory of Spatial Statistics: a Concise Introduction, by M. N. M. van Lieshout. Boca Raton, FL: Chapman & Hall/CRC Press, 2019, xiii + 168 pp., $69.95(P), ISBN: 978-0-36-714639-9. Journal of the American Statistical Association, 115(530):1033–1034, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Song:2024:SMM

[SCMW24] Fangda Song, Jing Chu, Shuangge Ma, and Yingying Wei. Survival mixed membership blockmodel. Journal of the American Statistical Association, 119(546):1647–1656, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shi:2022:DFC

[SDH22] Hongjian Shi, Mathias Drton, and Fang Han. Distributionfree consistent independence tests via center-outward ranks and signs. Journal of the American Statistical Association, 117(537):395–410, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sen:2024:BRM

[Sen24] Bodhisattva Sen. Book review: Mathematical Foundations of Infinite-Dimensional Statistical Models, Evarist Giné and Richard Nickl, New York, NY: Cambridge University Press, 2021, xiv + 690 pp., $52.99(P), ISBN 978-1-108-99413-2. Journal of the American Statistical Association, 119(546): 1687–1689, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Seo:2024:BRM

[Seo24] Insuk Seo. Book review: Martingale Methods in Statistics Yoichi Nishiyama, Boca Raton, FL: Chapman & Hall/CRC Press, 2022, xiii + 245 pp., $120.00(H), ISBN: 978-1-466-58281-1. Journal of the American Statistical Association, 119(545):787–789, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schweinberger:2023:DTT

[SF23] Michael Schweinberger and Cornelius Fritz. Discussion of “A Tale of Two Datasets: Representativeness and Generalisability of Inference for Samples of Networks” by Pavel N. Krivitsky, Pietro Coletti, and Niel Hens. Journal of the American Statistical Association, 118(544):2225–2227, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [KCH23b, KCH23a].

Shen:2020:DFR

[SH20] Xiaotong Shen and Hsin-Cheng Huang. Discussion of “From Fixed-X to Random-X Regression: Bias-Variance Decompositions, Covariance Penalties, and Prediction Error Estimation”. Journal of the American Statistical Association, 115 (529):152–156, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [RT20a, RT20b].

Shafer:2021:CGS

[Sha21] Glenn Shafer. Comment on “A Gibbs Sampler for a Class of Random Convex Polytopes,” by Pierre E. Jacob, Ruobin Gong, Paul T. Edlefsen, and Arthur P. Dempster. Journal of the American Statistical Association, 116(535):1196–1197, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stephenson:2020:RCS

[SHO20] Briana J. K. Stephenson, Amy H. Herring, and Andrew Olshan. Robust clustering with subpopulation-specific deviations. Journal of the American Statistical Association, 115 (530):521–537, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sung:2020:CCE

[SHR^{+}20a] Chih-Li Sung, Ying Hung, William Rittase, Cheng Zhu, and C. F. J. Wu. Calibration for computer experiments with binary responses and application to cell adhesion study. Journal of the American Statistical Association, 115(532):1664–1674, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sung:2020:GGP

[SHR^{+}20b] Chih-Li Sung, Ying Hung, William Rittase, Cheng Zhu, and C. F. Jeff Wu. A generalized Gaussian process model for computer experiments with binary time series. Journal of the American Statistical Association, 115(530):945–956, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Simpson:2023:IPD

[SHWB23] Matthew Simpson, Scott H. Holan, Christopher K. Wikle, and Jonathan R. Bradley. Interpolating population distributions using public-use data: an application to income segregation using American community survey data. Journal of the American Statistical Association, 118(541):84–96, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Song:2023:IMM

[SJW23] Xiaoyu Song, Jiayi Ji, and Pei Wang. iProMix: a mixture model for studying the function of ACE2 based on bulk proteogenomic data. Journal of the American Statistical Association, 118(541):43–55, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sang:2022:SFI

[SKL22] Hejian Sang, Jae Kwang Kim, and Danhyang Lee. Semiparametric fractional imputation using Gaussian mixture models for handling multivariate missing data. Journal of the American Statistical Association, 117(538):654–663, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sun:2020:GVS

[SL20] Ryan Sun and Xihong Lin. Genetic variant set-based tests using the generalized Berk–Jones statistic with application to a genome-wide association study of breast cancer. Journal of the American Statistical Association, 115(531):1079–1091, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shi:2022:TME

[SL22a] Chengchun Shi and Lexin Li. Testing mediation effects using logic of Boolean matrices. Journal of the American Statistical Association, 117(540):2014–2027, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shi:2022:CRC

[SL22b] Peng Shi and Gee Y. Lee. Copula regression for compound distributions with endogenous covariates with applications in insurance deductible pricing. Journal of the American Statistical Association, 117(539):1094–1109, 2022. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shin:2022:NPB

[SL22c] Minsuk Shin and Jun S. Liu. Neuronized priors for Bayesian sparse linear regression. Journal of the American Statistical Association, 117(540):1695–1710, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Solea:2022:CGG

[SL22d] Eftychia Solea and Bing Li. Copula Gaussian graphical models for functional data. Journal of the American Statistical Association, 117(538):781–793, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shi:2021:SRU

[SLC21] Xu Shi, Xiaoou Li, and Tianxi Cai. Spherical regression under mismatch corruption with application to automated knowledge translation. Journal of the American Statistical Association, 116(536):1953–1964, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Salvana:2023:STC

[SLG23] Mary Lai O. Salvaña, Amanda Lenzi, and Marc G. Genton. Spatio-temporal cross-covariance functions under the Lagrangian framework with multiple advections. Journal of the American Statistical Association, 118(544):2746–2761, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shen:2020:FIF

[SLgX20] Jieli Shen, Regina Y. Liu, and Min ge Xie. i fusion: Individualized fusion learning. Journal of the American Statistical Association, 115(531):1251–1267, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shi:2024:SEA

[SLL^{+}24] Chengchun Shi, Shikai Luo, Yuan Le, Hongtu Zhu, and Rui Song. Statistically efficient advantage learning for offline reinforcement learning in infinite horizons. Journal of the American Statistical Association, 119(545):232–245, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shi:2020:SRP

[SLS20] Chengchun Shi, Wenbin Lu, and Rui Song. A sparse random projection-based test for overall qualitative treatment effects. Journal of the American Statistical Association, 115 (531):1201–1213, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Su:2020:TES

[SLSH20] Lin Su, Wenbin Lu, Rui Song, and Danyang Huang. Testing and estimation of social network dependence with time to event data. Journal of the American Statistical Association, 115(530):570–582, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Song:2024:GET

[SLZ24] Shanshan Song, Yuanyuan Lin, and Yong Zhou. A general M-estimation theory in semi-supervised framework. Journal of the American Statistical Association, 119(546):1065–1075, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schulz:2021:DRE

[SM21] Juliana Schulz and Erica E. M. Moodie. Doubly robust estimation of optimal dosing strategies. Journal of the American Statistical Association, 116(533):256–268, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sun:2021:REA

[SMMH21] Yifei Sun, Charles E. McCulloch, Kieren A. Marr, and Chiung-Yu Huang. Recurrent events analysis with data collected at informative clinical visits in electronic health records. Journal of the American Statistical Association, 116 (534):594–604, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Simoneau:2020:EOD

[SMN^{+}20] Gabrielle Simoneau, Erica E. M. Moodie, Jagtar S. Nijjar, Robert W. Platt, and the Scottish Early Rheumatoid Arthritis Inception Cohort Investigators. Estimating optimal dynamic treatment regimes with survival outcomes. Journal of the American Statistical Association, 115(531):1531–1539, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Scheike:2023:EEF

[SMO23] Thomas H. Scheike, Torben Martinussen, and Brice Ozenne. Efficient estimation in the Fine and Gray model. Journal of the American Statistical Association, 118(544):2482–2490, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sarkar:2021:BCD

[SPMC21] Abhra Sarkar, Debdeep Pati, Bani K. Mallick, and Raymond J. Carroll. Bayesian copula density deconvolution for zero-inflated data in nutritional epidemiology. Journal of the American Statistical Association, 116(535):1075–1087, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schochet:2022:DBR

[SPMK22] Peter Z. Schochet, Nicole E. Pashley, Luke W. Miratrix, and Tim Kautz. Design-based ratio estimators and central limit theorems for clustered, blocked RCTs. Journal of the American Statistical Association, 117(540):2135–2146, 2022. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shen:2020:DCM

[SPV20] Cencheng Shen, Carey E. Priebe, and Joshua T. Vogelstein. From distance correlation to multiscale graph correlation. Journal of the American Statistical Association, 115(529): 280–291, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Su:2021:EBC

[SQPQ21] Qihui Su, Zhongling Qin, Liang Peng, and Gengsheng Qin. Efficiently backtesting conditional value-at-risk and conditional expected shortfall. Journal of the American Statistical Association, 116(536):2041–2052, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Song:2020:PHD

[SR20] Hyebin Song and Garvesh Raskutti. PUlasso: Highdimensional variable selection with presence-only data. Journal of the American Statistical Association, 115(529):334– 347, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stensrud:2023:CSE

[SRS^{+}23] Mats J. Stensrud, James M. Robins, Aaron Sarvet, Eric J. Tchetgen Tchetgen, and Jessica G. Young. Conditional separable effects. Journal of the American Statistical Association, 118(544):2671–2683, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Safikhani:2022:JSB

[SS22] Abolfazl Safikhani and Ali Shojaie. Joint structural break detection and parameter estimation in high-dimensional nonstationary VAR models. Journal of the American Statistical Association, 117(537):251–264, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sansom:2021:CPF

[SSB21] Philip G. Sansom, David B. Stephenson, and Thomas J. Bracegirdle. On constraining projections of future climate using observations and simulations from multiple climate models. Journal of the American Statistical Association, 116(534): 546–557, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sun:2022:CSD

[SSL22] Yan Sun, Qifan Song, and Faming Liang. Consistent sparse deep learning: Theory and computation. Journal of the American Statistical Association, 117(540):1981–1995, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shi:2021:SIH

[SSLL21] Chengchun Shi, Rui Song, Wenbin Lu, and Runze Li. Statistical inference for high-dimensional models via recursive online-score estimation. Journal of the American Statistical Association, 116(535):1307–1318, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sekhon:2021:INC

[SST21] Jasjeet S. Sekhon and Yotam Shem-Tov. Inference on a new class of sample average treatment effects. Journal of the American Statistical Association, 116(534):798–804, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shaikh:2021:RTO

[ST21] Azeem M. Shaikh and Panos Toulis. Randomization tests in observational studies with staggered adoption of treatment. Journal of the American Statistical Association, 116 (536):1835–1848, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Song:2023:HAM

[ST23a] Zexi Song and Zhiqiang Tan. Hamiltonian-assisted Metropolis sampling. Journal of the American Statistical Association, 118(542):1176–1194, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sun:2023:UPD

[ST23b] Cheng-Yu Sun and Boxin Tang. Uniform projection designs and strong orthogonal arrays. Journal of the American Statistical Association, 118(541):417–423, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Storlie:2020:PIM

[STC^{+}20] Curtis B. Storlie, Terry M. Therneau, Rickey E. Carter, Nicholas Chia, John R. Bergquist, Jeanne M. Huddleston, and Santiago Romero-Brufau. Prediction and inference with missing data in patient alert systems. Journal of the American Statistical Association, 115(529):32–46, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stein:2023:EWM

[Ste23] Michael L. Stein. Editorial: What makes for a great applications and case studies paper? Journal of the American Statistical Association, 118(541):1–2, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stokes:2022:BRS

[Sto22] S. Lynne Stokes. Book review: Sampling: Design and Analysis, 3rd ed. Sharon L. Lohr, Boca Raton, FL: Chapman & Hall/CRC Press, 2022, xxiii + 650 pp., $79.95(H), ISBN 978-0-367-27950-9. Journal of the American Statistical Association, 117(540):2287–2288, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sugasawa:2021:GHM

[Sug21] Shonosuke Sugasawa. Grouped heterogeneous mixture modeling for clustered data. Journal of the American Statistical Association, 116(534):999–1010, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Scealy:2021:ASA

[SW21a] Janice L. Scealy and Andrew T. A. Wood. Analogues on the sphere of the affine-equivariant spatial median. Journal of the American Statistical Association, 116(535):1457–1471, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sun:2021:STS

[SW21b] Yilun Sun and Lu Wang. Stochastic tree search for estimating optimal dynamic treatment regimes. Journal of the American Statistical Association, 116(533):421–432, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Scealy:2023:SMC

[SW23] Janice L. Scealy and Andrew T. A. Wood. Score matching for compositional distributions. Journal of the American Statistical Association, 118(543):1811–1823, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shi:2023:DCE

[SWL^{+}23] Chengchun Shi, Xiaoyu Wang, Shikai Luo, Hongtu Zhu, Jieping Ye, and Rui Song. Dynamic causal effects evaluation in A/B testing with a reinforcement learning framework. Journal of the American Statistical Association, 118 (543):2059–2071, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sung:2020:MFA

[SWPH20] Chih-Li Sung, Wenjia Wang, Matthew Plumlee, and Benjamin Haaland. Multiresolution functional ANOVA for largescale, many-input computer experiments. Journal of the American Statistical Association, 115(530):908–919, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schnell:2024:ORT

[SWR24] Patrick M. Schnell, Matthew Wascher, and Grzegorz A. Rempala. Overcoming repeated testing schedule bias in estimates of disease prevalence. Journal of the American Statistical Association, 119(545):1–13, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

She:2022:GOR

[SWS22] Yiyuan She, Zhifeng Wang, and Jiahui Shen. Gaining outlier resistance with progressive quantiles: Fast algorithms and theoretical studies. Journal of the American Statistical Association, 117(539):1282–1295, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shu:2020:DCD

[SWZ20] Hai Shu, Xiao Wang, and Hongtu Zhu. D-CCA: a decomposition-based canonical correlation analysis for highdimensional datasets. Journal of the American Statistical Association, 115(529):292–306, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stensrud:2022:SEC

[SYD^{+}22] Mats J. Stensrud, Jessica G. Young, Vanessa Didelez, James M. Robins, and Miguel A. Hernán. Separable effects for causal inference in the presence of competing events. Journal of the American Statistical Association, 117(537):175–183, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sun:2021:TII

[SZ21] Qiang Sun and Heping Zhang. Targeted inference involving high-dimensional data using nuisance penalized regression. Journal of the American Statistical Association, 116 (535):1472–1486, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sun:2020:AHR

[SZF20] Qiang Sun, Wen-Xin Zhou, and Jianqing Fan. Adaptive Huber regression. Journal of the American Statistical Association, 115(529):254–265, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shi:2024:PCI

[SZS^{+}24a] Chengchun Shi, Jin Zhu, Ye Shen, Shikai Luo, Hongtu Zhu, and Rui Song. Off-policy confidence interval estimation with confounded Markov decision process. Journal of the American Statistical Association, 119(545):273–284, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shi:2024:LWD

[SZS24b] Peng Shi, Wei Zhang, and Kun Shi. Leveraging weather dynamics in insurance claims triage using deep learning. Journal of the American Statistical Association, 119(546):825–838, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Su:2024:ECT

[SZZ24] Chang Su, Jingfei Zhang, and Hongyu Zhao. Estimating cell-type-specific gene co-expression networks from bulk gene expression data with an application to alzheimer’s disease. Journal of the American Statistical Association, 119(546): 811–824, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tabouy:2020:VIS

[TBC20] Timothée Tabouy, Pierre Barbillon, and Julien Chiquet. Variational inference for stochastic block models from sampled data. Journal of the American Statistical Association, 115 (529):455–466, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tang:2020:IMT

[TBQ20] Xiwei Tang, Xuan Bi, and Annie Qu. Individualized multilayer tensor learning with an application in imaging analysis. Journal of the American Statistical Association, 115(530): 836–851, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tong:2023:MFC

[TCYL23] Zhaoxue Tong, Zhanrui Cai, Songshan Yang, and Runze Li. Model-free conditional feature screening with FDR control. Journal of the American Statistical Association, 118(544): 2575–2587, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Trippe:2024:CCE

[TDB24] Brian L. Trippe, Sameer K. Deshpande, and Tamara Broderick. Confidently comparing estimates with the c-value. Journal of the American Statistical Association, 119(546):983– 994, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tendijck:2023:MEB

[TET^{+}23] Stan Tendijck, Emma Eastoe, Jonathan Tawn, David Randell, and Philip Jonathan. Modeling the extremes of bivariate mixture distributions with application to oceanographic data. Journal of the American Statistical Association, 118 (542):1373–1384, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tian:2023:RVS

[TF23a] Ye Tian and Yang Feng. RaSE: a variable screening framework via random subspace ensembles. Journal of the American Statistical Association, 118(541):457–468, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tian:2023:TLU

[TF23b] Ye Tian and Yang Feng. Transfer learning under highdimensional generalized linear models. Journal of the American Statistical Association, 118(544):2684–2697, 2023. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tang:2021:IDV

[TFCF21] Francesca Tang, Yang Feng, Hamza Chiheb, and Jianqing Fan. The interplay of demographic variables and social distancing scores in deep prediction of U.S. COVID-19 cases. Journal of the American Statistical Association, 116(534): 492–506, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tchetgen:2021:AGC

[TFS21] Eric J. Tchetgen Tchetgen, Isabel R. Fulcher, and Ilya Shpitser. Auto-G-computation of causal effects on a network. Journal of the American Statistical Association, 116(534): 833–844, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tuo:2023:RKH

[THP^{+}23] Rui Tuo, Shiyuan He, Arash Pourhabib, Yu Ding, and Jianhua Z. Huang. A reproducing kernel Hilbert space approach to functional calibration of computer models. Journal of the American Statistical Association, 118(542):883–897, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tang:2023:SAO

[THXZ23] Weijing Tang, Kevin He, Gongjun Xu, and Ji Zhu. Survival analysis via ordinary differential equations. Journal of the American Statistical Association, 118(544):2406–2421, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tokdar:2024:HTD

[TJC24] Surya T. Tokdar, Sheng Jiang, and Erika L. Cunningham. Heavy-tailed density estimation. Journal of the American Statistical Association, 119(545):163–175, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tian:2023:CGL

[TJH^{+}23] Chuan Tian, Duo Jiang, Austin Hammer, Thomas Sharpton, and Yuan Jiang. Compositional graphical lasso resolves the impact of parasitic infection on gut microbial interaction networks in a zebrafish model. Journal of the American Statistical Association, 118(543):1500–1514, 2023. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tu:2022:CMH

[TJK22] Wei Tu, Bei Jiang, and Linglong Kong. Comments on “Measuring Housing Vitality from Multi-Source Big Data and Machine Learning”. Journal of the American Statistical Association, 117(539):1060–1062, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [ZXS^{+}22a] and rejoinder [ZXS^{+}22b]..

Tang:2023:MTP

[TL23] Xiwei Tang and Lexin Li. Multivariate temporal point process regression. Journal of the American Statistical Association, 118(542):830–845, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tian:2024:AMD

[TLT^{+}24] Yuqi Tian, Chun Li, Shengxin Tu, Nathan T. James, FrankE. Harrell, and BryanE. Shepherd. Addressing multiple detection limits with semiparametric cumulative probability models. Journal of the American Statistical Association, 119(546): 864–874, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tao:2024:AGM

[TLX24] Jun Tao, Bing Li, and Lingzhou Xue. An additive graphical model for discrete data. Journal of the American Statistical Association, 119(545):368–381, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tian:2022:PNF

[TMNM22] Qinglong Tian, Fanqi Meng, Daniel J. Nordman, and William Q. Meeker. Predicting the number of future events. Journal of the American Statistical Association, 117(539): 1296–1310, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tenzer:2022:TIU

[TMZ22] Yaniv Tenzer, Micha Mandel, and Or Zuk. Testing independence under biased sampling. Journal of the American Statistical Association, 117(540):2194–2206, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tan:2023:SRR

[TSW23] Kean Ming Tan, Qiang Sun, and Daniela Witten. Sparse reduced rank Huber regression in high dimensions. Journal of the American Statistical Association, 118(544):2383–2393, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tian:2021:ESM

[TTL^{+}21] Ting Tian, Jianbin Tan, Wenxiang Luo, Yukang Jiang, Minqiong Chen, Songpan Yang, Canhong Wen, Wenliang Pan, and Xueqin Wang. The effects of stringent and mild interventions for coronavirus pandemic. Journal of the American Statistical Association, 116(534):481–491, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Thomas:2020:EEF

[TTPD20] Emma G. Thomas, Lorenzo Trippa, Giovanni Parmigiani, and Francesca Dominici. Estimating the effects of fine particulate matter on 432 cardiovascular diseases using multioutcome regression with tree-structured shrinkage. Journal of the American Statistical Association, 115(532):1689–1699, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tucker:2023:VSG

[TWM23] Danielle C. Tucker, Yichao Wu, and Hans-Georg Müller. Variable selection for global Fréchet regression. Journal of the American Statistical Association, 118(542):1023–1037, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tang:2021:IMV

[TXQ21] Xiwei Tang, Fei Xue, and Annie Qu. Individualized multidirectional variable selection. Journal of the American Statistical Association, 116(535):1280–1296, 2021. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Teng:2024:TWT

[TZ24] Hao Yang Teng and Zhengjun Zhang. Two-way truncated linear regression models with extremely thresholding penalization. Journal of the American Statistical Association, 119 (546):887–903, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tao:2020:ODT

[TZL20] Ran Tao, Donglin Zeng, and Dan-Yu Lin. Optimal designs of two-phase studies. Journal of the American Statistical Association, 115(532):1946–1959, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Vatter:2020:SCS

[Vat20] Thibault Vatter. Simulating copulas: Stochastic models, sampling algorithms, and applications. Journal of the American Statistical Association, 115(529):481–482, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Vazquez-Bare:2023:CSE

[VB23] Gonzalo Vazquez-Bare. Causal spillover effects using instrumental variables. Journal of the American Statistical Association, 118(543):1911–1922, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Viviano:2024:FPT

[VB24] Davide Viviano and Jelena Bradic. Fair policy targeting. Journal of the American Statistical Association, 119(545): 730–743, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Vanegas:2022:MQS

[VBM22] Laura Jula Vanegas, Merle Behr, and Axel Munk. Multiscale quantile segmentation. Journal of the American Statistical Association, 117(539):1384–1397, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Vansteelandt:2021:DKM

[VD21] Stijn Vansteelandt and Oliver Dukes. Discussion of Kallus and Mo, Qi, and Liu: New objectives for policy learning. Journal of the American Statistical Association, 116(534):

675–679, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Kal21a, MQL21a].

Vansteelandt:2024:ALC

[VDVM24] Stijn Vansteelandt, Oliver Dukes, Kelly Van Lancker, and Torben Martinussen. Assumption-lean Cox regression. Journal of the American Statistical Association, 119(545):475– 484, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wager:2020:CVR

[Wag20] Stefan Wager. Cross-validation, risk estimation, and model selection: Comment on a paper by Rosset and tibshirani. Journal of the American Statistical Association, 115(529): 157–160, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2020:BRM

[Wan20] Qing Wang. Book review: Multivariate Kernel Smoothing and Its Applications, by José E. Chacón and Tarn Duong. Boca Raton, FL: Chapman & Hall/CRC Press, 2018, xxi + 226 pp., $99.95 (H), ISBN: 978-1-49-876301-1. Journal of the American Statistical Association, 115(529):486, 2020. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2023:SSM

[Wan23] Shulei Wang. Self-supervised metric learning in multi-view data: a downstream task perspective. Journal of the American Statistical Association, 118(544):2454–2467, 2023. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wilms:2023:SIE

[WBBM23] Ines Wilms, Sumanta Basu, Jacob Bien, and David S. Matteson. Sparse identification and estimation of large-scale vector AutoRegressive moving averages. Journal of the American Statistical Association, 118(541):571–582, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2022:BER

[WBK^{+}22] Zeya Wang, Veerabhadran Baladandayuthapani, Ahmed O. Kaseb, Hesham M. Amin, Manal M. Hassan, Wenyi Wang, and Jeffrey S. Morris. Bayesian edge regression in undirected graphical models to characterize interpatient heterogeneity in cancer. Journal of the American Statistical Association, 117(538):533–546, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wei:2024:SCD

[WC24] Zeyu Wei and Yen-Chi Chen. Skeleton clustering: Dimensionfree density-aided clustering. Journal of the American Statistical Association, 119(546):1124–1135, 2024. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2021:OEW

[WCL21] Shulei Wang, T. Tony Cai, and Hongzhe Li. Optimal estimation of Wasserstein distance on a tree with an application to microbiome studies. Journal of the American Statistical Association, 116(535):1237–1253, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2024:VSH

[WCNL24] Jia Wang, Xizhen Cai, Xiaoyue Niu, and Runze Li. Variable selection for high-dimensional nodal attributes in social networks with degree heterogeneity. Journal of the American Statistical Association, 119(546):1322–1335, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2024:HAM

[WCZZ24] Shiying Wang, Todd Constable, Heping Zhang, and Yize Zhao. Heterogeneity analysis on multi-state brain functional connectivity and adolescent neurocognition. Journal of the American Statistical Association, 119(546):851–863, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wu:2021:RTW

[WD21] Jason Wu and Peng Ding. Randomization tests for weak null hypotheses in randomized experiments. Journal of the American Statistical Association, 116(536):1898–1913, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Weiss:2020:DBA

[Wei20] Christian H. Weiß. Distance-based analysis of ordinal data and ordinal time series. Journal of the American Statistical Association, 115(531):1189–1200, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Westling:2022:NTC

[Wes22] Ted Westling. Nonparametric tests of the causal null with nondiscrete exposures. Journal of the American Statistical Association, 117(539):1551–1562, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2022:BRS

[WFL22] Zixiao Wang, Yi Feng, and Lin Liu. Book review: Semiparametric Regression with R Jaroslaw Harezlak, David Ruppert, and Matt P. Wand, New York, NY: Springer, 2018, xi + 331 pp., $149.99(P), $109.00(e-Book), ISBN 978-1-4939-8851-8(P)/978-1-4939-8853-2(e-Book). Journal of the American Statistical Association, 117(540):2283–2287, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wu:2024:BSB

[WGK24] Ben Wu, Ying Guo, and Jian Kang. Bayesian spatial blind source separation via the thresholded Gaussian process. Journal of the American Statistical Association, 119(545):422– 433, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Williamson:2023:GFI

[WGSC23] Brian D. Williamson, Peter B. Gilbert, Noah R. Simon, and Marco Carone. A general framework for inference on algorithm-agnostic variable importance. Journal of the American Statistical Association, 118(543):1645–1658, 2023. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wrobel:2024:PAE

[WHC^{+}24] Julia Wrobel, Emily C. Hector, Lorin Crawford, Lucy D’Agostino McGowan, Natalia da Silva, Jeff Goldsmith, Stephanie Hicks, Michael Kane, Youjin Lee, Vinicius Mayrink, Christopher J. Paciorek, Therri Usher, and Julian Wolfson. Partnering with authors to enhance reproducibility at JASA. Journal of the American Statistical Association, 119(546):795–797, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wikle:2022:MMA

[WHHZ22] Nathan B. Wikle, Ephraim M. Hanks, Lucas R. F. Henneman, and Corwin M. Zigler. A mechanistic model of annual sulfate concentrations in the United States. Journal of the American Statistical Association, 117(539):1082–1093, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2020:DIT

[WHX20] Jingshen Wang, Xuming He, and Gongjun Xu. Debiased inference on treatment effect in a high-dimensional model. Journal of the American Statistical Association, 115(529): 442–454, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Williams:2021:DGS

[Wil21] Jonathan P. Williams. Discussion of “A Gibbs Sampler for a Class of Random Convex Polytopes”. Journal of the American Statistical Association, 116(535):1198–1200, 2021. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [JGED21a, JGED21b].

Wilson:2020:MTS

[WIS20] DouglasR. Wilson, JosephG. Ibrahim, and Wei Sun. Mapping tumor-specific expression QTLs in impure tumor samples. Journal of the American Statistical Association, 115(529):79– 89, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wilson:2020:IPA

[WJIS20] Douglas R. Wilson, Chong Jin, Joseph G. Ibrahim, and Wei Sun. ICeD-T provides accurate estimates of immune cell abundance in tumor samples by allowing for aberrant gene expression patterns. Journal of the American Statistical Association, 115(531):1055–1065, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2022:WBS

[WJM22] Lazhi Wang, David E. Jones, and Xiao-Li Meng. Warp bridge sampling: The next generation. Journal of the American Statistical Association, 117(538):835–851, 2022. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2021:TOF

[WJW^{+}21] Zhuo Wang, Yujing Jiang, Hui Wan, Jun Yan, and Xuebin Zhang. Toward optimal fingerprinting in detection and attribution of changes in climate extremes. Journal of the American Statistical Association, 116(533):1–13, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Westling:2024:ITS

[WLGC24] Ted Westling, Alex Luedtke, Peter B. Gilbert, and Marco Carone. Inference for treatment-specific survival curves using machine learning. Journal of the American Statistical Association, 119(546):1541–1553, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wu:2024:MGP

[WMK^{+}24] Xiao Wu, Fabrizia Mealli, Marianthi-Anna Kioumourtzoglou, Francesca Dominici, and Danielle Braun. Matching on generalized propensity scores with continuous exposures. Journal of the American Statistical Association, 119(545):757–772, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2024:KEB

[WNK24] Yue Wang, Bin Nan, and John D. Kalbfleisch. Kernel estimation of bivariate time-varying coefficient model for longitudinal data with terminal event. Journal of the American Statistical Association, 119(546):1102–1111, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wong:2024:BRH

[Won24] Raymond K. W. Wong. Book review: Handbook of Matching and Weighting Adjustments for Causal Inference, José R. Zubizarreta, Elizabeth A. Stuart, Dylan S. Small, and Paul R. Rosenbaum, eds., Boca Raton, FL: Chapman & Hall/CRC Press, 2023, xi + 621 pp., $250.00(H), ISBN: 978-0-367-60952-8. Journal of the American Statistical Association, 119(545): 791, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2020:RTF

[WPB^{+}20a] Lan Wang, Bo Peng, Jelena Bradic, Runze Li, and Yunan Wu. Rejoinder to “A tuning-free robust and efficient approach to high-dimensional regression”. Journal of the American Statistical Association, 115(532):1726–1729, 2020. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [WPB^{+}20b].

Wang:2020:TFR

[WPB^{+}20b] Lan Wang, Bo Peng, Jelena Bradic, Runze Li, and Yunan Wu. A tuning-free robust and efficient approach to highdimensional regression. Journal of the American Statistical Association, 115(532):1700–1714, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See comments [Loh20, LS20, FMW20] and rejoinder [WPB^{+}20a].

Weitkamp:2024:DDB

[WPTM24] Christoph Alexander Weitkamp, Katharina Proksch, Carla Tameling, and Axel Munk. Distribution of distances based object matching: Asymptotic inference. Journal of the American Statistical Association, 119(545):538–551, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2023:SVS

[WQ23] Wenbo Wang and Xingye Qiao. Set-valued support vector machine with bounded error rates. Journal of the American Statistical Association, 118(544):2847–2859, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2023:TEA

[WRG23] Tianhao Wang, Sarah J. Ratcliffe, and Wensheng Guo. Timeto-event analysis with unknown time origins via longitudinal biomarker registration. Journal of the American Statistical Association, 118(543):1968–1983, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2023:MRI

[WSM^{+}23] Bingkai Wang, Ryoko Susukida, Ramin Mojtabai, Masoumeh Amin-Esmaeili, and Michael Rosenblum. Model-robust inference for clinical trials that improve precision by stratified randomization and covariate adjustment. Journal of the American Statistical Association, 118(542):1152–1163, 2023.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Williams:2020:BAM

[WST^{+}20] Jonathan P. Williams, Curtis B. Storlie, Terry M. Therneau, Clifford R. Jack, Jr., and Jan Hannig. A Bayesian approach to multistate hidden Markov models: Application to dementia progression. Journal of the American Statistical Association, 115(529):16–31, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2022:DOM

[WSX22] Yaping Wang, Fasheng Sun, and Hongquan Xu. On design orthogonality, maximin distance, and projection uniformity for computer experiments. Journal of the American Statistical Association, 117(537):375–385, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2020:PPK

[WTW20] Wenjia Wang, Rui Tuo, and C. F. Jeff Wu. On prediction properties of kriging: Uniform error bounds and robustness. Journal of the American Statistical Association, 115(530): 920–930, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Waite:2022:MER

[WW22] Timothy W. Waite and David C. Woods. Minimax efficient random experimental design strategies with application to model-robust design for prediction. Journal of the American Statistical Association, 117(539):1452–1465, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wu:2023:MAU

[WWF23] Yunan Wu, Lan Wang, and Haoda Fu. Model-assisted uniformly honest inference for optimal treatment regimes in high dimension. Journal of the American Statistical Association, 118(541):305–314, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2023:RSR

[WWL23] Xinhe Wang, Tingyu Wang, and Hanzhong Liu. Rerandomization in stratified randomized experiments. Journal of the American Statistical Association, 118(542):1295–1304, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2024:HNP

[WWLT24] Lijia Wang, Y. X. Rachel Wang, Jingyi Jessica Li, and Xin Tong. Hierarchical Neyman–Pearson classification for prioritizing severe disease categories in COVID-19 patient data. Journal of the American Statistical Association, 119(545):39– 51, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wu:2022:MBM

[WWW22] Tung-Yu Wu, Y. X. Rachel Wang, and Wing H. Wong. Minibatch Metropolis–Hastings with reversible SGLD proposal. Journal of the American Statistical Association, 117(537): 386–394, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2022:LRC

[WWZ22] Jiayi Wang, Raymond K. W. Wong, and Xiaoke Zhang. Lowrank covariance function estimation for multidimensional functional data. Journal of the American Statistical Association, 117(538):809–822, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wu:2023:STM

[WZC23] Ruijia Wu, Linjun Zhang, and T. Tony Cai. Sparse topic modeling: Computational efficiency, near-optimal algorithms, and statistical inference. Journal of the American Statistical Association, 118(543):1849–1861, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2023:FNC

[WZL^{+}23] Jiangzhou Wang, Jingfei Zhang, Binghui Liu, Ji Zhu, and Jianhua Guo. Fast network community detection with profilepseudo likelihood methods. Journal of the American Statistical Association, 118(542):1359–1372, 2023. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2022:HDV

[WZLL22] Di Wang, Yao Zheng, Heng Lian, and Guodong Li. Highdimensional vector autoregressive time series modeling via tensor decomposition. Journal of the American Statistical Association, 117(539):1338–1356, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wu:2020:MLO

[WZW20] Peng Wu, Donglin Zeng, and Yuanjia Wang. Matched learning for optimizing individualized treatment strategies using electronic health records. Journal of the American Statistical Association, 115(529):380–392, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xia:2023:ISE

[XC23a] Fan Xia and Kwun Chuen Gary Chan. Identification, semiparametric efficiency, and quadruply robust estimation in mediation analysis with treatment-induced confounding. Journal of the American Statistical Association, 118(542):1272–1281, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xia:2023:DSF

[XC23b] Yin Xia and T. Tony Cai. Discussion of “A Scale-Free Approach for False Discovery Rate Control in Generalized Linear Models” by Dai, Lin, Xing, and Liu. Journal of the American Statistical Association, 118(543):1569–1572, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [DLXL23c].

Xia:2020:GGF

[XCS20] Yin Xia, T. Tony Cai, and Wenguang Sun. GAP: a general framework for information pooling in two-sample sparse inference. Journal of the American Statistical Association, 115 (531):1236–1250, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xu:2023:MRS

[XCZW23] Tianchen Xu, Yuan Chen, Donglin Zeng, and Yuanjia Wang. Mixed-response state-space model for analyzing multidimensional digital phenotypes. Journal of the American Statistical Association, 118(544):2288–2300, 2023. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xu:2024:TSD

[XDM24] Yuchen Xu, Marie-Christine Düker, and David S. Matteson. Testing simultaneous diagonalizability. Journal of the American Statistical Association, 119(546):1513–1525, 2024. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xia:2020:SCI

[XLLJ20] Yin Xia, Lexin Li, Samuel N. Lockhart, and William J. Jagust. Simultaneous covariance inference for multimodal integrative analysis. Journal of the American Statistical Association, 115(531):1279–1291, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xu:2023:SGF

[XLWG23] Ganggang Xu, Chen Liang, Rasmus Waagepetersen, and Yongtao Guan. Semiparametric goodness-of-fit test for clustered point processes with a shape-constrained pair correlation function. Journal of the American Statistical Association, 118(543):2072–2087, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xie:2020:CAV

[XLYT20] Jinhan Xie, Yuanyuan Lin, Xiaodong Yan, and Niansheng Tang. Category-adaptive variable screening for ultra-high dimensional heterogeneous categorical data. Journal of the American Statistical Association, 115(530):747–760, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xue:2021:IMB

[XQ21] Fei Xue and Annie Qu. Integrating multisource block-wise missing data in model selection. Journal of the American Statistical Association, 116(536):1914–1927, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xie:2021:NEC

[XS21] Bingying Xie and Jun Shao. Nonparametric estimation of conditional expectation with auxiliary information and dimension reduction. Journal of the American Statistical Association, 116(535):1346–1357, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xie:2022:DCI

[XS22] Dongyue Xie and Matthew Stephens. Discussion of “Confidence Intervals for Nonparametric Empirical Bayes Analysis”. Journal of the American Statistical Association, 117 (539):1186–1191, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [IW22a] and rejoinder [IW22b].

Xue:2023:CIT

[XSP23] Haoran Xue, Xiaotong Shen, and Wei Pan. Causal inference in transcriptome-wide association studies with invalid instruments and GWAS summary data. Journal of the American Statistical Association, 118(543):1525–1537, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xue:2022:HMA

[XTK^{+}22] Fei Xue, Xiwei Tang, Grace Kim, Karestan C. Koenen, Chantel L. Martin, Sandro Galea, Derek Wildman, Monica Uddin, and Annie Qu. Heterogeneous mediation analysis on epigenomic PTSD and traumatic stress in a predominantly African American cohort. Journal of the American Statistical Association, 117(540):1669–1683, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xiao:2024:MAL

[XWMD24] Qian Xiao, Yaping Wang, Abhyuday Mandal, and Xinwei Deng. Modeling and active learning for experiments with quantitative-sequence factors. Journal of the American Statistical Association, 119(545):407–421, 2024. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xie:2020:BRG

[XX20] Fangzheng Xie and Yanxun Xu. Bayesian repulsive Gaussian mixture model. Journal of the American Statistical Association, 115(529):187–203, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xie:2021:BPC

[XX21] Fangzheng Xie and Yanxun Xu. Bayesian projected calibration of computer models. Journal of the American Statistical Association, 116(536):1965–1982, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xie:2023:EER

[XX23] Fangzheng Xie and Yanxun Xu. Efficient estimation for random dot product graphs via a one-step procedure. Journal of the American Statistical Association, 118(541):651–664, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xu:2024:CUS

[XYWQ24] Qi Xu, Yubai Yuan, Junhui Wang, and Annie Qu. Crowdsourcing utilizing subgroup structure of latent factor modeling. Journal of the American Statistical Association, 119 (546):1192–1204, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xue:2024:OLD

[XYY24] Kaijie Xue, Jin Yang, and Fang Yao. Optimal linear discriminant analysis for high-dimensional functional data. Journal of the American Statistical Association, 119(546):1055–1064, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xing:2023:CFD

[XZL23] Xin Xing, Zhigen Zhao, and Jun S. Liu. Controlling false discovery rate using Gaussian mirrors. Journal of the American Statistical Association, 118(541):222–241, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xu:2024:BCT

[XZLG24] Ganggang Xu, Jingfei Zhang, Yehua Li, and Yongtao Guan. Bias-correction and test for mark-point dependence with replicated marked point processes. Journal of the American Statistical Association, 119(545):217–231, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xia:2021:ICT

[XZWT21] Lucy Xia, Richard Zhao, Yanhui Wu, and Xin Tong. Intentional control of Type I error over unconscious data distortion: a Neyman–Pearson approach to text classification. Journal of the American Statistical Association, 116(533):68–81, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xue:2022:MAB

[XZZ^{+}22] Fei Xue, Yanqing Zhang, Wenzhuo Zhou, Haoda Fu, and Annie Qu. Multicategory angle-based learning for estimating optimal dynamic treatment regimes with censored data. Journal of the American Statistical Association, 117(539):1438–1451, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yao:2023:OEF

[Yao23] Ying Yang Fang Yao. Online estimation for functional data. Journal of the American Statistical Association, 118(543): 1630–1644, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yu:2020:DSP

[YB20] Bin Yu and Rebecca Barter. The data science process: One culture. Journal of the American Statistical Association, 115 (530):672–674, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yan:2021:RFS

[YB21] Xiaohan Yan and Jacob Bien. Rare feature selection in high dimensions. Journal of the American Statistical Association, 116(534):887–900, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yu:2024:VBF

[YB24] Weichang Yu and Howard D. Bondell. Variational Bayes for fast and accurate empirical likelihood inference. Journal of the American Statistical Association, 119(546):1089–1101, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yang:2020:QFS

[YBRM20] Hojin Yang, Veerabhadran Baladandayuthapani, Arvind U. K. Rao, and Jeffrey S. Morris. Quantile function on scalar regression analysis for distributional data. Journal of the American Statistical Association, 115(529):90–106, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yang:2020:CMO

[YD20] Shu Yang and Peng Ding. Combining multiple observational data sources to estimate causal effects. Journal of the American Statistical Association, 115(531):1540–1554, 2020. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yang:2024:EBA

[YDV24] Chun-Hao Yang, Hani Doss, and Baba C. Vemuri. An empirical Bayes approach to shrinkage estimation on the manifold of symmetric positive-definite matrices. Journal of the American Statistical Association, 119(545):259–272, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ye:2024:BRF

[Ye24] Ting Ye. Book review: Fundamentals of Causal Inference: With R, Babette A. Brumback, Boca Raton, FL: Chapman & Hall/CRC Press, 2022, xii + 236 pp., $74.95(H), ISBN 978-0-367-70505-3. Journal of the American Statistical Association, 119(545):790–791, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yee:2022:HDE

[Yee22] Thomas W. Yee. On the hauck-donner effect in Wald tests: Detection, tipping points, and parameter space characterization. Journal of the American Statistical Association, 117 (540):1763–1774, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yang:2020:NEC

[YFZ20] Lu Yang, Edward W. Frees, and Zhengjun Zhang. Nonparametric estimation of copula regression models with discrete outcomes. Journal of the American Statistical Association, 115(530):707–720, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yang:2020:TPU

[YLPC20] Bingduo Yang, Wei Long, Liang Peng, and Zongwu Cai. Testing the predictability of U.S. housing price index returns based on an IVX-AR model. Journal of the American Statistical Association, 115(532):1598–1619, 2020. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yu:2020:OSL

[YLSL20] Guan Yu, Quefeng Li, Dinggang Shen, and Yufeng Liu. Optimal sparse linear prediction for block-missing multi-modality data without imputation. Journal of the American Statistical Association, 115(531):1406–1419, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yu:2024:FCP

[YLX24] Xiufan Yu, Danning Li, and Lingzhou Xue. Fisher’s combined probability test for high-dimensional covariance matrices. Journal of the American Statistical Association, 119 (545):511–524, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yu:2023:PES

[YLXL23] Xiufan Yu, Danning Li, Lingzhou Xue, and Runze Li. Powerenhanced simultaneous test of high-dimensional mean vectors and covariance matrices with application to gene-set testing. Journal of the American Statistical Association, 118(544): 2548–2561, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yi:2021:FUU

[YNCK21] Dingdong Yi, Shaoyang Ning, Chia-Jung Chang, and S. C. Kou. Forecasting unemployment using Internet search data via PRISM. Journal of the American Statistical Association, 116(536):1662–1673, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yon:2023:PMS

[Yon23] George G. Vega Yon. Power and multicollinearity in small networks: a discussion of “Tale of two datasets: Representativeness and generalisability of inference for samples of networks” by krivitsky, coletti, and hens. Journal of the American Statistical Association, 118(544):2228–2231, 2023. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [KCH23b, KCH23a].

Yadlowsky:2021:EVR

[YPL^{+}21] Steve Yadlowsky, Fabio Pellegrini, Federica Lionetto, Stefan Braune, and Lu Tian. Estimation and validation of ratiobased conditional average treatment effects using observational data. Journal of the American Statistical Association, 116(533):335–352, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yuan:2023:HOJ

[YQ23] Yubai Yuan and Annie Qu. High-order joint embedding for multi-level link prediction. Journal of the American Statistical Association, 118(543):1692–1706, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yang:2023:RSR

[YQL23] Zihao Yang, Tianyi Qu, and Xinran Li. Rejective sampling, rerandomization, and regression adjustment in survey experiments. Journal of the American Statistical Association, 118 (542):1207–1221, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yao:2023:AEC

[YRTJ23] Shunan Yao, Bradley Rava, Xin Tong, and Gareth James. Asymmetric error control under imperfect supervision: a label-noise-adjusted Neyman–Pearson umbrella algorithm. Journal of the American Statistical Association, 118(543): 1824–1836, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yan:2021:CRC

[YS21] Bowei Yan and Purnamrita Sarkar. Covariate regularized community detection in sparse graphs. Journal of the American Statistical Association, 116(534):734–745, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ye:2024:OIS

[YSL24] Shengbin Ye, Thomas P. Senftle, and Meng Li. Operatorinduced structural variable selection for identifying materials genes. Journal of the American Statistical Association, 119(545):81–94, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yu:2022:SSS

[YSLW22] Jiahui Yu, Jian Shi, Anna Liu, and Yuedong Wang. Smoothing spline semiparametric density models. Journal of the American Statistical Association, 117(537):237–250, 2022.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yin:2024:CSA

[YSWB24] Mingzhang Yin, Claudia Shi, Yixin Wang, and David M. Blei. Conformal sensitivity analysis for individual treatment effects. Journal of the American Statistical Association, 119 (545):122–135, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ye:2023:TBP

[YSYZ23] Ting Ye, Jun Shao, Yanyao Yi, and Qingyuan Zhao. Toward better practice of covariate adjustment in analyzing randomized clinical trials. Journal of the American Statistical Association, 118(544):2370–2382, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yu:2022:ODS

[YWAZ22] Jun Yu, HaiYing Wang, Mingyao Ai, and Huiming Zhang. Optimal distributed subsampling for maximum quasilikelihood estimators with massive data. Journal of the American Statistical Association, 117(537):265–276, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yu:2022:MGI

[YWKZ22] Dengdeng Yu, Linbo Wang, Dehan Kong, and Hongtu Zhu. Mapping the genetic-imaging-clinical pathway with applications to Alzheimer’s disease. Journal of the American Statistical Association, 117(540):1656–1668, 2022. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yu:2020:EIG

[YWW^{+}20] Shan Yu, Guannan Wang, Li Wang, Chenhui Liu, and Lijian Yang. Estimation and inference for generalized geoadditive models. Journal of the American Statistical Association, 115(530):761–774, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yu:2020:CIS

[YYLL20] Guan Yu, Liang Yin, Shu Lu, and Yufeng Liu. Confidence intervals for sparse penalized regression with random designs. Journal of the American Statistical Association, 115(530):

794–809, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yang:2024:OSB

[YYZ24] Ying Yang, Fang Yao, and Peng Zhao. Online smooth backfitting for generalized additive models. Journal of the American Statistical Association, 119(546):1215–1228, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ying:2020:RAD

[YYZZ20] Zhiliang Ying, Wen Yu, Ziqiang Zhao, and Ming Zheng. Regression analysis of doubly truncated data. Journal of the American Statistical Association, 115(530):810–821, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yao:2020:PBR

[YZ20] Zhigang Yao and Zhenyue Zhang. Principal boundary on Riemannian manifolds. Journal of the American Statistical Association, 115(531):1435–1448, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yang:2022:SIU

[YZ22] Jun Yang and Zhou Zhou. Spectral inference under complex temporal dynamics. Journal of the American Statistical Association, 117(537):133–155, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yun:2022:DLD

[YZL22] Sooin Yun, Xianyang Zhang, and Bo Li. Detection of local differences in spatial characteristics between two spatiotemporal random fields. Journal of the American Statistical Association, 117(537):291–306, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yang:2023:EKL

[YZZ23] Guangyu Yang, Baqun Zhang, and Min Zhang. Estimation of knots in linear spline models. Journal of the American Statistical Association, 118(541):639–650, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zanella:2020:IPL

[Zan20] Giacomo Zanella. Informed proposals for local MCMC in discrete spaces. Journal of the American Statistical Association, 115(530):852–865, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhao:2022:RMA

[ZB22] Sihai Dave Zhao and William Biscarri. A regression modeling approach to structured shrinkage estimation. Journal of the American Statistical Association, 117(540):1684–1694, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2024:BLB

[ZBM^{+}24] Cong Zhang, Tejasv Bedi, Chul Moon, Yang Xie, Min Chen, and Qiwei Li. Bayesian landmark-based shape analysis of tumor pathology images. Journal of the American Statistical Association, 119(546):798–810, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2022:CAF

[ZC22] Xianyang Zhang and Jun Chen. Covariate adaptive false discovery rate control with applications to omics-wide multiple testing. Journal of the American Statistical Association, 117(537):411–427, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhu:2022:NFV

[ZCM22] Xuening Zhu, Zhanrui Cai, and Yanyuan Ma. Network functional varying coefficient model. Journal of the American Statistical Association, 117(540):2074–2085, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhu:2023:OCM

[ZCW23] Wanrong Zhu, Xi Chen, and Wei Biao Wu. Online covariance matrix estimation in stochastic gradient descent. Journal of the American Statistical Association, 118(541):393–404, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhao:2023:GUB

[ZCZZ23] Yize Zhao, Changgee Chang, Jingwen Zhang, and Zhengwu Zhang. Genetic underpinnings of brain structural connectome for Young adults. Journal of the American Statistical Association, 118(543):1473–1487, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhao:2024:AAE

[ZD24] Anqi Zhao and Peng Ding. To adjust or not to adjust? estimating the average treatment effect in randomized experiments with missing covariates. Journal of the American Statistical Association, 119(545):450–460, 2024. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2021:NPT

[ZDM21] Cheng Zhang, Vu Dinh, and Frederick A. Matsen Iv. Nonbifurcating phylogenetic tree inference via the adaptive LASSO. Journal of the American Statistical Association, 116(534): 858–873, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zheng:2023:GIA

[ZDNH23] Jiayin Zheng, Xinyuan Dong, Christina C. Newton, and Li Hsu. A generalized integration approach to association analysis with multi-category outcome: an application to a tumor sequencing study of colorectal cancer and smoking. Journal of the American Statistical Association, 118(541):29–42, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2023:CPG

[ZDY23] Jiawei Zhang, Jie Ding, and Yuhong Yang. Is a classification procedure good enough? — a goodness-of-fit assessment tool for classification learning. Journal of the American Statistical Association, 118(542):1115–1125, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2020:STS

[ZGZC20] Jin-Ting Zhang, Jia Guo, Bu Zhou, and Ming-Yen Cheng. A simple two-sample test in high dimensions based on L^{2} - norm. Journal of the American Statistical Association, 115 (530):1011–1027, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhao:2020:BRE

[Zha20] Yichuan Zhao. Book review: Empirical Likelihood Methods in Biomedicine and Health, by Albert Vexler and Jihnhee Yu. Boca Raton, FL: Chapman & Hall/CRC Press, 2018, xviii + 299 pp., $174.95(H), ISBN: 978-1-46-655503-7. Journal of the American Statistical Association, 115(530):1028–1029, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhao:2022:BRD

[Zha22] Ying-Qi Zhao. Book review: Dynamic Treatment Regimes: Statistical Methods for Precision Medicine Anastasios A. Tsiatis, Marie Davidian, Shannon T. Holloway, and Eric B. Laber, Boca Raton, FL, Chapman & Hall/CRC Press, 2019, xv + 602 pp., $99.95(H), ISBN 978-1-49-876977-8. Journal of the American Statistical Association, 117(537):527, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2021:BHC

[ZHC21] Jincheng Zhou, James S. Hodges, and Haitao Chu. A Bayesian hierarchical CACE model accounting for incomplete noncompliance with application to a meta-analysis of epidural analgesia on cesarean section. Journal of the American Statistical Association, 116(536):1700–1712, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2022:DCD

[ZHW22] Jingnan Zhang, Xin He, and Junhui Wang. Directed community detection with network embedding. Journal of the American Statistical Association, 117(540):1809–1819, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhu:2023:PMM

[ZJLD23] Wanchuang Zhu, Yingkai Jiang, Jun S. Liu, and Ke Deng. Partition-Mallows model and its inference for rank aggregation. Journal of the American Statistical Association, 118 (541):343–359, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2023:DGA

[ZJLH23] Xingyu Zhou, Yuling Jiao, Jin Liu, and Jian Huang. A deep generative approach to conditional sampling. Journal of the American Statistical Association, 118(543):1837–1848, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2023:DLL

[ZK23] Daiwei Zhang and Jian Kang. Discussion of “LESA: Longitudinal Elastic Shape Analysis of Brain Subcortical Structures”. Journal of the American Statistical Association, 118(541):22– 24, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [ZWX^{+}23a].

Zhang:2023:BIQ

[ZKR23] Tao Zhang, Kengo Kato, and David Ruppert. Bootstrap inference for quantile-based modal regression. Journal of the American Statistical Association, 118(541):122–134, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zou:2022:ELR

[ZKZ22] Changliang Zou, Yuan Ke, and Wenyang Zhang. Estimation of low rank high-dimensional multivariate linear models for multi-response data. Journal of the American Statistical Association, 117(538):693–703, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2023:HDG

[ZL23] Jingfei Zhang and Yi Li. High-dimensional Gaussian graphical regression models with covariates. Journal of the American Statistical Association, 118(543):2088–2100, 2023. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhu:2023:CTI

[ZLD23] Yichen Zhu, Cheng Li, and David B. Dunson. Classification trees for imbalanced data: Surface-to-volume regularization. Journal of the American Statistical Association, 118 (543):1707–1717, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zu:2023:UHD

[ZLGY23] Tianhai Zu, Heng Lian, Brittany Green, and Yan Yu. Ultrahigh dimensional quantile regression for longitudinal data: an application to blood pressure analysis. Journal of the American Statistical Association, 118(541):97–108, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2021:GBE

[ZLQ^{+}21] Fan Zhou, Shikai Luo, Xiaohu Qie, Jieping Ye, and Hongtu Zhu. Graph-based equilibrium metrics for dynamic supplydemand systems with applications to ride-sourcing platforms. Journal of the American Statistical Association, 116(536): 1688–1699, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhong:2023:MFV

[ZLZ23a] Wenxuan Zhong, Yiwen Liu, and Peng Zeng. A model-free variable screening method based on leverage score. Journal of the American Statistical Association, 118(541):135–146, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2023:HCS

[ZLZ23b] Kun Zhou, Ker-Chau Li, and Qing Zhou. Honest confidence sets for high-dimensional regression by projection and shrinkage. Journal of the American Statistical Association, 118 (541):469–488, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2024:LCH

[ZLZ24] Xin Zhang, Jia Liu, and Zhengyuan Zhu. Learning coefficient heterogeneity over networks: a distributed spanning-tree-based fused-lasso regression. Journal of the American Statistical Association, 119(545):485–497, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhao:2022:VEP

[ZM22] Jiwei Zhao and Yanyuan Ma. A versatile estimation procedure without estimating the nonignorable missingness mechanism. Journal of the American Statistical Association, 117 (540):1916–1930, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2023:CSF

[ZM23] Qingzhao Zhang and Shuangge Ma. Comment on “A Scale-Free Approach for False Discovery Rate Control in Generalized Linear Models” by Chengguang Dai, Buyu Lin, Xin Xing, and Jun S. Liu. Journal of the American Statistical Association, 118(543):1566–1568, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [DLXL23c].

Zammit-Mangion:2022:DCS

[ZMNVF22] Andrew Zammit-Mangion, Tin Lok James Ng, Quan Vu, and Maurizio Filippone. Deep compositional spatial models. Journal of the American Statistical Association, 117(540):1787– 1808, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zeng:2024:SEA

[ZMZ24] Jing Zeng, Qing Mai, and Xin Zhang. Subspace estimation with automatic dimension and variable selection in sufficient dimension reduction. Journal of the American Statistical Association, 119(545):343–355, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2022:BRU

[ZNBR22] Yan Dora Zhang, Brian P. Naughton, Howard D. Bondell, and Brian J. Reich. Bayesian regression using a prior on the model fit: The R2-2D shrinkage prior. Journal of the American Statistical Association, 117(538):862–874, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2021:DCT

[ZP21] Bo Zhang and Hongming Pu. Discussion of Cui and Tchetgen Tchetgen (2020) and Qiu et al. (2020). Journal of the American Statistical Association, 116(533):196–199, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2024:FMC

[ZPYZ24] Bo Zhang, Guangming Pan, Qiwei Yao, and Wang Zhou. Factor modeling for clustering high-dimensional time series. Journal of the American Statistical Association, 119(546): 1252–1263, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zeng:2023:ZIL

[ZPZW23] Yanyan Zeng, Daolin Pang, Hongyu Zhao, and Tao Wang. A zero-inflated logistic normal multinomial model for extracting microbial compositions. Journal of the American Statistical Association, 118(544):2356–2369, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhong:2023:FSI

[ZQL^{+}23] Wei Zhong, Chen Qian, Wanjun Liu, Liping Zhu, and Runze Li. Feature screening for interval-valued response with application to study association between posted salary and required skills. Journal of the American Statistical Association, 118(542):805–817, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zito:2023:BMS

[ZROD23] Alessandro Zito, Tommaso Rigon, Otso Ovaskainen, and David B. Dunson. Bayesian modeling of sequential discoveries. Journal of the American Statistical Association, 118 (544):2521–2532, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2024:MSP

[ZRPB24] Shuang Zhou, Pallavi Ray, Debdeep Pati, and Anirban Bhattacharya. A mass-shifting phenomenon of truncated multivariate normal priors. Journal of the American Statistical Association, 119(545):582–596, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhu:2023:CMC

[ZSJW23] Yunzhang Zhu, Xiaotong Shen, Hui Jiang, and Wing Hung Wong. Collaborative multilabel classification. Journal of the American Statistical Association, 118(542):913–924, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2023:CEM

[ZSK23] Yichi Zhang, Weining Shen, and Dehan Kong. Covariance estimation for matrix-valued data. Journal of the American Statistical Association, 118(544):2620–2631, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2020:MET

[ZSL20] Jingfei Zhang, Will Wei Sun, and Lexin Li. Mixed-effect timevarying network model and application in brain connectivity analysis. Journal of the American Statistical Association, 115 (532):2022–2036, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2023:MOS

[ZSL^{+}23] B. Zhang, D. S. Small, K. B. Lasater, M. McHugh, J. H. Silber, and P. R. Rosenbaum. Matching one sample according to two criteria in observational studies. Journal of the American Statistical Association, 118(542):1140–1151, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhu:2020:HDC

[ZSP20] Yunzhang Zhu, Xiaotong Shen, and Wei Pan. On highdimensional constrained maximum likelihood inference. Journal of the American Statistical Association, 115(529):217– 230, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2022:HTS

[ZSW22] Likun Zhang, Benjamin A. Shaby, and Jennifer L. Wadsworth. Hierarchical transformed scale mixtures for flexible modeling of spatial extremes on datasets with many locations. Journal of the American Statistical Association, 117(539):1357–1369, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2023:POD

[ZSZL23] Jie Zhou, Will Wei Sun, Jingfei Zhang, and Lexin Li. Partially observed dynamic tensor response regression. Journal of the American Statistical Association, 118(541):424–439, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zheng:2020:NEM

[ZW20] Chaowen Zheng and Yichao Wu. Nonparametric estimation of multivariate mixtures. Journal of the American Statistical Association, 115(531):1456–1471, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhen:2023:CDG

[ZW23] Yaoming Zhen and Junhui Wang. Community detection in general hypergraph via graph embedding. Journal of the American Statistical Association, 118(543):1620–1629, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2022:HDS

[ZWKZ22] Zhengwu Zhang, Xiao Wang, Linglong Kong, and Hongtu Zhu. High-dimensional spatial quantile function-on-scalar regression. Journal of the American Statistical Association, 117 (539):1563–1578, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhao:2020:SEV

[ZWLS20] Hui Zhao, Qiwei Wu, Gang Li, and Jianguo Sun. Simultaneous estimation and variable selection for interval-censored data with broken adaptive ridge regression. Journal of the American Statistical Association, 115(529):204–216, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2022:AIC

[ZWS22] Yangfan Zhang, Runmin Wang, and Xiaofeng Shao. Adaptive inference for change points in high-dimensional data. Journal of the American Statistical Association, 117(540):1751–1762, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2021:SRI

[ZWSZ21] Bo Zhang, Jordan Weiss, Dylan S. Small, and Qingyuan Zhao. Selecting and ranking individualized treatment rules with unmeasured confounding. Journal of the American Statistical Association, 116(533):295–308, 2021. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2023:TIL

[ZWSZ23] Yu Zhou, Lan Wang, Rui Song, and Tuoyi Zhao. Transformationinvariant learning of optimal individualized decision rules with time-to-event outcomes. Journal of the American Statistical Association, 118(544):2632–2644, 2023. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2023:LLE

[ZWX^{+}23a] Zhengwu Zhang, Yuexuan Wu, Di Xiong, Joseph G. Ibrahim, Anuj Srivastava, and Hongtu Zhu. LESA: Longitudinal elastic shape analysis of brain subcortical structures. Journal of the American Statistical Association, 118(541):3–17, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See discussion [AL23, CHDP23, ZK23] and rejoinder [ZWX^{+}23b].

Zhang:2023:RLL

[ZWX^{+}23b] Zhengwu Zhang, Yuexuan Wu, Di Xiong, Joseph G. Ibrahim, Anuj Srivastava, and Hongtu Zhu. Rejoinder: LESA: Longitudinal elastic shape analysis of brain subcortical structures. Journal of the American Statistical Association, 118(541): 25–28, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [ZWX^{+}23a, ZWX^{+}23a, AL23, CHDP23, ZK23].

Zhang:2022:SIT

[ZWZ22] Yingying Zhang, Huixia Judy Wang, and Zhongyi Zhu. Single-index thresholding in quantile regression. Journal of the American Statistical Association, 117(540):2222–2237, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2024:SCR

[ZWZ24] Le Zhou, Boxiang Wang, and Hui Zou. Sparse convoluted rank regression in high dimensions. Journal of the American Statistical Association, 119(546):1500–1512, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2024:TSD

[ZX24] Ting Zhang and Beibei Xu. Tail spectral density estimation and its uncertainty quantification: Another look at tail dependent time series analysis. Journal of the American Statistical Association, 119(546):1424–1433, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2022:MHV

[ZXS^{+}22a] Yang Zhou, Lirong Xue, Zhengyu Shi, Libo Wu, and Jianqing Fan. Measuring housing vitality from multi-source big data and machine learning. Journal of the American Statistical Association, 117(539):1045–1059, 2022. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). See comments [TJK22, Ban22] and rejoinder [ZXS^{+}22b].

Zhou:2022:R

[ZXS^{+}22b] Yang Zhou, Lirong Xue, Zhengyu Shi, Libo Wu, and Jianqing Fan. Rejoinder. Journal of the American Statistical Association, 117(539):1066–1067, 2022. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [ZXS^{+}22a, TJK22, Ban22].

Zhao:2024:ETA

[ZYZ24] Bingxin Zhao, Xiaochen Yang, and Hongtu Zhu. Estimating trans-ancestry genetic correlation with unbalanced data resources. Journal of the American Statistical Association, 119(546):839–850, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhao:2022:GCE

[ZZ22] Bingxin Zhao and Hongtu Zhu. On genetic correlation estimation with summary statistics from genome-wide association studies. Journal of the American Statistical Association, 117 (537):1–11, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2023:WRT

[ZZ23a] Yao Zhang and Qingyuan Zhao. What is a randomization test? Journal of the American Statistical Association, 118 (544):2928–2942, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2023:CFR

[ZZ23b] Le Zhou and Hui Zou. Cross-fitted residual regression for high-dimensional heteroscedasticity pursuit. Journal of the American Statistical Association, 118(542):1056–1065, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zheng:2022:RPT

[ZZH22] Jiayin Zheng, Yingye Zheng, and Li Hsu. Risk projection for time-to-event outcome leveraging summary statistics with source individual-level data. Journal of the American Statistical Association, 117(540):2043–2055, 2022. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhao:2024:ELF

[ZZL24] Junlong Zhao, Yang Zhou, and Yufeng Liu. Estimation of linear functionals in high-dimensional linear models: From sparsity to nonsparsity. Journal of the American Statistical Association, 119(546):1579–1591, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2020:PMA

[ZZLC20] Xinyu Zhang, Guohua Zou, Hua Liang, and Raymond J. Carroll. Parsimonious model averaging with a diverging number of parameters. Journal of the American Statistical Association, 115(530):972–984, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2024:EOI

[ZZQ24] Wenzhuo Zhou, Ruoqing Zhu, and Annie Qu. Estimating optimal infinite horizon dynamic treatment regimes via pTlearning. Journal of the American Statistical Association, 119(545):625–638, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2020:MFF

[ZZXL20] Tingyou Zhou, Liping Zhu, Chen Xu, and Runze Li. Modelfree forward screening via cumulative divergence. Journal of the American Statistical Association, 115(531):1393–1405, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).